A Bibliography of Papers in Lecture Notes in Computer Science (2012): Volumes 7500–7549

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

14 October 2017 Version 1.01

Title word cross-reference

#W [1] [554]. 2 [53]. 2\frac{1}{2} [225]. 2^{n} [540]. 3 [135, 120, 214, 216, 211, 139, 110, 61, 202]. 4 [134]. ^{2} [430]. ^{T M} [305]. Audi [358]. β [686]. δ [529]. i [684]. k [538, 554, 41]. O [358]. O(1.89^{n}) [555]. T_{1} [81].

-Biomarkers [684]. -D [216]. -DoF [53]. -Hard [554]. -Matchings [554]. -Path [538]. -TRIMAX [529].

/Game [332]. /Therapy [332].

11th [727]. 128 [704]. 12th [712, 737]. 13th [724]. 14th [722]. 15th [716, 717, 718]. 16th [713]. 17th [740]. 19th [746].

2 [705, 706]. 2.0 [638]. 2012 [249]. 20th [321]. 2nd [736].

31st [735]. 35th [730]. 384-bit [462]. 3D [116, 47]. 3rd [721].

4th [731, 720, 733].

512 [705]. 5th [714].

6th [725, 734].

7th [721, 738].

8th [745, 742].

9th [741, 719, 726].

= [496].

AAL [233]. ABA [618]. Abdominal [86]. Abnormality [94]. Absolute [90]. Abstract [280]. Abuse [16]. Accelerated [127]. Acceptance [199]. Access [237, 463, 245]. Accounting [526]. Accuracy [79]. Accurate [530]. Achievements [170]. ACIVS [722]. ACL [605]. ACNS [734]. Acquisition [222, 132]. Action [369]. Activation [393, 68]. Active [365, 213, 460]. Activities [202]. Activity [98, 68]. Actually [481]. Actuator [54]. AD [101]. Adaptation [595]. Adaptations [239]. Adaption [418]. Adaptive [127, 179, 94, 364, 150, 606]. Adaptivity [193]. ADBIS [713]. Additive [707]. Adjusting [219]. ADNTIIC [744]. Adolescent [124]. Adoption [419]. Adults [335, 66]. Advanced [470, 488, 722]. Advances [169, 730, 723, 744, 713]. Adventure [205]. Adventures [207]. Adversaries [460]. Advertising [574]. Affect [13]. Affective [9]. after [317]. against [472, 581, 691]. Age [335]. Age-Related [335]. Agent [262, 7, 742, 4, 6, 655, 608, 609, 612, 741]. Agents [13, 11, 602, 10, 607, 618, 623, 712, 614, 12, 620]. Aggregate [35]. Aggregation [383, 398, 161]. Agreement [509, 461]. Agreements [234]. AI [730, 391]. AICI [733]. Airway [145]. Alcazar [640]. Alert [473]. Algebra [568]. Algorithm [549, 543, 81, 708, 353, 55, 547, 59, 35, 681, 437, 230, 525, 445, 439, 156, 377, 656, 685, 166]. Algorithms [551, 345, 180, 36, 40, 181, 539, 149, 182, 528, 737]. Aliasing [300]. ALignment [30, 528, 115]. Alignments [176, 527]. Almost [702]. alpha [442]. ALS-Based [362]. Alternating [297]. Alternative [465, 70]. Alzheimer [65]. Am [137]. AmbiLearn [433]. America [736]. Analyse [332]. Analysing [529]. Analysis [72, 250, 112, 440, 468, 159, 658, 104, 466, 654, 636, 36, 162, 666, 161, 251, 300, 589, 203, 239, 225, 151, 4, 707, 453, 446, 715]. Analytic [121]. Analytical [37]. Analytics [477, 340]. Analyzing [479, 341]. Anaplastic [104]. Anatomical [145, 119, 139]. Anatomy [417]. Ancient [3]. Angular [80]. Animated [20]. Annotate [675]. Annotating [224]. Annotation [380, 688, 323]. Annotation-Based [380]. Annual [730]. Answer [1]. Answering [645, 285, 287, 4, 275]. Answers [152]. Anticipate [614]. Aortic [108]. Appearance [13, 120, 134]. Appearances [119]. Application [173, 107, 430, 133, 163, 706, 226, 69, 629, 195, 707, 308, 148]. Application-Level [148]. Applications [351, 584, 402, 633, 340, 695, 732, 714, 275, 694]. Applying [187, 486, 50]. Appraisals [6]. Apprentice [2]. Apprenticeship [372]. Approach [591, 674, 572, 233, 281, 494, 468, 593, 104, 120, 402, 373, 15, 438, 245, 293, 61, 236, 148]. Approaches [17, 50]. Approximability [542]. Approximate [312]. Approximation [551, 136]. APRIL [365]. Architecting [585]. Architectural [197]. Architecture [582, 188, 664, 665, 100]. Architectures [187, 586, 734, 248, 406, 30]. Argentina [744, 743]. Argentinization [648]. ArgMAS [742]. Arguing [623]. Argument [625]. Argumentation [279, 278, 255, 281, 280, 626, 604, 271, 607, 617, 618, 619, 609, 620, 277, 627, 628, 629, 621, 742]. Argumentation-Based [627]. Arguments [624]. Arithmetic [505]. Arm [49]. ARMADILLO2 [699]. Array [449]. Arrays [447]. Arterial [82]. Article [361]. Articulated [102]. Articulation [18]. Artificial [730, 724, 733]. ARX [707]. ASM [236]. ASM-Based [236]. Aspects [650, 432, 726]. Assembly [149, 164, 536]. Assessment [85, 431]. Assimilating [422]. Assist [50]. Assisted [716, 717, 718]. Assistive [208]. Association [573, 343, 453]. Assume [406]. Assume-Guarantee [406]. Assumption [503]. Assurance [220]. Assured [459]. Asynchronous [164]. Atelic [238]. Atlas [71]. Atlases [75, 86, 77]. Atomic [500]. Atrium [106]. Atrophy [67]. Attack [697, 706, 698]. Attacker [158]. Attacks [281, 692, 472, 705, 704, 703, 694]. Attempt [283]. Attitude [7]. Attribute [463, 117]. Attribute-Based [463]. Attribute-Guided [117]. Audio [20, 657, 14, 358]. Audio-Visual [657]. Authenticated [702, 513]. Authentication [449, 710, 709, 461, 160]. Authoring [205]. Autistic [5]. Automata [379, 297]. Automated [65, 1, 114]. Automatic [26, 645, 104, 135, 662, 227, 119, 75, 106, 480, 30]. Automatically [183]. Automation [61]. Automaton [296]. Autonomous [5, 580, 366]. Availability [411]. Averaging [130, 398]. Avoiding [407, 386]. Aware [331, 362, 360, 464, 156].

B [403, 121]. B-spline [121]. Back [20, 83, 314, 412, 515, 575, 597, 615, 630, 646, 689]. Bag [593]. Bag-Oriented [593]. Balance [655, 416]. Balancing [445, 156]. Bands [291]. Bangalore [726]. Bangladesh [643]. Barreto [498]. Base [397]. Based [504, 365, 279, 278, 591, 449, 174, 674, 250, 380, 193, 93, 574, 492, 233, 459, 680, 130, 430, 159, 363, 688, 535, 187, 215, 594, 46, 411, 720, 178, 469, 348, 116, 119, 588, 710, 362, 59, 134, 179, 66, 78, 463, 618, 437, 292, 230, 106, 163, 41, 490, 438, 443, 354, 111, 423, 333, 611, 118, 627, 390, 612, 501, 196, 652, 505, 371, 184, 471, 114, 195, 151, 447, 442, 4, 236, 448, 150, 49, 47, 394, 446, 460, 309, 186, 71, 444].

Batch [686]. Battery [370]. Bayesian [671, 374, 345, 373, 348, 133, 367, 346, 377]. before [514]. Behavior [13, 17, 45, 665, 18]. Behavioral [472, 475]. Behaviour [379]. Behavioural [432]. behind [489]. Being [425]. Belief [256, 288, 381, 289]. bent [499]. Betaville [427]. BetterRelations [573]. between [288, 258, 623, 438, 391, 474, 624]. beyond [550]. BI [723, 249, 251]. Bicliques [705]. Bidirectional [272, 376]. Big [477, 245, 439, 247]. Bilateral [217]. Bimodal [230]. Binary [349, 229]. Bio [45]. Bio-monitoring [45]. Bioinformatic [687]. Bioinformatics [745, 670, 673, 737, 675]. Biomarkers [684, 101]. Biomedical [678]. Biomolecular [678]. Biostatistics [745, 670]. Bisimilarity [258]. bit [462]. Bitvector [353]. Blackjack [391]. Bladder [684]. Blended [422]. Blind [462, 444]. Block [514]. Blocks [500]. Blood [82]. Bloom [535]. Blur [218]. Body [85, 19]. Boltzmann [110]. Bonding [57]. Bone [93, 100]. Bones [125]. Boolean [388]. Boosted [684]. Booster [307]. Boosting [184]. Bootstrapping [368]. Botnets [474]. Bounding [295, 503]. Bounds [349, 531]. Boxicity [551]. Brain [676, 66, 70, 113, 76, 117, 715, 143]. Brainstem [124]. Breakdown [67]. Breaking [160]. Breathing [103]. Bremen [727, 732]. Bridging [488]. Bristol [728, 729]. Brno [722]. Broadening [739]. Browser [241]. Browsing [657]. Bruijn [534, 535, 149, 536]. Bug [176]. Building [644, 5]. Bundling [148]. Business [30, 245, 185]. Bytecode [300].

C [24]. C5 [684]. CA [712]. Calculi [254]. Calibration [111, 213]. Calisthenics [384]. Camellia [696]. Camera [214]. Canada [714]. Cancer [684]. Canny [441]. Canonical [38]. Capturing [221]. Cardiac [109, 106, 111, 110]. Cards [171, 160]. Care [579]. Carlo [354, 368]. Case [686, 438, 410, 401]. Case/Control [686]. Cases [287, 403, 301]. Categorical [344]. Categorization [119, 350]. Catheter [108]. Catheterisation [109]. Causality [440]. CBTC [411]. Cellular [154]. Centered [645, 648, 208]. Centric [463, 236]. Century [321]. Cerebral [82]. Certain [285]. Challenge [600]. Challenges [577, 340, 432]. Chance [13]. Change [237]. Changes [335]. Channel [225]. Channels [686, 460]. Characterisation [76]. Characteristics [46]. Characterization [282, 100, 131, 225, 475]. Characterizing [452]. Characters [19, 531, 8]. Chat [389]. Check [514]. Check-before-Output [514]. Checker [306]. Checking [308, 309]. Chemical [532]. Chengdu [733]. Children [5, 334]. Chile [736]. China [733]. Chinese [166]. Chip [151, 685]. ChIP-Seq [685]. Chit [389]. Chit-Chat [389]. Chordal [547]. Choreographies [409]. CIBB [745]. Ciphers [514]. Ciphertext [501, 701]. CISD [430]. City [3]. Civil [414]. Class [385, 363, 98]. Classes [349]. Classical [40]. Classification [23, 352, 363, 373, 89, 350, 99, 79]. Classifications [448]. Classifier [360]. Classifiers [349, 346]. Classroom [422, 197]. Clean [287]. Client [236]. Client-Centric [236]. CLIIQ [530]. Clique [541]. Close [212]. Close-Up [212]. Closing [19]. Cloud [459, 653, 234, 236, 450]. Cloud-Based [459].

Clustering [566, 569, 119, 75, 486, 390, 451]. CMS [723, 232]. Co [157, 225, 18]. Co-Articulation [18]. Co-occurrence [225]. Co-simulation [157]. Coating [54]. Cockpits [402]. Code [212]. Coded [599]. Codes [369]. Coding [219]. Coefficient [552]. Cognition [482]. Cognitive [322, 67, 391, 64]. Collaborating [623]. Collaboration [59]. Collaborative [658, 660, 635, 626, 655, 190, 192, 656]. Collecting [573]. Collections [492, 533, 220]. Collision [706, 704]. Colloquium [726]. Color [686, 217, 448]. Color-Texture [448]. Column [484]. Column-Store [484]. Combining [65, 179, 347]. Command [580]. Commercial [457]. Commodity [171]. Common [544]. Communicability [744, 635, 663, 634, 636, 633]. Communication [594, 654, 611, 687, 202, 151]. Communication-Based [594]. Community [681]. Compact [388, 532]. Companion [12]. Comparative [253, 530, 76]. Comparing [533]. Comparison [533, 345, 361, 17, 391]. Comparisons [527]. Compatibility [524, 531]. Compatible [574]. Compensation [108]. Competence [13, 426, 430, 198, 421]. Competence-Based [430]. Competitive [186]. Compile [400]. Compile-Time [400]. Complementing [680]. Complete [500]. Complex [28, 592, 740, 593, 594, 606]. Complexities [4]. Complexity [259, 693, 262, 557, 546, 289, 589]. Components [468, 475]. Composite [407, 454]. Compositional [592, 296]. Comprehensive [246, 450]. Compressed [127, 533]. Compression [550, 223]. Computation [514, 15, 738, 519]. Computational [669, 733, 745]. Computer [716, 717, 718, 655, 734, 423, 435, 196, 474, 743, 734]. Computer-Assisted [716, 717, 718]. Computer-Based [196]. Computing [716, 717, 718, 739, 243, 408, 727, 555, 653, 719, 234, 9, 651, 726, 236, 148, 450]. Concept [453]. Concepts [722, 426, 321]. Conceptual [723, 479, 635, 489, 481, 495, 244, 232, 482, 480, 735]. Concurrency [195]. Concurrent [276, 180]. Conditional [254, 260, 289]. Conditionals [383]. Conditions [41]. Conference [735, 716, 717, 718, 722, 744, 728, 729, 730, 721, 727, 736, 725, 734, 733, 732, 713, 712, 719, 714, 724]. Confidence [291, 398]. Configurable [219]. Confirmation [602]. Conflict [255, 11, 523]. Conflict-Free [523]. Conflict-Tolerant [255]. Conformance [306]. Conjunction [715]. Conjunctive [285]. Connectivity [66, 78, 67]. Considering [281]. Consistency [286]. Constrained [295, 81]. Constraint [601, 184, 480, 394]. Constraint-Based [394]. Constraints [388]. Constructing [119, 77]. Construction [296, 212, 359]. Constructionist [417]. Constructor [304]. Constructors [654]. Contagion [8]. Containment [39]. Contaminant [293]. Content [605]. Contents [649, 638]. Context [494, 658, 362, 208, 99, 656]. Context-Aware [362]. Context-Sensitive [208]. Contexts [208]. Contextual [242]. Contextualized [424]. Continuous [288, 274, 392]. Contractibility [556]. Contracts [496]. Control [51, 686, 191, 463, 580, 443, 157, 43, 313, 416, 48, 61, 49, 47, 56]. Controls [662]. Conventional [237]. Conversation [16]. Conversational [4].

Conversion [95, 101]. Converting [706]. Convolution [216]. Cooperative [627]. Coordinate [614]. Coordinates [500]. Coordination [601, 18]. Cope [183]. Coprocessing [26]. Córdoba [743]. Coregulation [529]. Corporate [419]. Correcting [369]. Correction [218, 447]. Correctness [298]. Correlation [693, 473]. Correlations [479]. Cost [203]. Counterexample [591]. Counterexample-Based [591]. Countermeasure [469]. Counting [554]. Coupled [377]. Course [671]. Courses [656]. Cover [548]. Coverage [307]. CP [159]. criteria [625]. Critical [402]. Cross [491, 165, 225]. Cross-Channel [225]. Cross-Language [491]. Cross-Platform [165]. Crowds [572]. Cruz [712]. Cryptanalysis [693, 696, 695, 699, 691]. Cryptographic [506]. Cryptography [505]. Cryptology [736]. CScale [584]. CSPConCheck [306]. CT [135, 116, 100, 86]. Cube [221]. Cubes [479]. Cubic [121]. Cultural [743, 649, 634, 632]. Culture [649, 633]. Cup [92]. Cure [310]. Current [40]. Currents [113]. Curricular [194]. Curriculum [422]. Curves [500, 507, 498, 499]. Curvilinear [107]. Customized [472]. Cuts [228]. Cyber [577, 472, 157]. Cyber-Attacks [472]. Cyber-Physical [577, 157]. Cycling [703]. Czech [722, 321]. Czechoslovakia [321].

D [27, 135, 120, 214, 134, 216, 211, 139, 110, 61, 225, 202]. D-ReServE [27]. DAGs [479]. Damper [58]. Dark [317]. Darmstadt [721]. Data [483, 565, 250, 529, 559, 693, 563, 566, 233, 678, 569, 30, 31, 477, 570, 479, 338, 274, 670, 34, 243, 676, 237, 564, 116, 89, 573, 36, 73, 37, 38, 438, 305, 495, 245, 161, 560, 709, 347, 366, 439, 247, 344, 341, 94, 675, 532, 283, 393, 488, 685]. Data-Analysis [36]. Data-Driven [366]. Data-Intensive [243]. Database [72, 240, 38]. Databases [26, 286, 263, 276, 728, 729, 268, 373, 41, 713, 473, 238]. Datalog [258]. Dataset [65]. Datasets [114]. DC [746]. Dealing [426]. Debugging [167]. December [744]. Decidability [272]. Deciding [258]. Decision [366, 364, 6]. Decision-Making [366]. Declarative [338]. Decoding [46]. Decomposition [528, 229]. Deconfounding [68]. Decryption [485]. Deductive [286]. DEEN [681]. Deep [429]. Default [440]. Defeasible [271, 627]. Defect [55]. Defense [474]. Definite [300]. Definition [430, 613]. Definitorial [546]. Deformable [141, 114, 143]. Deformation [122, 121]. Deformations [97]. Degenerate [540]. del [745, 720, 257]. DEL-Sequents [257]. Delay [155]. Delegation [604]. Delta [586]. Delta-Oriented [586]. Democracy [641]. Demons [119, 113]. Dempster [158]. Dempster-Shafer [158]. Denoising [80]. Dense [141]. Density [390]. Dependability [408]. Dependencies [287]. Dependency [364]. Depth [230, 213]. Derived [657]. Descent [395]. Describing [637]. Descriptive [342]. Design [319, 93, 459, 311, 603, 188, 648, 335, 443, 666, 327, 687, 196, 61, 321, 309, 334]. Designing [4, 324, 206, 421]. Desktop [665]. Destructors [654]. Detecting [666]. Detection [282, 135, 65, 55, 681, 162, 163, 530, 441, 226, 101, 94, 92, 68, 102]. Detector [119, 100]. Deterministic [36, 167]. Developer [175]. Developing [191, 384, 88]. Development [426, 740, 645, 55, 404, 588, 52, 732, 567, 424, 401, 198, 446]. Devices [457]. DGraph [149]. Diagnosis [602, 282, 78, 44]. Dialogic [4]. Dialogue [609]. Dialogues [627, 625]. Dictionaries [127]. Dictionary [128]. Diffeomorphic [119, 117]. Difference [325]. Differences [139, 707]. Different [209, 416, 436]. Different-Level [436]. Differential [437, 696, 695, 691, 707]. Differential-Linear [695]. Differentially [359]. Differentiated [46]. Diffusion [127, 85, 123, 74, 128, 129]. Diffusion-Weighted [129]. Diffusivity [85]. Digital [582, 194, 638, 335, 220, 462, 323, 641]. Dimensional [109, 566, 251]. Dimensions [483]. Direct [533]. Directed [555, 289]. Direction [74]. Discharging [545]. disciplinary [588]. disclosure [467]. Discovering [489, 673, 247, 342]. Discovery [678, 339, 728, 729, 343, 364, 488]. Discrete [666, 313, 124]. Discriminant [672]. Discussing [624]. Disease [65]. Diseases [670]. Disjoint [553]. Disorders [50]. Dispensing [56]. Displacement [141]. Display [328]. Displays [6]. Disruptions [225]. Distance [288, 503, 636, 490]. Distance-Based [490]. Distance-Bounding [503]. Distinction [238]. Distributed [601, 27, 593, 594, 584, 474]. Distribution [37, 460]. Distributions [291]. DL [605]. DMC [521]. dMPI [167]. DNA [533]. Do [70, 12]. Document [220, 239]. Documentation [613]. Documents [350, 664]. Does [72, 548]. DOF [60, 53]. Domain [483, 492, 430, 95, 32, 404, 410, 560, 589, 185, 565, 574]. Domain-Driven [185]. Domain-Specific [492, 589]. Domains [392]. dominating [542]. Don’t [414]. Drift [295, 282]. Drill [87]. Driven [572, 424, 366, 185, 452]. Driver [165]. DSP [215, 446]. DTI [65, 77]. Dual [398, 704]. Dual-Stream [704]. DUET [444]. Dummy [514]. DWI [80]. Dynamic [440, 153, 289, 495, 156, 621, 108, 53, 155]. Dynamical [282, 134]. Dynamics [669, 521].

e-Culture [649, 633]. e-Learning [663, 721, 635]. E-mail [657, 645]. EAP [128]. Early [67, 101]. East [713]. ECAs [16]. ECDM [723, 237]. ECDM-NoCoDA [723]. Echo [109]. Echocardiography [134]. ECML [728, 729]. Economic [574]. Edge [441, 542]. Education [721, 433, 417, 204]. Educational [654]. Edutainment [199, 721]. Effect [669, 14, 196, 16, 6]. Effectiveness [194]. Effects [674, 686, 223, 335, 68]. Efficiency [510, 159]. Efficient [568, 535, 32, 98, 710, 505, 343, 519, 92]. Efficiently [297]. eGovernment [641]. EHR [650]. Elastography [90]. Elderly [384]. Electric [370]. Electricity [366]. Electromagnetic [58]. Electrophysiology [110]. Elements [335]. ElimLin [708]. Elliptic [500]. Embedded [468, 215, 404, 179, 443]. Embedding [352, 354]. Embodied [19]. Embodiment [12]. Emotion [633, 6]. Emotional [8]. Emotions [10]. Empirically [657]. Employ [283]. Enable [88]. Encoding [291, 388]. Encounters [7]. Encryption [504, 502, 746, 702, 513]. End [585].

End-User [585]. Energy [32]. Enforcement [458]. Enforcing [466]. Engagement [421]. Engine [241]. Engineering [731, 174, 494, 658, 33, 659, 720, 181, 567, 661, 656]. Enhance [473]. Enhanced [307]. Enhancement [107, 129]. Enhancing [414, 433]. Enjoyment [199]. Enrichment [405]. Ensemble [354, 435]. Ensembles [119]. Ensuring [311]. Enterprise [665, 590]. Entertainment [324, 727]. Entity [287]. Entropy [267]. Enumeration [312]. Environment [27, 329, 433, 179, 495, 309]. Environmental [588]. Environments [233, 190, 148]. EP [110]. Epistemic [260, 264]. Equivalence [436]. Equivalences [383]. Era [634]. eRobotics [61]. Error [72, 144, 369]. Error-Correcting [369]. ESP [429]. Essential [88]. Estimating [522]. Estimation [81, 363, 144, 98, 62, 82, 390, 140, 218, 94]. Estimator [219]. EUMAS [741]. European [741, 728, 729, 713, 724, 65, 654]. Evaluating [174, 223, 432, 431]. Evaluation [319, 568, 221, 265, 411, 638, 493, 109, 157, 198, 302, 450]. Event [403, 470, 471]. Event-B [403]. Events [318]. Evidential [289]. Evolution [240, 595, 237, 437, 521]. Evolutionary [169, 181]. Evolving [172, 586, 284, 452, 186, 406]. Exact [535, 268, 547, 123, 738]. Example [404, 428]. Examples [670]. Exception [400]. Exceptional [375]. Exchange [308]. Execution [392]. Exergames [208, 203]. Exhaustive [375]. Exhibition [327]. Exome [675]. Exoskeleton [46, 48, 49]. Expansion [634]. Experience [209]. Experiences [559]. Experimental [345]. Experiments [671]. Exploit [267]. Exploiting [265, 40]. Exploration [569]. Explorations [570]. Exploratory [181]. Exponential [541]. Exponentiation [509]. Exponents [682]. Exporting [583]. Expose [176]. Exposing [457]. Expressed [530]. Expressing [605]. Expression [529, 300]. Expressions [480]. Expressive [435]. Extended [298]. Extending [568, 572, 349, 271, 495]. Extension [350, 637]. Extensions [297]. External [32]. Extract [485]. Extract-Transform-Load [485]. Extracting [529, 562, 523, 560]. Extraction [552, 561, 377]. Extractors [460]. Extreme [593]. Extremities [59]. Extremity [46]. Eye [212].

fAARS [331]. Face [20]. Facilitating [167]. Fact [33]. Fact-Oriented [33]. Factorization [369, 362]. Failure [408]. Fair [308]. Fairness [360, 308]. Fairness-Aware [360]. Family [702, 705]. FAS [407]. Fast [518, 520, 369, 353, 362, 681, 553, 123, 166, 709, 486, 110, 746]. Faster [507, 540, 685]. Fault [282, 402, 514, 401, 186]. Fault-Tolerant [402]. Faults [407]. Feature [676, 78, 181, 118, 637, 364, 448, 355]. Feature-Oriented [637]. Features [354]. Federation [650]. Feedback [547, 362, 313]. FES [51, 43]. FES-Involved [51]. Few [549]. Fiber [113, 129, 77]. Field [509, 57, 410]. Fields [500, 380, 121]. Filter [535, 41, 217]. Filtering [228, 480, 80]. Finalist [697]. Financial [578]. Finding [456, 540, 452, 685]. Fingerprint [532]. Finite [379]. Finite-State [379]. First [13, 7, 634, 268, 243]. First-Order [268]. Fish [222]. Fish-Inspired [222].

Fissure [115]. Fixed [542]. Flexible [275, 60, 88]. Florence [735, 723]. Flow [468, 466, 82, 69, 56, 124]. Fluoroscopy [107]. fMRI [68, 77]. fMRI-Guided [77]. FO [24]. Focal [447]. FOCUS [24]. Focusing [208]. Folded [679]. Follow [328]. Foolproof [702]. Forbes.com [361]. Forbid [662]. Forecast [32]. Forensic [159]. Forests [135, 136]. Form [214]. Formal [380, 401]. Formalization [259]. Formation [656]. Foundations [585, 484]. FPS [325]. Fractured [293]. Fragment [613]. Framework [382, 227, 271, 588, 165, 358, 118, 560, 452, 30]. Frameworks [255, 617, 277, 621]. France [716, 717, 718, 715, 724]. Free [523, 87, 214, 545]. Free-Form [214]. Frequency [46]. Frequent [343]. Front [21, 42, 558, 576, 598, 616, 631, 647, 668, 690, 711, 63, 84, 105, 126, 147, 168, 189, 210, 231, 252, 273, 294, 315, 336, 357, 378, 399, 413, 434, 455, 476, 497, 516, 537]. FSE [746]. Full [692]. Fully [1]. Function [72, 503, 103, 704, 44, 698, 444]. Functional [101, 77]. Functions [288, 289, 499, 709]. Furhat [20]. Fusion [562, 73, 75, 326, 620, 71]. Fuzzy [98, 617, 62, 284, 283, 450].

GA [175]. Gain [214]. Gait [446]. Game [600, 319, 193, 322, 429, 430, 335, 199, 573, 200, 427, 332, 422, 423, 333, 416, 197, 196, 321, 428, 192, 206]. Game-Based [193, 333]. GameDays [721]. Games [419, 426, 194, 325, 310, 381, 420, 415, 335, 721, 331, 642, 732, 332, 422, 205, 432, 204, 418, 425, 192, 334, 421]. Gaming [201, 209, 317]. Gaps [488]. Garda [745, 720]. Gargnano [745]. Garner [166]. Gathering [318]. Gaussian [444]. GCM [703]. Gender [16]. Gene [529, 525, 526]. General [381, 385]. Generalization [349]. Generalized [257]. Generalizing [277]. Generate [301]. Generating [10]. Generation [177, 659, 609, 1, 439, 234]. Generative [73]. Generic [375]. Genes [669]. Genetic [670, 180, 101, 283]. Genetics [66]. Genome [164, 536]. Genomic [678, 688, 670]. genre [319]. Geodesic [130, 145]. Geographical [638]. Geometry [196, 131, 113, 132, 146]. German [730]. Germany [730, 721, 727, 725, 732]. Gesture [89]. Get [34]. GHASH [703]. Gibbs [120]. Girly [199]. Giving [320]. Gland [99]. Global [578, 120, 634]. Globally [141]. GMM [116]. Goal [183]. Goals [258, 496]. Goats [428]. Good [72]. Google [562]. GOSPL [33]. GOST [692]. Governance [612]. GPGPU [395]. GPU [216]. Gradient [608, 74, 442, 395]. Gradual [212]. Grande [744]. Graph [535, 522, 228, 149]. Graphical [289, 677]. Graphics [165, 171]. Graphs [556, 295, 534, 547, 524, 376, 545, 251, 536, 218]. Grass [328]. Grey [437]. Grid [637]. Grids [390, 665]. Grøstl [697, 698]. Group [660, 152, 486]. Groups [242]. Groupwise [142, 79, 146, 125]. Guarantee [406]. Guarantees [343]. Guidance [87, 109]. Guide [107]. Guide-Wire [107]. Guided [451, 117, 77]. Guidelines [657]. Guides [644]. Gwangju [719]. GWAP [429].

H [545]. H-Minor [545]. Half [520]. Half-Sibling [520]. Handling [400]. Hard [554]. HARDI [66, 75]. HARDI-Based [66]. Hardware [219, 505].

Harmonic [292]. Hash [704, 709, 698]. Hashes [703]. Hashing [518, 498, 552, 700]. HCITOCH [743]. Head&Neck [114]. Health [415, 721, 425]. Healthcare [410]. Held [715]. Help [548]. Heritage [743, 634, 632]. Heterogeneous [670, 154, 148]. Heuristic [520, 368]. HEX [265]. HEX-Program [265]. Hidden [673]. Hierarchical [504, 296, 479, 145, 151, 117, 102]. Hierarchically [86]. Hierarchy [550]. High [193, 194, 513, 54, 118, 171, 218]. High-Power [54]. High-Resolution [193]. High-Schools [194]. Highdimensional [670]. Hippocampal [133]. Histograms [359]. History [3, 639, 321]. HOG [226]. Home [443]. Homogeneous [154]. Homomorphic [552]. Homomorphism [620]. Honeynet [473]. Honeypot [473]. Honeypot/Honeynet [473]. HTML5 [204]. Huerta [744]. Human [59, 15, 14, 62, 423, 45, 50, 226, 393, 186, 743]. Human-Computer [743]. Human-Robot [59]. Human-Robot-Interaction-Technologies [50]. Humanoid [213]. Humans [15, 190]. Hybrid [602, 33, 491, 382, 345, 109, 437, 162, 313, 327]. Hyper [140]. Hyper-Parameters [140]. Hyperbolic [69]. Hyperelliptic [499]. Hypergraph [355]. Hypermedia [636]. Hypotheses [602, 397]. Hypothesis [541, 244].

I-alpha [442]. ICEC [727]. ICIRA [714]. ICMP [159]. ICMP-Based [159]. ICOCOON [329]. ICTAC [726]. ICU [677]. IDEA [700]. Identification [672, 322, 292, 211]. Identify [158, 675]. Identifying [325, 139]. Identity [236]. Idiopathic [124]. IFIP [719]. II [717, 729, 714]. III [718]. Illumina [686]. Image [716, 717, 718, 107, 222, 59, 220, 138, 226, 225, 202, 442, 715, 77, 92]. Image- [107]. Images [135, 98, 227, 224, 216, 74, 217, 110, 229, 448, 117, 129, 143]. Imaging [127, 85, 109, 100, 101, 132]. Imminent [577, 67]. Impact [506, 209]. Impairment [67]. Imperfect [368]. Implantation [108]. Implementability [596]. Implementation [507, 459, 219, 513, 48]. Implementing [193, 400]. Implications [14, 397, 484]. Implicit [362]. Importance [174]. Important [600]. Importing [583]. Impossible [696]. Impression [13]. Impressions [7]. Improve [40, 665]. Improved [692, 509, 55, 697, 115, 531, 447, 442, 396]. Improvement [159, 441]. Improves [526, 79]. Improving [602, 222, 178, 398, 226]. in-Browser [241]. In-Game [422]. in-the-Loop [175]. Inapproximability [542]. Incentive [574]. Incidence [437]. Including [528]. Incomplete [264, 381]. Inconsistency [259]. Incorporating [133, 122]. Incremental [591, 25, 97, 364, 18]. Incrementality [595]. Indefinite [263]. Independent [596, 560]. Index [24, 469, 53]. Indexes [25, 533]. India [726]. Indifferentiable [498]. Individual [469]. Induced [540]. Infectious [670]. Infective [514]. Inference [298, 289, 70, 526]. Inferring [652]. Information [562, 459, 468, 688, 264, 32, 523, 466, 670, 381, 638, 654, 662, 214, 465, 736, 713, 138, 590, 442, 377, 248, 318]. Information-Gain [214]. Information-Gathering [318]. Infrared [447]. Infrastructure [509].

Infrastructures [653]. Initialising [146]. Inner [504]. Input [177]. Insider [158]. Inspired [222]. Instance [550]. Instruction [1]. Instructional [194]. Instructions [40]. Integrated [688, 649, 245]. Integrating [670]. Integration [582, 678, 32, 635, 38, 115]. Integrity [358]. Intelligence [745, 730, 733, 245, 724]. Intelligent [722, 658, 44, 712, 714]. Intelligibility [20]. Intensive [420, 243]. Intention [62]. Inter [599, 139]. Inter-Point [139]. Inter-task [599]. Interaction [678, 743, 200, 190, 323, 50, 679, 474, 521, 316]. Interactions [658, 614, 528]. Interactive [175, 744, 402, 384, 328, 424, 19, 629, 324, 206, 316]. Interconnection [151]. Interdisciplinary [418]. Interests [664]. Interface [635, 663, 62]. Interfaces [744, 564, 209]. Interference [176, 58]. Interlocking [87]. Internal [614]. Internalization [632]. International [735, 731, 716, 717, 718, 745, 722, 746, 744, 743, 720, 243, 721, 727, 736, 725, 734, 733, 732, 742, 712, 719, 737, 418, 726, 232, 714, 738, 715, 237]. Interpersonal [7]. Interpretability [284]. Interpretation [104]. Intersection [524]. Intervals [291]. Intervention [716, 717, 718]. Interventions [90]. Intra [132, 92]. Intra-image [92]. Intra-operative [132]. Intramedullary [87]. Introducing [407]. Introduction [208]. Invasive [684, 431]. Inventive [38]. Inverse [367]. Inversion [517]. Investigators [410]. Investment [469]. Involved [322, 51]. IPEC [738]. Iris [212]. Isoforms [530]. Issues [284]. Italy [735, 731, 745, 723, 720, 650]. Itemsets [343]. Iteration [371]. ITrace [159]. IV [512]. IVA [712].

Jacobian [500]. Java [300, 195]. Java-Based [195]. Javascript [162]. JELIA [724]. Join [25, 31]. Joins [22]. Joint [113]. Judgments [7]. July [745]. June [745]. Just [37]. Just-In-Time [37]. Justifications [623].

KALCAS [30]. Kansei [438]. Kernel [680, 97]. Kernel-Based [680]. Kernelization [545]. Kernels [387]. Key [449, 509, 512, 461, 460]. Key-IV [512]. Keyword [478, 491, 490]. KI [730]. Kidneys [135]. Knowledge [256, 678, 420, 728, 729, 197, 508, 397, 561, 150, 451]. Knowledge-Base [397]. Knowledge-Based [150]. Knowledge-Guided [451]. Knowledge-Intensive [420]. Koblitz [507]. Korea [719].

Label [351, 363, 73, 75, 71]. Label-Noise [351]. labeled [523]. Labeling [566, 145, 617, 82]. Laboratories [194]. Landmark [119]. Landmark-Local [119]. Lands [321]. Language [28, 491, 564, 405]. Laparoscopic [96]. Lapin [710]. Large [485, 25, 592, 578, 349, 369, 65, 216, 152, 139, 480, 451, 740]. Large-Margin [349]. Large-Scale [592, 451, 740]. Laser [57]. Latent [104]. Lateral [69]. Latin [736]. LATINCRYPT [736]. Lattice [504, 620, 505, 110, 453]. Lattice-Based [504, 505]. Lattice-Boltzmann [110]. Layer [29, 22]. Lazy [302]. LBM [110]. LBM-EP [110]. Leaf [687]. Leakage [502]. Leakage-Resilient [502]. Learned [187, 371, 393]. Learning [337, 365, 599, 672, 193, 372, 430, 95, 660, 338, 663, 369, 339, 353, 370, 655, 382, 636, 728, 729, 201, 345, 227, 348, 433, 376, 385, 608, 398, 417, 422, 128, 367, 118, 432, 196, 679, 391, 629, 377, 108, 356, 92, 64, 656, 143, 102, 635, 721]. Learning-Based [365]. Learnt [194, 424]. LED [54]. Left [134, 106]. Lengths [66]. Lesion [98, 118]. Lesions [76]. Lessons [187, 424, 194]. Level [230, 234, 148, 436]. Level-Set [230]. Leveraging [22]. Library [506]. Lifecycle [404]. Light [211]. Like [250, 12]. Limb [51, 52, 48]. Limitation [473]. Line [187, 662, 636, 702, 586, 522, 406]. Linear [674, 693, 299, 695, 313, 691, 353]. Link [521]. Linkage [56]. Linked [569, 570, 573, 560]. Linking [403]. Lip [20]. Lip-Reading [20]. Live [361]. Liver [91]. Ljubljana [737, 738]. Load [485, 156]. Lobe [115]. Local [601, 634, 632, 119, 129, 102]. Localisation [186]. Locality [518, 151]. Locality-Sensitive [518]. Locally [302]. Locate [59]. Location [331]. Location-Aware [331]. Log [113]. Log-Demons [113]. Logic [253, 279, 278, 261, 262, 264, 280, 267, 269, 272, 41, 347]. Logic-Based [279, 278, 41]. Logical [428]. Logics [254, 264, 297, 724]. Logistic [351]. Longitudinal [9]. Look [432]. Loop [175, 19]. Loops [299]. Low [203]. Low-Cost [203]. Lower [51, 46, 531]. LPN [710]. LSCITS [581]. LTE [461]. LTL [296]. Lung [120, 115].

Maastricht [741]. Machine [337, 672, 338, 339, 353, 382, 612, 679, 728, 729]. Machine-Learning [382, 679]. Machine-to-Machine [612]. Machines [683, 165]. MACs [703]. Magnetic [676]. Magneto [58]. Magneto-Rheological [58]. Magnitude [272]. mail [657, 645]. Maintenance [170, 635]. Make [13]. Making [366, 439, 6]. Malicious [162]. Malware [456, 475]. Management [259, 650, 740, 237, 725, 687, 470, 471, 236]. Managing [276]. Mandibular [93]. Manifold [143]. Manipulator [53]. Manipulators [436]. Manufacturing [430]. Many [72, 148]. Many-Task [148]. Map [82, 47]. Map-Based [47]. Mapping [599, 492, 233]. Mappings [511]. Maps [66]. Marburg [725]. March [740, 746]. Margin [349]. Markets [578, 366]. Markov [120, 62, 347, 364]. MAS [603]. MASKETEER [305]. Masks [15]. Massively [427]. Master [43]. Master-Slave [43]. Mastering [192]. Matching [287, 118, 664]. Matchings [554]. Material [328]. Materialization [563]. Mathematical [734]. Matrices [225]. Matter [83, 314, 412, 515, 575, 597, 615, 630, 646, 689, 21, 42, 558, 576, 598, 616, 631, 647, 668, 690, 711, 63, 84, 105, 126, 147, 168, 189, 210, 231, 252, 273, 294, 315, 336, 357, 378, 399, 413, 434, 455, 476, 497, 516, 537, 75, 142, 131, 76]. Maude [301]. MAX [543]. MAX-SAT [543]. Maximum [267, 540]. May [742]. MBIA [715]. MCI [95, 78, 101]. McOE [702]. MDE [405]. MDP [369]. Me [12]. Means [256]. Measurement [57, 211]. Measures [571, 493, 519]. Mechanical [111]. Mechanics [319, 196]. Mechanism [574, 58]. Mechanisms [465]. Media [479, 350, 293]. Mediated [639]. Medical [442, 716, 717, 718]. Medicine [203]. Medulloblastoma [104]. Meet [706].

Meet-In-The-Middle [706]. Meeting [745, 329]. Meets [170, 670, 545, 324]. Mega [477]. Mega-modeling [477]. Memetic [656]. Message [501, 167]. Message-Based [501]. Metabolic [98]. Metaheuristic [188]. Metaprogramming [400]. MetaVals [418]. Method [494, 33, 490, 350, 110, 447]. Methodology [430, 609, 695, 424]. Methods [745, 339, 35, 361, 734, 70, 354, 679]. Methylation [686]. Metric [253, 138, 356]. Metro [223]. MIB [159]. MIB-ITrace-CP [159]. MICCAI [716, 717, 718, 715]. Micro [100, 218]. Micro-architecture [100]. Micro-graphs [218]. Microarray [671, 674, 686]. Microblog [241]. Microcontroller [513]. Middle [706]. Mild [67]. Mind [19]. Mind-Body [19]. Minimal [270]. Minimization [173]. Minimum [141]. Mining [374, 338, 373, 375, 438, 358, 453]. Minor [548, 545]. miRNA [674]. Mirror [427]. Misregistration [14]. Mitigation [163]. Mixed [566]. MJPEG [223]. MMM [734]. MMM-ACNS [734]. Mobile [565, 201, 438, 45, 457]. Modal [303, 262, 350, 114]. modality [442]. Mode [440, 48]. Model [380, 572, 660, 667, 479, 120, 584, 655, 411, 270, 493, 134, 73, 179, 437, 375, 495, 190, 332, 164, 244, 371, 183, 471, 113, 246, 694, 308, 309, 474, 102, 106]. Model-Based [411, 179, 471]. Model-Driven [572]. Modeling [235, 735, 723, 603, 338, 472, 98, 243, 404, 15, 17, 245, 128, 251, 91, 393, 248, 58, 309, 587, 477]. Modelling [408, 588, 669, 589, 403, 232, 482]. Models [253, 671, 674, 559, 233, 32, 298, 489, 481, 133, 385, 38, 734, 62, 181, 610, 677, 247, 391, 341, 364, 146, 521]. Modes [700]. MoDIC [723]. Modified [217]. Modular [591]. modularization [175]. Molecular [521]. Molina [640]. Monadic [280]. Monitoring [409, 296, 45]. Monotone [553]. Monte [354, 368]. Monterey [740]. Montreal [714]. MORE-BI [723, 249]. Morphology [93]. Morphometry [69]. MOSKitt4ME [494]. Most [256]. Motif [685]. Motion [223, 219, 46, 203, 393, 108, 436]. Motivation [659, 423]. Movements [52]. Moving [386]. MPI [167]. MR [117, 143, 102]. MRF [118]. MRI [81, 65, 106, 70, 128, 140, 76]. MRI-TRUS [140]. MS [76]. MSC [295]. MSOL [556]. MSP430X [513]. Multi [671, 319, 565, 25, 566, 574, 363, 169, 523, 655, 98, 588, 608, 609, 106, 100, 350, 620, 742, 251, 401, 612, 131, 531, 114, 442, 86, 71, 60, 88, 625, 587, 741]. Multi-Agent [742, 655, 608, 609, 612, 741]. Multi-agents [620]. Multi-Atlas [71]. Multi-class [363, 98]. Multi-criteria [625]. Multi-Detector [100]. Multi-dimensional [566, 251]. Multi-disciplinary [588]. Multi-DOF [60]. Multi-domain [565]. Multi-genre [319]. Multi-labeled [523]. Multi-modal [350, 114]. Multi-modality [442]. Multi-model-Based [106]. Multi-objective [169]. Multi-organ [86]. Multi-piercing [88]. Multi-robotic [401]. Multi-sample [671]. Multi-scale [131]. Multi-state [531]. Multi-table [25]. Multi-view [587]. Multiagent [600, 660, 611]. Multiatlas [72]. Multibit [532]. Multicenter [65]. Multicuts [549]. Multidimensional [245]. Multimodal [384, 73, 64, 18, 715]. Multiparty [622]. Multiplayer [207, 192]. Multiple [409, 349, 75, 499, 146]. Multiplication [507]. Multipoint [329].

Multiprocessor [151]. Multiscale [134, 122]. Multivariate [508]. Musculoskeletal [50]. Music [339]. Mutations [675]. Mutual [138]. Myths [632].

Naehrig [498]. Nails [87]. Naive [277]. Nameless [302]. Narrative [642]. Narrowing [304]. Nations [634]. Natural [634, 564, 62, 302]. Nature [639]. Navigating [284]. Navigation [251]. NCCPIS [157]. Nearest [35, 387]. Necessities [277]. Need [70]. Negligibility [272]. Negotiation [6]. Neighbor [35]. Neighborhoods [356]. Neighbors [387]. Neighbourhood [136]. Nested [254]. Netherlands [741]. Network [158, 440, 468, 159, 345, 35, 681, 78, 734, 445, 67, 341, 151, 346, 719]. Network-Based [78]. Networked [157]. Networks [172, 379, 583, 571, 66, 669, 289, 347, 154, 447, 452, 155, 521]. Neural [669, 447]. Neurogenetic [669]. Neuroimaging [79, 64]. News [452]. Next [659, 439]. Nice [716, 717, 718, 715]. NoCoDA [723, 237]. Nogoods [394]. Noise [351, 385, 58]. Noisy [217, 460]. Non [467, 374, 299, 353, 237, 62, 266, 431, 146]. Non-disclosure [467]. Non-invasive [431]. Non-linear [353]. Non-optimal [266]. Non-rigid [146]. Non-Spurious [374]. Non-termination [299]. Non-verbal [62]. Nonblocker [545]. Nonlinear [292, 653, 48]. Nonmonotonic [270]. Nonparametric [367]. Nonrigid [138]. Nonuniformity [447]. Nonverbal [13]. Norm [10]. Norm-Related [10]. Normalcy [472]. Normalization [74]. Note [349]. Notes [321]. Novel [479, 120, 88]. November [741]. NPC [719]. NXT [195].

Obfuscated [162]. Object [214, 298]. objective [169]. Observations [696]. Obtained [218]. Occupational [334]. occurrence [225]. October [735, 716, 717, 718, 723, 736, 734, 733, 714, 715]. Odds [581]. ODF [128]. Off [636, 422]. Off-Game [422]. Off-Line [636]. OLAP [250]. OLAP-Like [250]. Older [335]. OLED [55]. On-Line [662, 702]. One [262]. One-Agent [262]. Ones [154]. Online [601, 398, 108]. ontinuo [24]. Ontological [233, 247, 484]. Ontologies [483, 268]. Ontology [33, 652]. Open [170]. OpenStreetMap [643]. Operation [740]. Operations [153]. operative [132]. Operators [188]. Opponent [325]. Opportunities [432]. Optic [92]. Optical [221]. Optimal [141, 525, 501, 140, 182, 266]. Optimality [619]. Optimised [177]. Optimization [601, 81, 187, 169, 370, 216, 437, 118, 185, 132, 460, 53]. Optimized [116]. Optimizing [176]. Orchestration [467]. Order [280, 268, 272, 118]. orders [290]. Ordinal [275]. organ [86]. Organizations [603]. Oriented [593, 33, 298, 408, 586, 637]. Orthopaedic [116]. OSEK [309]. OSEK/VDX [309]. Other [614, 703]. Otsu [55]. Our [639]. Outcome [677]. Outcomes [279, 622, 196]. Outer [88]. Output [369, 514, 709]. Overall [402]. OWFs [510]. OWL [605]. Oxford [740].

P2P [286, 156]. PAC [385]. PAC-Learning [385]. Packet [153]. Paint [326]. Pairs [549, 1]. Palpation [90]. Panel [684, 55]. Panic [414]. Papers [745, 740, 746, 744, 743, 741, 742]. Parallel [166, 164, 719, 53]. Parameter [542, 133, 111]. Parameterized [551, 543, 544, 541, 557, 546, 539, 528, 738]. Parameters [234, 140, 346]. Parametric [267, 128, 653]. Parametrizable [319]. Pareto [182, 619]. Pareto-optimal [182]. Parity [310]. Parochialism [659]. Part [716, 717, 718, 728, 729, 714]. Partial [381, 290]. Participative [420]. Participatory [427]. Particle [116]. Partners [5]. Parts [146]. Passing [167]. Passive [48]. Patch [71]. Patch-Based [71]. Path [107, 538, 66]. Pathological [60]. Pathway [579]. Pathways [672, 673]. Pathwidth [555]. Patients [677]. Pattern [24, 176, 221, 375]. Patterns [400, 666, 567, 452]. People [384, 6]. Perceived [13, 199]. Performance [327, 171, 343, 665, 64, 53]. Performances [219, 435]. Perfusivity [85]. Person [435]. Person-Computer [435]. Personalities [15]. Personality [7, 15]. Personalization [418]. Personalized [361]. Persons [386]. Perspective [578]. Perspectives [682]. Persuasion [622, 625]. Pervasive [328]. PET [98]. Petersburg [734]. Phase [56]. Philosopher [324]. Philosophy [324]. Phones [438]. Phosphor [54]. Photography [638]. Phylogenetic [518, 680, 519]. Phylogeny [517]. Physical [577, 157, 334]. Physiological [682]. Physiology [423]. Pictures [327]. Piece [327]. piercing [88]. Piezoelectric [54]. Pilot [209]. Pipeline [672]. Pisa [731]. PKDD [728, 729]. Plan [626, 392]. Planar [556, 549]. Plane [447]. Planning [260, 96, 214, 386, 627, 266, 364, 436]. Plates [93]. Platform [331, 165, 204]. Platforms [674, 245]. Play [428]. Player [325, 332]. Playgrounds [323]. Playing [639, 330]. Plots [318]. Pneumoperitoneum [96]. Point [250, 139]. Point-Based [250]. Poland [713]. Policies [467, 466, 370]. Policy [458, 371]. Polyaffine [122]. Polygonal [107]. Polynomial [551, 549, 550, 542, 703]. Polynomial-Time [549]. Polynomials [508]. Polytomies [525]. POMDP [611]. POMDP-Based [611]. Popular [519]. Popularity [156]. Popularity-Aware [156]. Population [93, 522, 75, 530]. Population-Based [93]. Portable [446]. Possibilistic [271]. Possibility [291]. Postulational [290]. Potential [444]. Potentials [203]. Power [54]. Poznań [713]. Practical [619, 699, 484]. Pragmatics [587]. Pre [290]. Pre-orders [290]. Precedence [14]. Precision [346]. Predict [679, 283]. Prediction [95, 688, 686, 684, 676, 291, 386, 97, 677, 665, 359]. Predictive [374]. Predictors [64]. Predicts [67]. Preface [249]. Preference [365]. Preferences [325, 269, 277, 12, 624]. Preferential [253]. Preferred [279]. Preimage [706, 698]. Preimages [705, 511]. Prejudice [360]. Premises [283]. Preprocessing [548]. Preschoolers [4]. Prescription [44]. Prescription-Diagnosis [44]. Presence [686, 209]. PRESENT [694]. Preservation [220]. Preserving [23, 517, 38, 161]. Primary [433]. Prime [500]. Primitives [393]. Printing [54]. Priors [92]. Privacy [23, 305, 161]. Private [359]. ProB [308]. Probabilistic [592, 267, 408, 73, 669, 41, 341].

Probabilities [347]. Probability [363]. Problem [173, 263, 541, 188, 39, 450]. Problem-Specific [188]. Problems [601, 170, 353, 557, 548, 524, 340, 653, 312]. Procedure [431, 675]. Procedures [109]. Proceedings [716, 717, 718, 728, 729, 714, 735, 731, 722, 723, 720, 730, 721, 727, 736, 725, 734, 733, 732, 713, 712, 719, 737, 726, 738, 715, 724]. Process [235, 303, 494, 603, 57, 493, 404, 613, 321]. ProCesses [30, 322, 185, 364]. Processing [26, 37]. Processors [40]. Procrustes [139]. Product [504, 187, 586, 406]. Production [50]. Profiles [660, 472]. Profiling [129]. Prognosis [282]. Program [265, 392]. Programming [667, 584]. Programs [269, 180, 167]. Progress [736]. Progression [684]. Project [652]. Projected [20, 359]. Projection [202]. Promotes [655]. Promotion [634]. Proof [546]. Properties [592, 36]. Proposal [613]. Propositional [261]. Propositions [425]. Prostate [227, 140]. Prostatic [99]. Protection [414, 472, 514]. Protein [680, 679]. Proteins [679]. Proteomic [678]. Protocol [449, 710, 155]. Protocols [503, 161, 687, 461, 308, 460]. Proton [676]. Provable [691]. Provenance [240, 246]. Provide [48]. Providing [494]. Proximity [571]. Pseudo [388, 706, 698]. Pseudo-Boolean [388]. Pseudoknots [528]. Pseudorandom [503]. PTL [261]. Public [645]. Pulmonary [115]. Puppetry [327]. Putting [175]. Puzzle [207, 204]. Puzzle-it [204]. PVS [410].

Quadratic [509]. Qualitative [272, 458, 394]. Quality [93, 638, 220]. Quantification [530, 293]. Quantitative [373, 100, 458, 453]. Quantized [313]. Quartets [531]. Queries [24, 25, 568, 28, 31, 36]. Query [28, 285, 268, 287, 37, 275, 396]. Querying [478, 274]. QueryPOMDP [611]. Quest [595]. Question [1, 4]. Question-Answer [1]. Question/Answering [4].

r [24]. RABAC [463]. Radiation [87]. Radiation-Free [87]. Radiometric [80]. Rail [225]. Railway [223, 404]. Random [674, 135, 179, 511]. Randomised [182]. Randomized [539]. Randomness [174]. Ranking [363, 241]. Ranks [275]. Rare [670]. Rate [501]. Ray [107, 109]. RBAC [464]. RDF [247]. Re [175]. Re-modularization [175]. Reachability [311]. Reading [20]. Reads [149]. Real [509, 46, 109, 179, 216, 190, 91, 641]. Real-Time [46, 109, 179, 91]. Realistic [177, 98]. Realities [632]. Reality [191, 661, 331]. Realizer [18]. Really [70, 651]. Realtime [435]. Reason [628]. Reasoning [602, 267, 270, 272, 394, 624]. Rebound [697]. Recognition [137, 45, 392]. Recognizing [614]. Recommendation [339, 362, 361]. Recommendations [486, 242]. Reconciliation [559, 525, 526]. Reconfigurable [505]. Reconstructing [521]. Reconstruction [517, 518, 520, 103, 214, 522, 74, 640]. Recovered [185]. Rectal [227]. Rectangle [536]. Rectangular [229]. Recursion [39]. Recursivity [649]. Reduced [693, 228, 696, 704, 346, 698].

Reduced-Round [696]. Reducing [524]. Reduction [144, 116, 185]. Refactoring [178]. Reference [153]. Refinement [262]. Reformulation [268]. Regional [139, 143]. Registration [144, 120, 116, 141, 123, 142, 140, 138, 115, 122, 113, 114, 442, 117, 143, 77, 146, 124, 125]. Regression [351, 683, 97]. Regularity [510]. Regularization [123, 121]. Regularizer [360]. Regulations [259]. Rehabilitation [51, 46, 43, 52, 44, 203, 48]. Reinforcement [365, 599, 369, 608, 367]. Related [10, 335, 512]. Relation [258]. Relational [29, 267, 383]. Relationship [639, 438]. Relationships [619]. Relaxation [81]. Relaxed [31]. RELAXing [183]. Reliability [674, 470]. Reliable [27, 153, 584, 85]. Remainder [166]. Remote [90, 160]. Remover [360]. Reorientation [123]. Replacement [23]. Repositories [493]. Representation [535, 221, 134, 289, 302, 532]. Representing [291]. Republic [722]. Requirements [470, 590, 248]. Research [54, 438, 60]. ReServE [27]. Resilient [502, 731, 405]. Resolution [193, 287, 129]. Resonance [676]. Resource [235, 665]. Resourcing [626]. Response [469]. Resting [68]. Restoration [217]. Restoring [286]. Restricted [556, 544, 524, 511]. Results [259, 542, 493, 567]. Retention [23]. Retinal [92]. Return [469]. Reusable [570, 400]. Revealing [589]. Revenue [574]. Reverification [592]. Reverse [181]. Reviews [642]. Revised [745, 740, 746, 744, 743, 741, 742]. Revision [290, 397]. Revisited [708]. Rewrite [303]. RFID [449, 34]. Rheological [58]. Ricci [69, 124]. Ridge [97]. Rigid [88, 125, 146]. Rigid-Flexible [88]. RIGiM [723]. Ring [710]. Ring-LPN [710]. RIPEMD [704]. RIPEMD-128 [704]. Risks [581, 457]. Riva [720]. RNA [527, 528]. Road [35]. Road-Network [35]. Robot [51, 46, 59, 62, 52, 45, 44, 48, 50, 49, 436, 60]. Robotic [90, 401]. Robotics [337, 714]. Robots [195, 213]. Robust [172, 351, 215, 353, 70, 82, 225, 160, 102]. Rocchio [350]. ROIs [77]. Role [655, 463]. Role-Centric [463]. Roles [496]. Room [329]. Rough [454]. Round [696, 698]. Round-Reduced [698]. Rounds [514]. Routing [445, 155]. Rule [233, 373, 652, 453]. Rules [374, 343, 482]. Russia [734].

S [304]. S-Narrowing [304]. Saarbrücken [730]. Safely [59]. Safety [592, 404, 403]. sample [671]. Sampling [221, 141, 343, 60]. Santa [712]. Santiago [736]. SAT [173, 543, 388]. Satisfiability [39, 266]. SBSE [170, 171]. Scalability [208]. Scalable [584, 248, 725]. Scalar [507]. Scale [592, 682, 578, 100, 456, 451, 131, 740]. Scaling [276]. Scans [106]. Scenario [430]. Scenarios [609]. Schedule [176]. Scheduling [665, 150, 148]. Schema [485, 240]. Schemas [480]. Scheme [145, 462, 323, 160, 436]. Schemes [702]. School [433]. Schooling [194]. Schools [194]. Science [202]. Scientific [244, 651]. Scoliosis [124]. Score [348]. Score-Based [348]. Scoring [374, 152]. Scripted [1]. Scripting [429]. Scrum [661]. SCS [484]. Search [565, 174, 572, 574, 187, 491, 188, 178, 493, 35, 179, 490, 354, 567, 368, 184, 396, 739, 720]. Search-Based [174, 178, 179]. Searches [532]. Searching [381, 182]. SeCoGIS [723]. Second [13, 744, 743, 280, 715].

Secondo [435]. Sector [645]. Secure [503, 709]. Secured [467]. Security [600, 506, 650, 583, 662, 178, 465, 736, 734, 163, 458, 691, 470, 701, 471, 457, 700]. Segmentation [72, 107, 135, 227, 134, 133, 230, 106, 228, 99, 226, 115, 79, 86, 71]. Segura [640]. Selected [745, 740, 746, 744, 743, 741, 742]. Selection [25, 26, 31, 676, 469, 36, 78, 140, 355, 625]. Selective [269]. Self [682, 606, 138, 150, 436]. Self-adaptive [606]. Self-Motion [436]. Self-Scheduling [150]. Self-similarity [682, 138]. Semantic [483, 492, 570, 667, 104, 491, 560, 47]. Semantics [253, 279, 278, 255, 380, 429, 383, 605, 269, 270, 607, 38, 619, 277, 301, 302, 621]. sEMG [43, 49]. sEMG-Based [49]. Semi [30, 376, 499, 377, 488, 355]. Semi-automatic [30]. Semi-bent [499]. Semi-structured [488]. Semi-supervised [376, 377, 355]. Sensing [127]. Sensitive [518, 208]. Sensor [158, 233, 215, 445]. Sentiment [352]. Separation [444]. September [731, 722, 743, 728, 729, 720, 730, 721, 727, 725, 732, 713, 712, 719, 737, 726, 738, 724]. Septic [677]. Seq [685]. Sequence [352, 533, 528]. Sequences [680, 666, 685]. Sequencing [675]. Sequent [254]. Sequential [250]. Sequents [257]. SERENE [731]. Series [529, 682, 451]. Serious [419, 426, 430, 420, 422, 205, 416, 197, 432, 204, 418, 425, 321, 192, 421, 732]. Serpent [694]. Service [235, 409, 27, 408, 407, 358, 234]. Service-Oriented [408]. Services [467, 177, 564, 407, 232, 637]. Set [547, 230, 707, 542]. Sets [299, 369, 265, 553, 512, 377, 454]. SGDA [732]. SHA [705, 706]. SHA-2 [705, 706]. Shafer [158]. Shape [112, 130]. Shape-Based [130]. Shapes [139]. Shared [47]. Sharing [574, 459]. Sheath [88]. Short [639, 152, 709]. Short-Output [709]. Shortgun [149]. SibJoin [520]. Sibling [520]. Signature [112, 462]. Significance [452]. Significance-Driven [452]. SIMD [40]. Similarity [253, 28, 493, 41, 22, 682, 138]. Simple [299, 681, 386]. SimpleDB [29]. SimpleSQL [29]. Simulation [96, 51, 191, 116, 588, 653, 91, 110, 61, 157]. Simulation-Based [116]. Simultaneous [228, 122]. Sink [549]. Sites [489, 679]. Situated [191]. Size [72, 522]. Skein [705]. Skein-512 [705]. Skeleton [114]. Skill [348]. Skills [414]. Skin [118]. Slave [43]. Sliced [484]. Slides [104]. Sliding [48]. Slovenia [737, 738]. Small [548, 389, 164]. Smart [215, 443, 328, 366, 47, 160]. Smartphones [651]. SmartTestGen [307]. Smoothing [344]. SNI [442]. Social [5, 583, 571, 479, 200, 350, 323, 341, 316]. Society [641]. Soft [414, 97]. Software [174, 175, 746, 187, 658, 170, 188, 632, 659, 720, 178, 586, 666, 664, 661, 656, 406, 731]. Soil [416]. Solids [653]. Solution [222]. Solvers [173]. Solving [340]. Somatic [675]. Some [546]. Sosemanuk [512]. Source [549]. Source-Sink [549]. Sources [409, 563, 561, 444]. Space [582, 535, 557, 428]. Space-Efficient [535]. Spanning [141]. SPARQL [568]. Sparse [599, 134, 552, 390, 393, 64, 71]. Spatial [502, 478, 14, 74, 209, 80]. Spatio [386]. Spatiotemporal [103]. Speaker [17]. Special [707]. Specific [492, 188, 589, 86]. Specifications [380]. Specified [301]. Spectra [186, 355]. Spectra-Based [186]. Spectral [112, 119]. Spectroscopy [676]. Spectrum [127, 682]. Speed [513, 218]. Spelling [490]. Spiking [669]. Spin [82]. Spine [102]. spline [121]. Sports [721, 203]. Spotting [325]. Spray [155]. Spurious [374]. SQL [285]. Squaring [429]. SSBSE [720]. SSMTA* [35]. St [734]. Stable [278, 277]. Stacked [581]. Stakeholder [664]. State [379, 313, 643, 68, 531]. States [381, 614]. Static [25, 239]. Statins [677]. Statistical [98]. Statistics [122]. Status [642]. Stealing [701]. Steering [594]. Stitching [218]. Stochastic [398, 293, 395, 474]. Storage [22]. Store [34, 484]. Stories [321]. Story [424, 318]. StoryTec [205]. Storytelling [206, 316]. Strategies [626, 604, 179]. Strategy [51, 391, 56]. Stream [704, 283]. Streams [31, 274]. Strength [162]. Strengthen [449]. Strengths [573]. String [40]. Structural [335, 78, 307, 101]. Structure [107, 345, 99, 528, 94, 694, 92]. Structure-Sequence [528]. Structured [565, 372, 566, 211, 156, 561, 64, 488]. Structures [532]. Student [199]. Students [660, 661]. Studies [194, 423, 653]. Study [51, 662, 224, 438, 181, 290, 209, 410, 401, 677, 8, 450]. Studying [69]. Stuttering [310]. Subfield [133]. Subgraph [548, 540]. Subject [86]. Subject-Specific [86]. Subjective [347]. Subsequence [24]. Subset [547]. Succinct [534, 532]. Suggestion [490, 396]. Suggestions [546, 224]. Suite [173, 307]. SUM [725, 620]. Summation [553]. Superior [154]. Superstring [544]. Supervised [352, 227, 376, 377, 355]. Support [568, 678, 494, 570, 660, 329, 683, 22, 661]. Supported [655]. Supporting [240, 410, 687]. Surface [225]. Surfaces [124]. Surgery [96, 88]. Surgical [89]. Surveillance [223, 215]. Survey [571]. Survival [683]. Susceptibility [76]. Sustainability [574]. Sustainable [430]. SWAD [215]. SWAD-Based [215]. Swap [544]. Symbolically [295]. Symmetric [117]. Symmetries [593]. Symposium [720, 738]. Synchronization [449]. Syncretic [620]. Synergy [649]. Synthesis [582, 38]. System [194, 459, 649, 274, 411, 654, 55, 57, 384, 109, 224, 162, 326, 211, 152, 443, 157, 43, 423, 435, 44, 401, 203, 202, 446, 56]. Systems [172, 485, 279, 600, 731, 303, 722, 577, 592, 740, 282, 468, 578, 593, 741, 221, 329, 635, 594, 636, 606, 404, 179, 208, 292, 580, 358, 284, 313, 742, 611, 713, 612, 304, 665, 629, 156, 590, 283, 474]. Systems- [404]. Systems-of-Systems [172].

T1 [113]. table [25]. Tables [560, 487, 562]. Tagging [224]. Taipei [742]. Taiwan [742]. Take [72]. Taken [538]. Talk [389]. Talking [481]. Talks [320]. Taming [582]. Tangible [334]. Target [144]. Targeted [472]. Task [604, 148, 68, 599]. Taxonomy [200]. Teaching [3, 661, 321, 195]. Team [655, 656]. Team-Role [655]. Teams [580, 652]. Technique [374]. Techniques [602, 539, 186]. Technologies [744, 654, 50]. Technology [443, 202]. Teleconference [329]. Telerobotic [191]. Telic [238]. Telic/Atelic [238]. Temperature [57]. Templates [72, 207]. Temporal [112, 682, 271, 386, 266, 297, 392, 238]. Tensor [362, 123, 100, 74]. termination [299]. Terms [499]. Tertiary [528]. Test [173, 177, 152, 307, 301]. Testing [180, 179, 163, 184, 406]. Text [533, 341]. Texture [448]. Thalamus [19]. Their [680, 709]. Thematic [569]. Theorem [166]. Theoretical [726]. Theory [158, 600]. Therapeutic [384, 332]. Therapy [332, 48, 334]. Thinking [428]. Third [732, 232]. Thoracic [94]. Threads [176]. Three [264, 466, 109]. Three-Dimensional [109]. Three-Valued [264, 466]. Threshold [426]. Tight [343]. Tile [342]. Time [551, 295, 671, 250, 549, 13, 529, 281, 81, 682, 541, 46, 408, 109, 179, 37, 555, 292, 400, 62, 313, 91, 451]. Time-Constrained [295]. Time-Course [671]. Time-Frequency [46]. Time-Probabilistic [408]. Time-Variant [292]. Time-Varying [62]. ting [358]. Tissue [97, 91]. ToF [214]. Tolerance [291]. Tolerant [255, 402, 401, 155]. Tomographic [132]. Tool [221, 33, 329, 320, 326, 157, 687, 306]. Tools [585]. Top [568, 41]. Top- [41]. Top-K [568]. Total [98]. Toulouse [724]. Tourism [743, 632]. Trabecular [100]. Trace [499]. Traceability [34]. Traceback [159]. Tracing [501]. Tracked [34]. Tracker [215]. Tracking [223, 108]. Tract [75, 131]. Tractable [287]. Tractography [142]. Trading [510, 439]. Traffic [259]. Training [419, 191, 721, 384, 320, 190, 416, 44, 347]. Traitor [501]. Trajectories [393]. Trajectory [386]. Trans [227, 331]. Trans-reality [331]. Transactions [276, 283]. Transcatheter [108]. Transfer [599, 95]. Transform [485, 111]. Transformation [451]. Transient [57]. Transition [371, 364]. Translate [297]. Translation [382, 165]. Translational [53]. Transport [293]. Trauma [116]. Treated [677]. Tree [518, 680, 524, 141, 354, 527, 526, 368, 528, 532, 519]. Tree-Based [354]. Trees [523, 145, 684, 525, 375, 526, 532, 342]. Triangular [253]. Triangulations [524]. Trichotomic [636]. Triclusters [529]. TRIMAX [529]. Trimming [429]. Triple [34]. Triples [573]. Triplet [151]. Triplet-Based [151]. Triplets [531]. TRUS [140]. Trust [333, 464, 628]. Trust-Aware [464]. Tumor [676]. Turning [194]. Tutorial [171]. Two [686, 670, 17, 56]. Two-Color [686]. Two-Phase [56]. Type [635, 298, 445]. Types [676, 416]. Typicality [261].

UK [740, 728, 729]. UKCF [165]. Ultra [578]. Ultra-Large-Scale [578]. Ultrasound [227, 116, 137, 90]. Ultrasound-CT [116]. UMTS [461]. UNAF [707]. Unbiased [142]. Uncertain [274, 364]. Uncertainties [526]. Uncertainty [582, 276, 291, 725, 133, 118, 293, 183, 140]. Uncertainty-Based [118]. Undecidability [313]. Underdetermined [444]. Undergoing [334]. Understandability [481]. Understanding [481, 480, 487]. Unfounded [265]. Unified [190]. Uniform [121]. Uniformity [267]. Units [26]. Universal [709]. Unordered [527]. Unpartitioned [393]. Unrooted [524, 527]. Unscented [111]. Unsupervised [387]. UOWHFs [510]. Update [263]. Upper [52, 48]. Upper-Limb [52]. Upscaling [293]. Uruk [3]. USA [746, 712]. Usability [635, 663, 438]. Usage [370]. Use [235, 709, 439, 677, 359]. Useful [651].

User [7, 645, 658, 648, 585, 224, 208, 486, 12, 160]. User-Centered [208]. Users [659]. Using [599, 483, 379, 127, 177, 518, 81, 135, 472, 369, 65, 120, 383, 605, 178, 676, 116, 141, 180, 35, 386, 211, 62, 100, 305, 43, 149, 118, 251, 397, 628, 171, 266, 301, 184, 90, 113, 532, 202, 428, 160, 86, 64, 80, 475, 102, 146, 124, 444]. utlier [358]. utlier-Mining-as-a-Service [358].

Validity [431]. Value [38, 425]. Valued [264, 466]. Values [619]. Valve [108]. Variables [153]. Variant [292]. Various [700]. Varying [62]. Vascularized [91]. VDX [309]. Vector [683]. Vehicles [370]. Ventricle [134, 69]. Verbal [15, 62]. Vergence [213]. Verification [591, 379, 312]. versus [659]. Vertex [548, 547]. Very [25]. via [112, 280, 287, 522, 398, 228, 527, 364, 108, 77, 167]. Video [322, 219, 89, 642, 212]. Videos [137]. View [544, 263, 214, 588, 36, 587]. View-Based [588]. View-Update [263]. Viewpoints [596]. Views [235]. Virtual [644, 13, 3, 7, 11, 104, 329, 10, 191, 165, 320, 15, 14, 1, 2, 190, 712, 661, 640, 8, 324, 6]. Vision [722]. Visual [20, 657, 104, 340, 567, 16]. Visualization [680, 676]. Visualizing [567]. Volumetric [67]. Volumetry [133]. Voxel [74]. Voxel-Wise [74]. Vulnerabilities [163]. Vulnerability [461].

WABI [737]. Wait [155]. Wall [640]. Warehouse [678]. Warehouses [483]. Warmth [13]. Washington [746]. Wavelet [451]. Wavelets [292]. Weather [330]. Web [739, 565, 559, 177, 563, 492, 570, 667, 638, 489, 163, 443, 456, 567, 487]. Web-Based [492]. Websites [162]. Weighted [554, 138, 76, 86, 129]. Weights [271]. Well [425]. Well-Being [425]. Wheelchair [47]. Where [137]. White [54, 75, 142, 131, 76]. Whole [675]. Whole-Exome [675]. Widgets [492]. Wire [107]. Wireless [158, 445, 43]. Wise [74]. WISM [723]. within [460]. without [514]. Work [208]. Workers [414, 50]. Workshop [731, 740, 746, 743, 741, 243, 237, 742, 737, 232, 715]. Workshops [723]. World [3, 427, 164]. Worlds [644, 383, 201]. WoW [198]. Wrist [125]. WSDL [637]. WSNs [161].

X [107, 109]. X-Ray [107, 109]. XDBMSs [22]. XML [490, 22, 239]. XPath [39]. Xtream [274].

Zero [693, 508]. Zero-Knowledge [508]. ZigBee [443].

References

Kuyten:2012:FAG

[1] Pascal Kuyten, Timothy Bickmore, Svetlana Stoyanchev, Paul Piwek, and Helmut Prendinger. Fully automated generation of question-answer pairs for scripted virtual instruction. Lecture Notes in Computer Science, 7502:1–14, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33197-8_1/.

Li:2012:VA

[2] Weizi Li and Jan M. Allbeck. The virtual apprentice. Lecture Notes in Computer Science, 7502:15–27, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33197-8_2/.

Bogdanovych:2012:CUT

[3] Anton Bogdanovych, Kiran Ijaz, and Simeon Simoff. The city of Uruk: Teaching ancient history in a virtual world. Lecture Notes in Computer Science, 7502:28–35, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33197-8_3/.

Tewari:2012:ADC

[4] Anuj Tewari, Ingrid Liu, Carrie Cai, and John Canny. An analysis of the dialogic complexities in designing a question/answering based conversational agent for preschoolers. Lecture Notes in Computer Science, 7502:36–45, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33197-8_4/.

Bernardini:2012:BAS

[5] Sara Bernardini, Kaska Porayska-Pomsta, Tim J. Smith, and Katerina Avramides. Building autonomous social partners for autistic children. Lecture Notes in Computer Science, 7502:46–52, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-33197-8_5/.

deMelo:2012:EVA

[6] Celso M. de Melo, Peter Carnevale, and Jonathan Gratch. The effect of virtual agents’ emotion displays and appraisals on people’s decision making in negotiation. Lecture Notes in Computer Science, 7502: 53–66, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33197-8_6/.

Cafaro:2012:FIU

[7] Angelo Cafaro, Hannes Högni Vilhjálmsson, Timothy Bickmore, and Dirk Heylen. First impressions: Users’ judgments of virtual agents’ personality and interpersonal attitude in first encounters. Lecture Notes in Computer Science, 7502:67–80, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33197-8_7/.

Tsai:2012:SEC

[8] Jason Tsai, Emma Bowring, Stacy Marsella, Wendy Wood, and Milind Tambe. A study of emotional contagion with virtual characters. Lecture Notes in Computer Science, 7502:81–88, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33197-8_8/.

Ring:2012:LAC

[9] Lazlo Ring, Timothy Bickmore, and Daniel Schulman. Longitudinal affective computing. Lecture Notes in Computer Science, 7502:89–96, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33197-8_9/.

Ferreira:2012:GNR

[10] Nuno Ferreira, Samuel Mascarenhas, Ana Paiva, Frank Dignum, John Mc Breen, and Nick Degens. Generating norm-related emotions in virtual agents. Lecture Notes in Computer Science, 7502:97–104, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33197-8_10/.

Campos:2012:VAC

[11] Henrique Campos, Joana Campos, Carlos Martinho, and Ana Paiva. Virtual agents in conflict. Lecture Notes in Computer Science, 7502: 105–111, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33197-8_11/.

Segura:2012:HDY

[12] Elena Márquez Segura, Michael Kriegel, Ruth Aylett, Amol Deshmukh, and Henriette Cramer. How do you like me in this: User embodiment preferences for companion agents. Lecture Notes in Computer Science, 7502:112–125, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-33197-8_12/.

Bergmann:2012:SCM

[13] Kirsten Bergmann, Friederike Eyssel, and Stefan Kopp. A second chance to make a first impression? How appearance and nonverbal behavior affect perceived warmth and competence of virtual agents over time. Lecture Notes in Computer Science, 7502:126–138, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33197-8_13/.

Krum:2012:SMV

[14] David M. Krum, Evan A. Suma, and Mark Bolas. Spatial misregistration of virtual human audio: Implications of the precedence effect. Lecture Notes in Computer Science, 7502:139–145, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33197-8_14/.

Krishnan:2012:VHP

[15] Vaishnavi Krishnan, Adriana Foster, Regis Kopper, and Benjamin Lok. Virtual human personality masks: a human computation approach to modeling verbal personalities in virtual humans. Lecture Notes in Computer Science, 7502:146–152, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33197-8_15/.

Silvervarg:2012:EVG

[16] Annika Silvervarg, Kristin Raukola, Magnus Haake, and Agneta Gulz. The effect of visual gender on abuse in conversation with ECAs. Lecture Notes in Computer Science, 7502:153–160, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33197-8_16/.

Lee:2012:MSB

[17] Jina Lee and Stacy Marsella. Modeling speaker behavior: a comparison of two approaches. Lecture Notes in Computer Science, 7502: 161–174, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33197-8_17/.

vanWelbergen:2012:IMR

[18] Herwin van Welbergen, Dennis Reidsma, and Stefan Kopp. An incremental multimodal realizer for behavior co-articulation and coordination. Lecture Notes in Computer Science, 7502:175–188, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33197-8_18/.

Ribeiro:2012:TCM

[19] Tiago Ribeiro, Marco Vala, and Ana Paiva. Thalamus: Closing the mindbody loop in interactive embodied characters. Lecture Notes in Computer Science, 7502:189–195, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33197-8_19/.

AlMoubayed:2012:LRF

[20] Samer Al Moubayed, Gabriel Skantze, and Jonas Beskow. Lip-Reading: Furhat audio visual intelligibility of a back projected animated face. Lecture Notes in Computer Science, 7502:196–203, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-33197-8_20/.

Anonymous:2012:FMa

[21] Anonymous. Front matter. Lecture Notes in Computer Science, 7502: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-33197-8/1.

Ribeiro:2012:LSL

[22] Leonardo Andrade Ribeiro and Theo Härder. Leveraging the storage layer to support XML similarity joins in XDBMSs. Lecture Notes in Computer Science, 7503:1–14, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33074-2_1/.

Andruszkiewicz:2012:RRP

[23] Piotr Andruszkiewicz. Retention replacement in privacy preserving classification. Lecture Notes in Computer Science, 7503:15–28, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33074-2_2/.

Andrzejewski:2012:FIF

[24] Witold Andrzejewski and Bartosz Bebel. FOCUS: An index FO r C ontinuo US subsequence pattern queries. Lecture Notes in Computer Science, 7503:29–42, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33074-2_3/.

Bouchakri:2012:SIS

[25] Rima Bouchakri and Ladjel Bellatreche. Static and incremental selection of multi-table indexes for very large join queries. Lecture Notes in Computer Science, 7503:43–56, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33074-2_4/.

Bress:2012:ASP

[26] Sebastian Breß, Felix Beier, and Hannes Rauhe. Automatic selection of processing units for coprocessing in databases. Lecture Notes in Computer Science, 7503:57–70, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33074-2_5/.

Brzezinski:2012:DRD

[27] Jerzy Brzeziński and Arkadiusz Danilecki. D-ReServE: Distributed reliable service environment. Lecture Notes in Computer Science, 7503: 71–84, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33074-2_6/.

Budikova:2012:QLC

[28] Petra Budikova, Michal Batko, and Pavel Zezula. Query language for complex similarity queries. Lecture Notes in Computer Science, 7503: 85–98, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33074-2_7/.

Calil:2012:SRL

[29] Andre Calil and Ronaldo dos Santos Mello. SimpleSQL: a relational layer for SimpleDB. Lecture Notes in Computer Science, 7503:99–110, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33074-2_8/.

Castellanos:2012:KFS

[30] Camilo Castellanos and Dario Correal. KALCAS: a FrameworK for Semi-automatic ALignment of Data and Business ProCesses ArchitectureS. Lecture Notes in Computer Science, 7503:111–124, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33074-2_9/.

Catania:2012:TRS

[31] Barbara Catania and Giovanna Guerrini. Towards relaxed selection and join queries over data streams. Lecture Notes in Computer Science, 7503: 125–138, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33074-2_10/.

Dannecker:2012:EIE

[32] Lars Dannecker and Elena Vasilyeva. Efficient integration of external information into forecast models from the energy domain. Lecture Notes in Computer Science, 7503:139–152, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33074-2_11/.

Debruyne:2012:GMT

[33] Christophe Debruyne and Robert Meersman. GOSPL: a method and tool for fact-oriented hybrid ontology engineering. Lecture Notes in Computer Science, 7503:153–166, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33074-2_12/.

Dobreva:2012:GTT

[34] Veneta Dobreva and Martina-Cezara Albutiu. Get tracked: a triple store for RFID traceability data. Lecture Notes in Computer Science, 7503: 167–180, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33074-2_13/.

Htoo:2012:ANN

[35] Htoo Htoo, Yutaka Ohsawa, and Noboru Sonehara. Aggregate nearest neighbor search methods using SSMTA* algorithm on road-network. Lecture Notes in Computer Science, 7503:181–194, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-33074-2_14/.

Huang:2012:DVS

[36] Rong Huang, Rada Chirkova, and Yahya Fathi. Deterministic view selection for data-analysis queries: Properties and algorithms. Lecture Notes in Computer Science, 7503:195–208, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33074-2_15/.

Ivanova:2012:JTD

[37] Milena Ivanova and Martin Kersten. Just-in-time data distribution for analytical query processing. Lecture Notes in Computer Science, 7503: 209–222, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33074-2_16/.

Kalinichenko:2012:SCM

[38] Leonid Kalinichenko and Sergey Stupnikov. Synthesis of the canonical models for database integration preserving semantics of the value inventive data models. Lecture Notes in Computer Science, 7503:223– 239, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33074-2_17/.

Kosa:2012:CSP

[39] Balázs Kósa. Containment and satisfiability problem for XPath with recursion. Lecture Notes in Computer Science, 7503:240–253, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33074-2_18/.

Ladra:2012:ESI

[40] Susana Ladra, Oscar Pedreira, and Jose Duato. Exploiting SIMD instructions in current processors to improve classical string algorithms. Lecture Notes in Computer Science, 7503:254–267, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33074-2_19/.

Lehrack:2012:TFL

[41] Sebastian Lehrack and Sascha Saretz. A top-k filter for logic-based similarity conditions on probabilistic databases. Lecture Notes in Computer Science, 7503:268–281, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33074-2_20/.

Anonymous:2012:FMb

[42] Anonymous. Front matter. Lecture Notes in Computer Science, 7503: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-33074-2/1.

Lou:2012:WMS

[43] Zuozheng Lou, Peng Yao, and Dingguo Zhang. Wireless master-slave FES rehabilitation system using sEMG control. Lecture Notes in Computer Science, 7507:1–10, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33515-0_1/.

Pan:2012:IPD

[44] Lizheng Pan, Aiguo Song, Guozheng Xu, Huijun Li, and Baoguo Xu. Intelligent prescription-diagnosis function for rehabilitation training robot system. Lecture Notes in Computer Science, 7507:11–20, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33515-0_2/.

Nergui:2012:HBR

[45] Myagmarbayar Nergui, Yuki Yoshida, Nevrez Imamoglu, Jose Gonzalez, and Wenwei Yu. Human behavior recognition by a bio-monitoring mobile robot. Lecture Notes in Computer Science, 7507:21–30, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33515-0_3/.

Fan:2012:DTF

[46] Yuanjie Fan and Yuehong Yin. Differentiated time-frequency characteristics based real-time motion decoding for lower extremity rehabilitation exoskeleton robot. Lecture Notes in Computer Science, 7507: 31–40, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33515-0_4/.

Wei:2012:SMB

[47] Zhixuan Wei, Weidong Chen, and Jingchuan Wang. 3D semantic mapbased shared control for smart wheelchair. Lecture Notes in Computer Science, 7507:41–51, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33515-0_5/.

Rahman:2012:NSM

[48] Mohammad Habibur Rahman, Maarouf Saad, Jean Pierre Kenné, and P. S. Archambault. Nonlinear sliding mode control implementation of an upper limb exoskeleton robot to provide passive rehabilitation therapy. Lecture Notes in Computer Science, 7507:52–62, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33515-0_6/.

Wang:2012:SBC

[49] Baocheng Wang, Chenguang Yang, Zhijun Li, and Alex Smith. sEMGbased control of an exoskeleton robot arm. Lecture Notes in Computer Science, 7507:63–73, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33515-0_7/.

Reinhart:2012:AAH

[50] Gunther Reinhart, Ruediger Spillner, and Yi Shen. Approaches of applying human-robot-interaction-technologies to assist workers with musculoskeletal disorders in production. Lecture Notes in Computer Science, 7507:74–84, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33515-0_8/.

Chen:2012:SSF

[51] Yixiong Chen, Jin Hu, Feng Zhang, and Zengguang Hou. Simulation study of an FES-involved control strategy for lower limb rehabilitation robot. Lecture Notes in Computer Science, 7507:85–95, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33515-0_9/.

Lv:2012:DRR

[52] Chao Lv, Le Xie, Wei Shao, Hai long Yu, Yuan Wang, Jin wu Wang, and Ning Nan. Development of a rehabilitation robot for upper-limb movements. Lecture Notes in Computer Science, 7507:96–102, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33515-0_10/.

Zhang:2012:DON

[53] Gang Zhang, PinKuan Liu, and Han Ding. Dynamic optimization with a new performance index for a 2-DoF translational parallel manipulator. Lecture Notes in Computer Science, 7507:103–115, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33515-0_11/.

Guo:2012:RPP

[54] Qiwei Guo, Yueming Hu, Zhifu Li, and Ge Ma. Research of piezoelectric printing actuator for high-power white LED phosphor coating. Lecture Notes in Computer Science, 7507:116–126, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33515-0_12/.

Gao:2012:DOP

[55] Jian Gao, Zhiliang Wang, Yanyun Liu, Chuanxia Jian, and Xin Chen. Development of OLED panel defect detection system through improved Otsu algorithm. Lecture Notes in Computer Science, 7507:127–134, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33515-0_13/.

Zhang:2012:LCS

[56] Jinsong Zhang and Jianhua Zhang. The linkage control strategy for the two-phase flow dispensing system. Lecture Notes in Computer Science, 7507:135–144, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-33515-0_14/.

Ge:2012:TTF

[57] Junfeng Ge, Yuneng Lai, Yuanhao Huang, and Jianhua Zhang. The transient temperature field measurement system for laser bonding process. Lecture Notes in Computer Science, 7507:145–152, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33515-0_15/.

Yan:2012:MEI

[58] Wei Yan, Enrong Wang, Yang Zhao, Rakheja Subhash, and Chunyi Su. Modeling of electromagnetic interference noise mechanism for magnetorheological damper. Lecture Notes in Computer Science, 7507:153–163, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33515-0_16/.

Hocherl:2012:IBA

[59] Johannes Höcherl and Thomas Schlegl. An image based algorithm to safely locate human extremities for human-robot collaboration. Lecture Notes in Computer Science, 7507:164–175, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33515-0_17/.

Zhou:2012:RMD

[60] Hangfei Zhou, Jian Fei, Gen Pan, Weixin Yan, Zhuang Fu, and Yanzheng Zhao. Research of a multi-DOF pathological sampling flexible robot. Lecture Notes in Computer Science, 7507:176–185, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-33515-0_18/.

Rossmann:2012:CSN

[61] Juergen Rossmann, Michael Schluse, Christian Schlette, and Ralf Waspe. Control by 3D simulation — a new eRobotics approach to control design in automation. Lecture Notes in Computer Science, 7507:186–197, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33515-0_19/.

Liu:2012:HIE

[62] Peter Liu and Chang-En Yang. Human intention estimation using timevarying fuzzy Markov models for natural non-verbal human robot interface. Lecture Notes in Computer Science, 7507:198–206, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33515-0_20/.

Anonymous:2012:FMc

[63] Anonymous. Front matter. Lecture Notes in Computer Science, 7507: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-33515-0/1.

Yan:2012:MNP

[64] Jingwen Yan, Shannon L. Risacher, and Sungeun Kim. Multimodal neuroimaging predictors for cognitive performance using structured sparse learning. Lecture Notes in Computer Science, 7509:1–17, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33530-3_1/.

Dyrba:2012:CDM

[65] Martin Dyrba, Michael Ewers, and Martin Wegrzyn. Combining DTI and MRI for the automated detection of Alzheimer’s disease using a large European multicenter dataset. Lecture Notes in Computer Science, 7509: 18–28, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33530-3_2/.

Jahanshad:2012:GPL

[66] Neda Jahanshad, Gautam Prasad, and Arthur W. Toga. Genetics of path lengths in brain connectivity networks: HARDI-based maps in 457 adults. Lecture Notes in Computer Science, 7509:29–40, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33530-3_3/.

Nir:2012:CNB

[67] Talia M. Nir, Neda Jahanshad, and Arthur W. Toga. Connectivity network breakdown predicts imminent volumetric atrophy in early mild cognitive impairment. Lecture Notes in Computer Science, 7509: 41–50, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33530-3_4/.

Yoldemir:2012:DER

[68] Burak Yoldemir, Bernard Ng, and Rafeef Abugharbieh. Deconfounding the effects of resting state activity on task activation detection in fMRI. Lecture Notes in Computer Science, 7509:51–60, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33530-3_5/.

Shi:2012:HRF

[69] Jie Shi, Paul M. Thompson, and Yalin Wang. Hyperbolic Ricci flow and its application in studying lateral ventricle morphometry. Lecture Notes in Computer Science, 7509:61–76, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33530-3_6/.

Landman:2012:DWR

[70] Bennett A. Landman, Xue Yang, and Hakmook Kang. Do we really need robust and alternative inference methods for brain MRI? Lecture Notes in Computer Science, 7509:77–93, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33530-3_7/.

Zhang:2012:SPB

[71] Daoqiang Zhang, Qimiao Guo, Guorong Wu, and Dinggang Shen. Sparse patch-based label fusion for multi-atlas segmentation. Lecture Notes in Computer Science, 7509:94–102, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33530-3_8/.

Awate:2012:HMT

[72] Suyash P. Awate, Peihong Zhu, and Ross T. Whitaker. How many templates does it take for a good segmentation?: Error analysis in multiatlas segmentation as a function of database size. Lecture Notes in Computer Science, 7509:103–114, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33530-3_9/.

Iglesias:2012:GMP

[73] Juan Eugenio Iglesias and Mert Rory Sabuncu. A generative model for probabilistic label fusion of multimodal data. Lecture Notes in Computer Science, 7509:115–133, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33530-3_10/.

Liu:2012:SND

[74] Wei Liu, Xiaozheng Liu, Xiaofu He, Zhenyu Zhou, and Ying Wen. Spatial normalization of diffusion tensor images with voxel-wise reconstruction of the diffusion gradient direction. Lecture Notes in Computer Science, 7509: 134–146, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33530-3_11/.

Jin:2012:APH

[75] Yan Jin, Yonggang Shi, Liang Zhan, and Junning Li. Automatic population HARDI white matter tract clustering by label fusion of multiple tract atlases. Lecture Notes in Computer Science, 7509:147–156, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33530-3_12/.

Strumia:2012:CCS

[76] Maddalena Strumia and Constantin Anastasopoulos. Comparative characterisation of susceptibility weighted MRI for brain white matter lesions in MS. Lecture Notes in Computer Science, 7509:157–166, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33530-3_13/.

Zhang:2012:CFA

[77] Tuo Zhang, Lei Guo, Hanbo Chen, Xintao Hu, and Kaiming Li. Constructing fiber atlases for functional ROIs via fMRI-guided DTI image registration. Lecture Notes in Computer Science, 7509:167–174, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33530-3_14/.

Jie:2012:SFS

[78] Biao Jie, Daoqiang Zhang, Chong-Yaw Wee, and Dinggang Shen. Structural feature selection for connectivity network-based MCI diagnosis. Lecture Notes in Computer Science, 7509:175–184, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-33530-3_15/.

Wang:2012:GSI

[79] Yaping Wang, Hongjun Jia, Pew-Thian Yap, and Bo Cheng. Groupwise segmentation improves neuroimaging classification accuracy. Lecture Notes in Computer Science, 7509:185–193, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33530-3_16/.

Yap:2012:DDU

[80] Pew-Thian Yap and Dinggang Shen. DWI denoising using spatial, angular, and radiometric filtering. Lecture Notes in Computer Science, 7509: 194–202, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33530-3_17/.

Cao:2012:MER

[81] Fang Cao, Olivier Commowick, and Elise Bannier. MRI estimation of T_{1} relaxation time using a constrained optimization algorithm. Lecture Notes in Computer Science, 7509:203–214, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33530-3_18/.

Maumet:2012:RCB

[82] Camille Maumet, Pierre Maurel, and Jean-Christophe Ferré. Robust cerebral blood flow map estimation in arterial spin labeling. Lecture Notes in Computer Science, 7509:215–224, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33530-3_19/.

Anonymous:2012:BMa

[83] Anonymous. Back matter. Lecture Notes in Computer Science, 7509: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-33530-3/1.

Anonymous:2012:FMd

[84] Anonymous. Front matter. Lecture Notes in Computer Science, 7509: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-33530-3/1.

Freiman:2012:RAP

[85] M. Freiman, S. D. Voss, R. V. Mulkern, J. M. Perez-Rossello, and M. J. Callahan. Reliable assessment of perfusivity and diffusivity from diffusion imaging of the body. Lecture Notes in Computer Science, 7510:1–9, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33415-3_1/.

Wolz:2012:MOA

[86] Robin Wolz, Chengwen Chu, Kazunari Misawa, Kensaku Mori, and Daniel Rueckert. Multi-organ abdominal CT segmentation using hierarchically weighted subject-specific atlases. Lecture Notes in Computer Science, 7510:10–17, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33415-3_2/.

Diotte:2012:RFD

[87] Benoit Diotte, Pascal Fallavollita, Lejing Wang, Simon Weidert, and Peter-Helmut Thaller. Radiation-free drill guidance in interlocking of intramedullary nails. Lecture Notes in Computer Science, 7510:18–25, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33415-3_3/.

Zuo:2012:DER

[88] Siyang Zuo, Takeshi Ohdaira, Kenta Kuwana, Yoshihiro Nagao, and Satoshi Ieiri. Developing essential rigid-flexible outer sheath to enable novel multi-piercing surgery. Lecture Notes in Computer Science, 7510: 26–33, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33415-3_4/.

Haro:2012:SGC

[89] Benjamín Béjar Haro, Luca Zappella, and René Vidal. Surgical gesture classification from video data. Lecture Notes in Computer Science, 7510:34–41, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33415-3_5/.

Schneider:2012:RUP

[90] Caitlin Schneider, Ali Baghani, Robert Rohling, and Septimiu Salcudean. Remote ultrasound palpation for robotic interventions using absolute elastography. Lecture Notes in Computer Science, 7510:42–49, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33415-3_6/.

Peterlik:2012:MRT

[91] Igor Peterlík, Christian Duriez, and Stéphane Cotin. Modeling and realtime simulation of a vascularized liver tissue. Lecture Notes in Computer Science, 7510:50–57, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33415-3_7/.

Xu:2012:EOC

[92] Yanwu Xu, Jiang Liu, Stephen Lin, Dong Xu, Carol Y. Cheung, Tin Aung, and Tien Yin Wong. Efficient optic cup detection from intra-image learning with retinal structure priors. Lecture Notes in Computer Science, 7510:58–65, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33415-3_8/.

Bousleiman:2012:PBD

[93] Habib Bousleiman, Christof Seiler, Tateyuki Iizuka, Lutz-Peter Nolte, and Mauricio Reyes. Population-based design of mandibular plates based on bone quality and morphology. Lecture Notes in Computer Science, 7510:66–73, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33415-3_9/.

Song:2012:TAD

[94] Yang Song, Weidong Cai, Yun Zhou, and Dagan Feng. Thoracic abnormality detection with data adaptive structure estimation. Lecture Notes in Computer Science, 7510:74–81, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33415-3_10/.

Cheng:2012:DTL

[95] Bo Cheng, Daoqiang Zhang, and Dinggang Shen. Domain transfer learning for MCI conversion prediction. Lecture Notes in Computer Science, 7510:82–90, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33415-3_11/.

Bano:2012:SPL

[96] J. Bano, A. Hostettler, S. A. Nicolau, S. Cotin, C. Doignon, H. S. Wu, and M. H. Huang. Simulation of pneumoperitoneum for laparoscopic surgery planning. Lecture Notes in Computer Science, 7510:91–98, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33415-3_12/.

Pan:2012:IKR

[97] Binbin Pan, James J. Xia, Peng Yuan, Jaime Gateno, Horace H. S. Ip, and Qizhen He. Incremental kernel ridge regression for the prediction of soft tissue deformations. Lecture Notes in Computer Science, 7510: 99–106, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33415-3_13/.

George:2012:FMC

[98] Jose George, Kathleen Vunckx, Elke Van de Casteele, and Sabine Tejpar. Fuzzy multi-class statistical modeling for efficient total lesion metabolic activity estimation from realistic PET images. Lecture Notes in Computer Science, 7510:107–114, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33415-3_14/.

Nguyen:2012:SCP

[99] Kien Nguyen, Anindya Sarkar, and Anil K. Jain. Structure and context in prostatic gland segmentation and classification. Lecture Notes in Computer Science, 7510:115–123, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33415-3_15/.

Liu:2012:QCT

[100] Yinxiao Liu, Punam K. Saha, and Ziyue Xu. Quantitative characterization of trabecular bone micro-architecture using tensor scale and multi-detector CT imaging. Lecture Notes in Computer Science, 7510: 124–131, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33415-3_16/.

Singh:2012:GSF

[101] Nikhil Singh, Angela Y. Wang, Preethi Sankaranarayanan, and P. Thomas Fletcher. Genetic, structural and functional imaging biomarkers for early detection of conversion from MCI to AD. Lecture Notes in Computer Science, 7510:132–140, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33415-3_17/.

Zhan:2012:RMS

[102] Yiqiang Zhan, Dewan Maneesh, Martin Harder, and Xiang Sean Zhou. Robust MR spine detection using hierarchical learning and local articulated model. Lecture Notes in Computer Science, 7510:141–148, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33415-3_18/.

Duong:2012:SRB

[103] D. Duong, D. Shastri, P. Tsiamyrtzis, and I. Pavlidis. Spatiotemporal reconstruction of the breathing function. Lecture Notes in Computer Science, 7510:149–156, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33415-3_19/.

Cruz-Roa:2012:VLS

[104] Angel Cruz-Roa, Fabio González, Joseph Galaro, Alexander R. Judkins, and David Ellison. A visual latent semantic approach for automatic analysis and interpretation of anaplastic medulloblastoma virtual slides. Lecture Notes in Computer Science, 7510:157–164, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33415-3_20/.

Anonymous:2012:FMe

[105] Anonymous. Front matter. Lecture Notes in Computer Science, 7510: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-33415-3/1.

Kutra:2012:AMM

[106] Dominik Kutra, Axel Saalbach, Helko Lehmann, Alexandra Groth, and Sebastian P. M. Dries. Automatic multi-model-based segmentation of the left atrium in cardiac MRI scans. Lecture Notes in Computer Science, 7511:1–8, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33418-4_1/.

Bismuth:2012:CSE

[107] Vincent Bismuth, Régis Vaillant, Hugues Talbot, and Laurent Najman. Curvilinear structure enhancement with the polygonal path image- application to guide-wire segmentation in X-ray fluoroscopy. Lecture Notes in Computer Science, 7511:9–16, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33418-4_2/.

Wang:2012:CTO

[108] Peng Wang, Yefeng Zheng, Matthias John, and Dorin Comaniciu. Catheter tracking via online learning for dynamic motion compensation in transcatheter aortic valve implantation. Lecture Notes in Computer Science, 7511:17–24, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33418-4_3/.

Housden:2012:ERT

[109] R. J. Housden, A. Arujuna, Y. Ma, N. Nijhof, G. Gijsbers, R. Bullens, and M. O’Neill. Evaluation of a real-time hybrid three-dimensional echo and X-ray imaging system for guidance of cardiac catheterisation procedures. Lecture Notes in Computer Science, 7511:25–32, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33418-4_4/.

Rapaka:2012:LEL

[110] S. Rapaka, T. Mansi, B. Georgescu, M. Pop, G. A. Wright, A. Kamen, and Dorin Comaniciu. LBM-EP: Lattice-boltzmann method for fast cardiac electrophysiology simulation from 3D images. Lecture Notes in Computer Science, 7511:33–40, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33418-4_5/.

Marchesseau:2012:CMP

[111] Stéphanie Marchesseau, Hervé Delingette, Maxime Sermesant, and Kawal Rhode. Cardiac mechanical parameter calibration based on the unscented transform. Lecture Notes in Computer Science, 7511:41–48, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33418-4_6/.

Bernardis:2012:TSA

[112] Elena Bernardis, Ender Konukoglu, Yangming Ou, Dimitris N. Metaxas, and Benoit Desjardins. Temporal shape analysis via the spectral signature. Lecture Notes in Computer Science, 7511:49–56, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33418-4_7/.

Siless:2012:JTB

[113] Viviana Siless, Joan Glaunès, Pamela Guevara, Jean-François Mangin, and Cyril Poupon. Joint T1 and brain fiber log-demons registration using currents to model geometry. Lecture Notes in Computer Science, 7511:57–65, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33418-4_8/.

Steger:2012:ASB

[114] Sebastian Steger and Stefan Wesarg. Automated skeleton based multimodal deformable registration of head&neck datasets. Lecture Notes in Computer Science, 7511:66–73, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33418-4_9/.

Schmidt-Richberg:2012:LRI

[115] Alexander Schmidt-Richberg, Jan Ehrhardt, René Werner, and Heinz Handels. Lung registration with improved fissure alignment by integration of pulmonary lobe segmentation. Lecture Notes in Computer Science, 7511:74–81, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33418-4_10/.

Hacihaliloglu:2012:UCR

[116] Ilker Hacihaliloglu, Anna Brounstein, Pierre Guy, Antony Hodgson, and Rafeef Abugharbieh. 3D ultrasound-CT registration in orthopaedic trauma using GMM registration with optimized particle simulationbased data reduction. Lecture Notes in Computer Science, 7511:82– 89, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33418-4_11/.

Wu:2012:HAG

[117] Guorong Wu, Minjeong Kim, Qian Wang, and Dinggang Shen. Hierarchical attribute-guided symmetric diffeomorphic registration for MR brain images. Lecture Notes in Computer Science, 7511:90–97, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33418-4_12/.

Mirzaalian:2012:UBF

[118] Hengameh Mirzaalian, Tim K. Lee, and Ghassan Hamarneh. Uncertaintybased feature learning for skin lesion matching using a high order MRF optimization framework. Lecture Notes in Computer Science, 7511: 98–105, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33418-4_13/.

Hanaoka:2012:ACA

[119] Shouhei Hanaoka, Yoshitaka Masutani, Mitsutaka Nemoto, and Yukihiro Nomura. Automatic categorization of anatomical landmark-local appearances based on diffeomorphic demons and spectral clustering for constructing detector ensembles. Lecture Notes in Computer Science, 7511: 106–113, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33418-4_14/.

El-Baz:2012:NAG

[120] Ayman El-Baz, Fahmi Khalifa, Ahmed Elnakib, Matthew Nitzken, and Ahmed Soliman. A novel approach for global lung registration using 3D Markov–Gibbs appearance model. Lecture Notes in Computer Science, 7511:114–121, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-33418-4_15/.

Shackleford:2012:ARU

[121] James A. Shackleford, Qi Yang, Ana M. Lourenço, Nadya Shusharina, and Nagarajan Kandasamy. Analytic regularization of uniform cubic B-spline deformation fields. Lecture Notes in Computer Science, 7511: 122–129, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33418-4_16/.

Seiler:2012:SMP

[122] Christof Seiler, Xavier Pennec, and Mauricio Reyes. Simultaneous multiscale polyaffine registration by incorporating deformation statistics. Lecture Notes in Computer Science, 7511:130–137, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-33418-4_17/.

Li:2012:FDT

[123] Junning Li, Yonggang Shi, Giang Tran, Ivo Dinov, Danny J. J. Wang, and Arthur W. Toga. Fast diffusion tensor registration with exact reorientation and regularization. Lecture Notes in Computer Science, 7511: 138–145, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33418-4_18/.

Zhang:2012:RBS

[124] Minqi Zhang, Fang Li, Ying He, Shi Lin, Defeng Wang, and Lok Ming Lui. Registration of brainstem surfaces in adolescent idiopathic scoliosis using discrete Ricci flow. Lecture Notes in Computer Science, 7511: 146–154, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33418-4_19/.

vandeGiessen:2012:GRR

[125] Martijn van de Giessen, Frans M. Vos, Cornelis A. Grimbergen, and Lucas J. van Vliet. Groupwise rigid registration of wrist bones. Lecture Notes in Computer Science, 7511:155–162, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33418-4_20/.

Anonymous:2012:FMf

[126] Anonymous. Front matter. Lecture Notes in Computer Science, 7511: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-33418-4/1.

Bilgic:2012:ADS

[127] Berkin Bilgic, Kawin Setsompop, Julien Cohen-Adad, Van Wedeen, and Lawrence L. Wald. Accelerated diffusion spectrum imaging with compressed sensing using adaptive dictionaries. Lecture Notes in Computer Science, 7512:1–9, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-33454-2_1/.

Merlet:2012:PDL

[128] Sylvain Merlet, Emmanuel Caruyer, and Rachid Deriche. Parametric dictionary learning for modeling EAP and ODF in diffusion MRI. Lecture Notes in Computer Science, 7512:10–17, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33454-2_2/.

Yap:2012:RED

[129] Pew-Thian Yap and Dinggang Shen. Resolution enhancement of diffusionweighted images by local fiber profiling. Lecture Notes in Computer Science, 7512:18–25, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-33454-2_3/.

Cardoso:2012:GSB

[130] M. Jorge Cardoso, Gavin Winston, Marc Modat, Shiva Keihaninejad, and John Duncan. Geodesic shape-based averaging. Lecture Notes in Computer Science, 7512:26–33, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33454-2_4/.

Savadjiev:2012:MSC

[131] Peter Savadjiev, Yogesh Rathi, Sylvain Bouix, Ragini Verma, and Carl-Fredrik Westin. Multi-scale characterization of white matter tract geometry. Lecture Notes in Computer Science, 7512:34–41, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33454-2_5/.

Vogel:2012:OAG

[132] Jakob Vogel, Tobias Reichl, José Gardiazabal, Nassir Navab, and Tobias Lasser. Optimization of acquisition geometry for intra-operative tomographic imaging. Lecture Notes in Computer Science, 7512:42– 49, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33454-2_6/.

Iglesias:2012:IPU

[133] Juan Eugenio Iglesias, Mert Rory Sabuncu, and Koen Van Leemput. Incorporating parameter uncertainty in Bayesian segmentation models: Application to hippocampal subfield volumetry. Lecture Notes in Computer Science, 7512:50–57, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33454-2_7/.

Huang:2012:DAM

[134] Xiaojie Huang, Donald P. Dione, Colin B. Compas, Xenophon Papademetris, and Ben A. Lin. A dynamical appearance model based on multiscale sparse representation: Segmentation of the left ventricle from 4D echocardiography. Lecture Notes in Computer Science, 7512: 58–65, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33454-2_8/.

Cuingnet:2012:ADS

[135] Rémi Cuingnet, Raphael Prevost, David Lesage, Laurent D. Cohen, and Benoît Mory. Automatic detection and segmentation of kidneys in 3D CT images using random forests. Lecture Notes in Computer Science, 7512:66–74, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33454-2_9/.

Konukoglu:2012:NAF

[136] Ender Konukoglu, Ben Glocker, Darko Zikic, and Antonio Criminisi. Neighbourhood approximation forests. Lecture Notes in Computer Science, 7512:75–82, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-33454-2_10/.

Kwitt:2012:RUV

[137] Roland Kwitt, Nuno Vasconcelos, Sharif Razzaque, and Stephen Aylward. Recognition in ultrasound videos: Where am I? Lecture Notes in Computer Science, 7512:83–90, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33454-2_11/.

Rivaz:2012:SSW

[138] Hassan Rivaz and D. Louis Collins. Self-similarity weighted mutual information: a new nonrigid image registration metric. Lecture Notes in Computer Science, 7512:91–98, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33454-2_12/.

Lekadir:2012:IPP

[139] Karim Lekadir, Alejandro F. Frangi, and Guang-Zhong Yang. Inter-point Procrustes: Identifying regional and large differences in 3D anatomical shapes. Lecture Notes in Computer Science, 7512:99–106, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33454-2_13/.

Risholm:2012:SOH

[140] Petter Risholm, Firdaus Janoos, Jennifer Pursley, Andriy Fedorov, and Clare Tempany. Selection of optimal hyper-parameters for estimation of uncertainty in MRI-TRUS registration of the prostate. Lecture Notes in Computer Science, 7512:107–114, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33454-2_14/.

Heinrich:2012:GOD

[141] Mattias P. Heinrich, Mark Jenkinson, Sir Michael Brady, and Julia A. Schnabel. Globally optimal deformable registration on a minimum spanning tree using dense displacement sampling. Lecture Notes in Computer Science, 7512:115–122, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33454-2_15/.

ODonnell:2012:UGR

[142] Lauren J. O’Donnell, William M. Wells III, Alexandra J. Golby, and Carl-Fredrik Westin. Unbiased groupwise registration of white matter tractography. Lecture Notes in Computer Science, 7512:123–130, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33454-2_16/.

Ye:2012:RML

[143] Dong Hye Ye, Jihun Hamm, Dongjin Kwon, Christos Davatzikos, and Kilian M. Pohl. Regional manifold learning for deformable registration of brain MR images. Lecture Notes in Computer Science, 7512: 131–138, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33454-2_17/.

Datteri:2012:ERT

[144] Ryan D. Datteri and Benoît M. Dawant. Estimation and reduction of target registration error. Lecture Notes in Computer Science, 7512: 139–146, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33454-2_18/.

Feragen:2012:HSG

[145] Aasa Feragen, Jens Petersen, Megan Owen, Pechin Lo, and Laura H. Thomsen. A hierarchical scheme for geodesic anatomical labeling of airway trees. Lecture Notes in Computer Science, 7512:147–155, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33454-2_19/.

Zhang:2012:IGN

[146] Pei Zhang, Pew-Thian Yap, Dinggang Shen, and Timothy F. Cootes. Initialising groupwise non-rigid registration using multiple parts+geometry models. Lecture Notes in Computer Science, 7512:156–163, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33454-2_20/.

Anonymous:2012:FMg

[147] Anonymous. Front matter. Lecture Notes in Computer Science, 7512: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-33454-2/1.

Xiao:2012:ALS

[148] Jian Xiao, Yu Zhang, Shuwei Chen, and Huashan Yu. An applicationlevel scheduling with task bundling approach for many-task computing in heterogeneous environments. Lecture Notes in Computer Science, 7513:1–13, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-35606-3_1/.

Meng:2012:DAS

[149] Jintao Meng, Jianrui Yuan, Jiefeng Cheng, Yanjie Wei, and Shengzhong Feng. DGraph: Algorithms for shortgun reads assembly using De Bruijn graph. Lecture Notes in Computer Science, 7513:14–21, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-35606-3_2/.

Wang:2012:KBA

[150] Yizhuo Wang, Weixing Ji, Feng Shi, Qi Zuo, and Ning Deng. Knowledgebased adaptive self-scheduling. Lecture Notes in Computer Science, 7513: 22–32, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-35606-3_3/.

Talpur:2012:CLA

[151] Shahnawaz Talpur, Feng Shi, and Yizhuo Wang. Communication locality analysis of triplet-based hierarchical interconnection network in chip multiprocessor. Lecture Notes in Computer Science, 7513:33–41, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-35606-3_4/.

Lee:2012:SSS

[152] Jae-Young Lee. A scoring system for short answers on the test in a large group. Lecture Notes in Computer Science, 7513:42–47, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-35606-3_5/.

Duncan:2012:RVD

[153] Ralph Duncan, Peder Jungck, Kenneth Ross, and Dwight Mulcahy. Reference variables for dynamic, reliable packet operations. Lecture Notes in Computer Science, 7513:48–60, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-35606-3_6/.

Salim:2012:HCN

[154] Shelly Salim, Christian H. W. Oey, and Sangman Moh. Are heterogeneous cellular networks superior to homogeneous ones? Lecture Notes in Computer Science, 7513:61–68, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-35606-3_7/.

Zhang:2012:DSW

[155] Longbo Zhang, Chen Yu, and Hai Jin. Dynamic spray and wait routing protocol for delay tolerant networks. Lecture Notes in Computer Science, 7513:69–76, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-35606-3_8/.

Soltani:2012:DPA

[156] Narjes Soltani, Ehsan Mousavi Khaneghah, Mohsen Sharifi, and Seyedeh Leili Mirtaheri. A dynamic popularity-aware load balancing algorithm for structured P2P systems. Lecture Notes in Computer Science, 7513:77–84, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-35606-3_9/.

Lin:2012:NCS

[157] Jinzhi Lin, Ying Wu, Gongyi Wu, and Jingdong Xu. NCCPIS: a cosimulation tool for networked control and cyber-physical system evaluation. Lecture Notes in Computer Science, 7513:85–93, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-35606-3_10/.

Ahmed:2012:DST

[158] Muhammad Ahmed, Xu Huang, and Dharmendra Sharma. Dempstershafer theory to identify insider attacker in wireless sensor network. Lecture Notes in Computer Science, 7513:94–100, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-35606-3_11/.

Cheng:2012:MIC

[159] Bo-Chao Cheng, Guo-Tan Liao, Ching-Kai Lin, Shih-Chun Hsu, Ping-Hai Hsu, and Jong Hyuk Park. MIB-ITrace-CP: An improvement of ICMPbased traceback efficiency in network forensic analysis. Lecture Notes in Computer Science, 7513:101–109, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-35606-3_12/.

Wang:2012:BRR

[160] Ding Wang, Chun guang Ma, Sen dong Zhao, and Chang li Zhou. Breaking a robust remote user authentication scheme using Smart Cards. Lecture Notes in Computer Science, 7513:110–118, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-35606-3_13/.

Memon:2012:APP

[161] Irfana Memon. An analysis of privacy preserving data aggregation protocols for WSNs. Lecture Notes in Computer Science, 7513:119–128, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-35606-3_14/.

Krishnaveni:2012:HOJ

[162] R. Krishnaveni, C. Chellappan, and R. Dhanalakshmi. Hybrid obfuscated Javascript strength analysis system for detection of malicious websites. Lecture Notes in Computer Science, 7513:129–137, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-35606-3_15/.

Lee:2012:DMW

[163] Taeseung Lee, Giyoun Won, Seongje Cho, Namje Park, and Dongho Won. Detection and mitigation of Web application vulnerabilities based on security testing. Lecture Notes in Computer Science, 7513:138–144, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-35606-3_16/.

Meng:2012:SWA

[164] Jintao Meng, Jianrui Yuan, Jiefeng Cheng, Yanjie Wei, and Shengzhong Feng. Small world asynchronous parallel model for genome assembly. Lecture Notes in Computer Science, 7513:145–155, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-35606-3_17/.

Jiang:2012:UNG

[165] Haitao Jiang, Yun Xu, Yin Liao, Guojie Jin, and Guoliang Chen. UKCF: a new graphics driver cross-platform translation framework for virtual machines. Lecture Notes in Computer Science, 7513:156–163, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-35606-3_18/.

Li:2012:FPG

[166] Yongnan Li, Limin Xiao, Aihua Liang, Yao Zheng, and Li Ruan. Fast parallel Garner Algorithm for Chinese Remainder Theorem. Lecture Notes in Computer Science, 7513:164–171, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-35606-3_19/.

Zhou:2012:DFD

[167] Xu Zhou, Kai Lu, Xicheng Lu, Xiaoping Wang, and Baohua Fan. dMPI: Facilitating debugging of MPI programs via deterministic message passing. Lecture Notes in Computer Science, 7513:172–179, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-35606-3_20/.

Anonymous:2012:FMh

[168] Anonymous. Front matter. Lecture Notes in Computer Science, 7513: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-35606-3/1.

Deb:2012:AEM

[169] Kalyanmoy Deb. Advances in evolutionary multi-objective optimization. Lecture Notes in Computer Science, 7515:1–26, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.com/chapter/10.1007/978-3-642-33119-0_1/.

DiPenta:2012:SMS

[170] Massimiliano Di Penta. SBSE meets software maintenance: Achievements and open problems. Lecture Notes in Computer Science, 7515:27–28, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/accesspage/chapter/10.1007/978-
3-642-33119-0_2.

Poulding:2012:THP

[171] Simon Poulding. Tutorial: High performance SBSE using commodity graphics cards. Lecture Notes in Computer Science, 7515:29, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/accesspage/chapter/10.1007/978-
3-642-33119-0_3.

Aitken:2012:ERN

[172] Jonathan M. Aitken, Rob Alexander, Tim Kelly, and Simon Poulding. Evolving robust networks for systems-of-systems. Lecture Notes in Computer Science, 7515:30–44, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33119-0_4/.

Arito:2012:ASS

[173] Franco Arito, Francisco Chicano, and Enrique Alba. On the application of SAT solvers to the test suite minimization problem. Lecture Notes in Computer Science, 7515:45–59, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33119-0_5/.

Barros:2012:EIR

[174] Márcio de Oliveira Barros. Evaluating the importance of randomness in search-based software engineering. Lecture Notes in Computer Science, 7515:60–74, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33119-0_6/.

Bavota:2012:PDL

[175] Gabriele Bavota, Filomena Carnevale, Andrea De Lucia, and Massimiliano Di Penta. Putting the developer in-the-loop: An interactive GA for software re-modularization. Lecture Notes in Computer Science, 7515: 75–89, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33119-0_7/.

Bhattacharya:2012:OTS

[176] Neelesh Bhattacharya, Olfat El-Mahi, Etienne Duclos, and Giovanni Beltrame. Optimizing threads schedule alignments to expose the interference bug pattern. Lecture Notes in Computer Science, 7515:90–104, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33119-0_8/.

Bozkurt:2012:ORT

[177] Mustafa Bozkurt and Mark Harman. Optimised realistic test input generation using Web services. Lecture Notes in Computer Science, 7515: 105–120, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33119-0_9/.

Ghaith:2012:ISS

[178] Shadi Ghaith and Mel Ó Cinnéide. Improving software security using search-based refactoring. Lecture Notes in Computer Science, 7515: 121–135, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33119-0_10/.

Iqbal:2012:CSB

[179] Muhammad Zohaib Iqbal, Andrea Arcuri, and Lionel Briand. Combining search-based and adaptive random testing strategies for environment model-based testing of real-time embedded systems. Lecture Notes in Computer Science, 7515:136–151, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33119-0_11/.

Hruba:2012:TCP

[180] Vendula Hrubá, Bohuslav Křena, Zdeněk Letko, Shmuel Ur, and Tomáš Vojnar. Testing of concurrent programs using genetic algorithms. Lecture Notes in Computer Science, 7515:152–167, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33119-0_12/.

Lopez-Herrejon:2012:REF

[181] Roberto Erick Lopez-Herrejon, José A. Galindo, David Benavides, and Sergio Segura. Reverse engineering feature models with evolutionary algorithms: An exploratory study. Lecture Notes in Computer Science, 7515: 168–182, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33119-0_13/.

Millard:2012:SPO

[182] Alan G. Millard, David R. White, and John A. Clark. Searching for paretooptimal randomised algorithms. Lecture Notes in Computer Science, 7515: 183–197, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33119-0_14/.

Ramirez:2012:ARG

[183] Andres J. Ramirez, Erik M. Fredericks, Adam C. Jensen, and Betty H. C. Cheng. Automatically RELAXing a goal model to cope with uncertainty. Lecture Notes in Computer Science, 7515:198–212, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33119-0_15/.

Sakti:2012:BSB

[184] Abdelilah Sakti, Yann-Gaël Guéhéneuc, and Gilles Pesant. Boosting search based testing by using constraint based testing. Lecture Notes in Computer Science, 7515:213–227, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33119-0_16/.

Tomasi:2012:DDR

[185] Alex Tomasi, Alessandro Marchetto, and Chiara Di Francescomarino. Domain-driven reduction optimization of recovered business processes. Lecture Notes in Computer Science, 7515:228–243, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.com/chapter/10.1007/978-3-642-33119-0_17/.

Yoo:2012:EHC

[186] Shin Yoo. Evolving human competitive spectra-based fault localisation techniques. Lecture Notes in Computer Science, 7515:244–258, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33119-0_18/.

Colanzi:2012:ASB

[187] Thelma Elita Colanzi and Silvia Regina Vergilio. Applying search based optimization to software product line architectures: Lessons learned. Lecture Notes in Computer Science, 7515:259–266, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-33119-0_19/.

Etemaadi:2012:PSS

[188] Ramin Etemaadi, Michael T. M. Emmerich, and Michel R. V. Chaudron. Problem-specific search operators for metaheuristic software architecture design. Lecture Notes in Computer Science, 7515:267–272, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33119-0_20/.

Anonymous:2012:FMi

[189] Anonymous. Front matter. Lecture Notes in Computer Science, 7515: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-33119-0/1.

Luna:2012:CVE

[190] Andrés Saraos Luna, Valérie Gouranton, and Bruno Arnaldi. Collaborative virtual environments for training: a unified interaction model for real humans and virtual humans. Lecture Notes in Computer Science, 7516:1–12, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33466-5_1/.

Gedeon:2012:DSV

[191] Tom Gedeon, Dingyun Zhu, and Stephane Bersot. Developing a situated virtual reality simulation for telerobotic control and training. Lecture Notes in Computer Science, 7516:13–22, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33466-5_2/.

Wendel:2012:GMC

[192] Viktor Wendel, Stefan Göbel, and Ralf Steinmetz. Game mastering in collaborative multiplayer serious games. Lecture Notes in Computer Science, 7516:23–34, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33466-5_3/.

Berger:2012:IHR

[193] Florian Berger and Wolfgang Müller. Implementing high-resolution adaptivity in game-based learning. Lecture Notes in Computer Science, 7516: 35–40, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33466-5_4/.

Brom:2012:THS

[194] Cyril Brom, Vít Šisler, Michaela Buchtová, Daniel Klement, and David Levčík. Turning high-schools into laboratories? Lessons learnt from studies of instructional effectiveness of digital games in the curricular schooling system. Lecture Notes in Computer Science, 7516:41–53, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33466-5_5/.

Szweda:2012:ANB

[195] Łukasz Szweda, Daniel Wilusz, and Jakub Flotyński. Application of NXT based robots for teaching Java-based concurrency. Lecture Notes in Computer Science, 7516:54–64, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33466-5_6/.

Plass:2012:ELM

[196] Jan L. Plass, Bruce D. Homer, Elizabeth O. Hayward, Jonathan Frye, and Tsu-Ting Huang. The effect of learning mechanics design on learning outcomes in a computer-based geometry game. Lecture Notes in Computer Science, 7516:65–71, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33466-5_7/.

Mildner:2012:SGA

[197] Philip Mildner, Christopher Campbell, Mark Himmelsbach, Christoph Malassa, and Marco Miczka. A serious game for architectural knowledge in the classroom. Lecture Notes in Computer Science, 7516:72– 77, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33466-5_8/.

Rausch:2012:ECD

[198] Steven Rausch, Uwe Fasshauer, and Alke Martens. Evaluation of competence development in WoW. Lecture Notes in Computer Science, 7516: 78–88, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33466-5_9/.

Giannakos:2012:GGP

[199] Michail N. Giannakos, Konstantinos Chorianopoulos, and Letizia Jaccheri. “this game is girly!” perceived enjoyment and student acceptance of edutainment. Lecture Notes in Computer Science, 7516:89–98, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33466-5_10/.

Konert:2012:TSG

[200] Johannes Konert, Stefan Göbel, and Ralf Steinmetz. Towards a social game interaction taxonomy. Lecture Notes in Computer Science, 7516: 99–110, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33466-5_11/.

Ganguin:2012:MWM

[201] Sonja Ganguin and Anna Hoblitz. Mobile worlds: Mobile gaming and learning? Lecture Notes in Computer Science, 7516:111–120, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33466-5_12/.

Takemata:2012:STC

[202] Kazuya Takemata, Akiyuki Minamide, and Sumio Nakamura. Science and technology communication activities by using 3D image projection system. Lecture Notes in Computer Science, 7516:121–124, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33466-5_13/.

Pogrzeba:2012:PLC

[203] Loreen Pogrzeba, Markus Wacker, and Bernhard Jung. Potentials of a low-cost motion analysis system for exergames in rehabilitation and sports medicine. Lecture Notes in Computer Science, 7516:125–133, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33466-5_14/.

Pranantha:2012:PIH

[204] Danu Pranantha, Francesco Bellotti, Ricardo Berta, and Alessandro De Gloria. Puzzle-it: An HTML5 serious games platform for education. Lecture Notes in Computer Science, 7516:134–143, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-33466-5_15/.

Mehm:2012:ASA

[205] Florian Mehm, Stefan Göbel, and Ralf Steinmetz. Authoring of serious adventure games in StoryTec. Lecture Notes in Computer Science, 7516: 144–154, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33466-5_16/.

Wu:2012:DIS

[206] Chun-Tsai Wu, Szu-Ming Chung, and Shao-Shiun Chang. Designing an interactive storytelling game. Lecture Notes in Computer Science, 7516: 155–160, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33466-5_17/.

Reuter:2012:TPT

[207] Christian Reuter, Viktor Wendel, Stefan Göbel, and Ralf Steinmetz. Towards puzzle templates for multiplayer adventures. Lecture Notes in Computer Science, 7516:161–163, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33466-5_18/.

Korn:2012:CSU

[208] Oliver Korn, Michael Brach, Albrecht Schmidt, Thomas Hörz, and Robert Konrad. Context-sensitive user-centered scalability: An introduction focusing on exergames and assistive systems in work contexts. Lecture Notes in Computer Science, 7516:164–176, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33466-5_19/.

Martin:2012:IDG

[209] Anna Lisa Martin and Josef Wiemeyer. The impact of different gaming interfaces on spatial experience and spatial presence — a pilot study. Lecture Notes in Computer Science, 7516:177–182, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33466-5_20/.

Anonymous:2012:FMj

[210] Anonymous. Front matter. Lecture Notes in Computer Science, 7516: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-33466-5/1.

Lee:2012:SIM

[211] Deokwoo Lee and Hamid Krim. System identification: 3D measurement using structured light system. Lecture Notes in Computer Science, 7517:1–11, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33140-4_1/.

Nemesin:2012:GIC

[212] Valérian Némesin, Stéphane Derrode, and Amel Benazza-Benyahia. Gradual Iris code construction from close-up eye video. Lecture Notes in Computer Science, 7517:12–23, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33140-4_2/.

Wang:2012:DVA

[213] Xin Wang, Boris Lenseigne, and Pieter Jonker. Depth from vergence and active calibration for humanoid robots. Lecture Notes in Computer Science, 7517:24–35, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33140-4_3/.

Foix:2012:IGV

[214] Sergi Foix, Simon Kriegel, Stefan Fuchs, Guillem Alenyà, and Carme Torras. Information-gain view planning for free-form object reconstruction with a 3D ToF camera. Lecture Notes in Computer Science, 7517: 36–47, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33140-4_4/.

DiCaterina:2012:DES

[215] Gaetano Di Caterina, Iain Hunter, and John J. Soraghan. DSP embedded smart surveillance sensor with robust SWAD-based tracker. Lecture Notes in Computer Science, 7517:48–58, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33140-4_5/.

Karas:2012:GOC

[216] Pavel Karas, David Svoboda, and Pavel Zemčík. GPU optimization of convolution for large 3-D real images. Lecture Notes in Computer Science, 7517:59–71, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33140-4_6/.

Malik:2012:MBF

[217] Krystyna Malik and Bogdan Smolka. Modified bilateral filter for the restoration of noisy color images. Lecture Notes in Computer Science, 7517:72–83, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33140-4_7/.

Soleimani:2012:CSB

[218] Seyfollah Soleimani, Jacob Premkumar Sukumaran, Koen Douterloigne, and Filip Rooms. Correction, stitching and blur estimation of micrographs obtained at high speed. Lecture Notes in Computer Science, 7517: 84–95, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33140-4_8/.

Elhamzi:2012:HIC

[219] Wajdi Elhamzi, Julien Dubois, Johel Miteran, Mohamed Atri, and Rached Tourki. Hardware implementation of a configurable motion estimator for adjusting the video coding performances. Lecture Notes in Computer Science, 7517:96–107, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33140-4_9/.

Huber-Mork:2012:QAD

[220] Reinhold Huber-Mörk and Alexander Schindler. Quality assurance for document image collections in digital preservation. Lecture Notes in Computer Science, 7517:108–119, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33140-4_10/.

Damghanian:2012:SPC

[221] Mitra Damghanian, Roger Olsson, and Mårten Sjöström. The sampling pattern cube — a representation and evaluation tool for optical capturing systems. Lecture Notes in Computer Science, 7517:120–131, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33140-4_11/.

Couillaud:2012:IIA

[222] Julien Couillaud, Alain Horé, and Djemel Ziou. Improving image acquisition: a fish-inspired solution. Lecture Notes in Computer Science, 7517: 132–141, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33140-4_12/.

Cozzolino:2012:EEM

[223] Angelo Cozzolino, Francesco Flammini, Valentina Galli, and Mariangela Lamberti. Evaluating the effects of MJPEG compression on motion tracking in metro railway surveillance. Lecture Notes in Computer Science, 7517:142–154, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-33140-4_13/.

Hradis:2012:AIS

[224] Michal Hradiš, Martin Kolář, Aleš Láník, Jiří Král, Pavel Zemčík, and Pavel Smrž. Annotating images with suggestions— user study of a tagging system. Lecture Notes in Computer Science, 7517:155–166, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33140-4_14/.

Soukup:2012:CCC

[225] Daniel Soukup and Reinhold Huber-Mörk. Cross-channel co-occurrence matrices for robust characterization of surface disruptions in 2\frac{1}{2}D rail image analysis. Lecture Notes in Computer Science, 7517:167–177, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33140-4_15/.

Salas:2012:IHI

[226] Yainuvis Socarrás Salas, David Vázquez Bermudez, and Antonio M. López Peña. Improving HOG with image segmentation: Application to human detection. Lecture Notes in Computer Science, 7517:178–189, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33140-4_16/.

Ghose:2012:SLF

[227] Soumya Ghose, Jhimli Mitra, Arnau Oliver, Robert Martí, Xavier Lladó, and Jordi Freixenet. A supervised learning framework for automatic prostate segmentation in trans rectal ultrasound images. Lecture Notes in Computer Science, 7517:190–200, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33140-4_17/.

Lerme:2012:SSF

[228] Nicolas Lermé and François Malgouyres. Simultaneous segmentation and filtering via reduced graph cuts. Lecture Notes in Computer Science, 7517: 201–212, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33140-4_18/.

Suk:2012:RDB

[229] Tomáš Suk, Cyril Höschl IV, and Jan Flusser. Rectangular decomposition of binary images. Lecture Notes in Computer Science, 7517: 213–224, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33140-4_19/.

Krumnikl:2012:NLS

[230] Michal Krumnikl, Eduard Sojka, and Jan Gaura. A new level-set based algorithm for bimodal depth segmentation. Lecture Notes in Computer Science, 7517:225–236, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33140-4_20/.

Anonymous:2012:FMk

[231] Anonymous. Front matter. Lecture Notes in Computer Science, 7517: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-33140-4/1.

Schewe:2012:TIW

[232] Klaus-Dieter Schewe and Qing Wang. Third International Workshop on Conceptual Modelling of Services (CMS 2012). Lecture Notes in Computer Science, 7518:1–2, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/accesspage/
chapter/10.1007/978-3-642-33999-8_1.

Buchmayr:2012:RBA

[233] Mario Buchmayr, Werner Kurschl, and Josef Küng. A rule based approach for mapping sensor data to ontological models in AAL environments. Lecture Notes in Computer Science, 7518:3–12, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33999-8_2/.

Rady:2012:PSL

[234] Mariam Rady. Parameters for service level agreements generation in cloud computing. Lecture Notes in Computer Science, 7518:13–22, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33999-8_3/.

Andersson:2012:RPU

[235] Birger Andersson, Maria Bergholtz, and Paul Johannesson. Resource, process, and use — views on service modeling. Lecture Notes in Computer Science, 7518:23–33, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33999-8_4/.

Vleju:2012:CCA

[236] Mircea Boris Vleju. A client-centric ASM-based approach to identity management in cloud computing. Lecture Notes in Computer Science, 7518:34–43, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33999-8_5/.

Grandi:2012:IWE

[237] Fabio Grandi, Giorgio Orsi, Letizia Tanca, and Riccardo Torlone. International Workshop on Evolution and Change in Data Management and on Non Conventional Data Access (ECDM — NoCoDa 2012). Lecture Notes in Computer Science, 7518:45–46, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/accesspage/chapter/10.1007/978-3-642-33999-8_6.

Terenziani:2012:TAD

[238] Paolo Terenziani. The Telic/Atelic distinction in temporal databases. Lecture Notes in Computer Science, 7518:47–56, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-33999-8_7/.

Solimando:2012:SAX

[239] Alessandro Solimando, Giorgio Delzanno, and Giovanna Guerrini. Static analysis of XML document adaptations. Lecture Notes in Computer Science, 7518:57–66, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-33999-8_8/.

Gao:2012:SDP

[240] Shi Gao and Carlo Zaniolo. Supporting database provenance under schema evolution. Lecture Notes in Computer Science, 7518:67–77, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33999-8_9/.

Frenot:2012:BMR

[241] Stéphane Frénot and Stéphane Grumbach. An in-browser microblog ranking engine. Lecture Notes in Computer Science, 7518:78–88, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33999-8_10/.

Stefanidis:2012:CRG

[242] Kostas Stefanidis, Nafiseh Shabib, Kjetil Nørvåg, and John Krogstie. Contextual recommendations for groups. Lecture Notes in Computer Science, 7518:89–97, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33999-8_11/.

Gil:2012:FIW

[243] David Gil, Juan Trujillo, and Il-Yeol Song. First international workshop on modeling for data-intensive computing. Lecture Notes in Computer Science, 7518:99, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/accesspage/
chapter/10.1007/978-3-642-33999-8_12.

Porto:2012:SHC

[244] Fabio Porto, Ana Maria de C. Moura, Bernardo Gonçalves, and Ramon Costa. A scientific hypothesis conceptual model. Lecture Notes in Computer Science, 7518:101–110, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33999-8_13/.

Mate:2012:IMM

[245] Alejandro Maté, Hector Llorens, and Elisa de Gregorio. An integrated multidimensional modeling approach to access big data in business intelligence platforms. Lecture Notes in Computer Science, 7518:111–120, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33999-8_14/.

Sultana:2012:CMP

[246] Salmin Sultana and Elisa Bertino. A comprehensive model for provenance. Lecture Notes in Computer Science, 7518:121–130, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33999-8_15/.

Rivero:2012:TDO

[247] Carlos R. Rivero, Inma Hernández, David Ruiz, and Rafael Corchuelo. Towards discovering ontological models from big RDF data. Lecture Notes in Computer Science, 7518:131–140, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33999-8_16/.

Wnuk:2012:TSI

[248] Krzysztof Wnuk, Markus Borg, and Saïd Assar. Towards scalable information modeling of requirements architectures. Lecture Notes in Computer Science, 7518:141–150, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33999-8_17/.

Jureta:2012:PMB

[249] Ivan J. Jureta, Stéphane Faulkner, and Esteban Zimányi. Preface to MORE-BI 2012. Lecture Notes in Computer Science, 7518:151, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/accesspage/chapter/10.1007/978-
3-642-33999-8_18.

Bebel:2012:OLA

[250] Bartosz Bebel, Mikołaj Morzy, Tadeusz Morzy, Zbyszko Królikowski, and Robert Wrembel. OLAP-like analysis of time point-based sequential data. Lecture Notes in Computer Science, 7518:153–161, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33999-8_19/.

Neubock:2012:MDN

[251] Thomas Neuböck, Bernd Neumayr, Thomas Rossgatterer, Stefan Anderlik, and Michael Schrefl. Multi-dimensional navigation modeling using BI analysis graphs. Lecture Notes in Computer Science, 7518: 162–171, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33999-8_20/.

Anonymous:2012:FMl

[252] Anonymous. Front matter. Lecture Notes in Computer Science, 7518: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-33999-8/1.

Alenda:2012:PSL

[253] Régis Alenda and Nicola Olivetti. Preferential semantics for the logic of comparative similarity over triangular and metric models. Lecture Notes in Computer Science, 7519:1–13, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33353-8_1/.

Alenda:2012:NSC

[254] Régis Alenda, Nicola Olivetti, and Gian Luca Pozzato. Nested sequent calculi for conditional logics. Lecture Notes in Computer Science, 7519: 14–27, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33353-8_2/.

Arieli:2012:CTS

[255] Ofer Arieli. Conflict-tolerant semantics for argumentation frameworks. Lecture Notes in Computer Science, 7519:28–40, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-33353-8_3/.

Askounis:2012:KMA

[256] Dimitris Askounis, Costas D. Koutras, and Yorgos Zikos. Knowledge means ‘ all ’, belief means ‘ most ’. Lecture Notes in Computer Science, 7519:41–53, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33353-8_4/.

Aucher:2012:GDS

[257] Guillaume Aucher, Bastien Maubert, and François Schwarzentruber. Generalized DEL-sequents. Lecture Notes in Computer Science, 7519:54– 66, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33353-8_5/.

Balbiani:2012:DBR

[258] Philippe Balbiani and Antoun Yaacoub. Deciding the bisimilarity relation between datalog goals. Lecture Notes in Computer Science, 7519: 67–79, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33353-8_6/.

Beck:2012:IMT

[259] Harald Beck, Thomas Eiter, and Thomas Krennwallner. Inconsistency management for traffic regulations: Formalization and complexity results. Lecture Notes in Computer Science, 7519:80–93, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33353-8_7/.

Andersen:2012:CEP

[260] Mikkel Birkegaard Andersen, Thomas Bolander, and Martin Holm Jensen. Conditional epistemic planning. Lecture Notes in Computer Science, 7519: 94–106, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33353-8_8/.

Booth:2012:PPT

[261] Richard Booth, Thomas Meyer, and Ivan Varzinczak. PTL: a propositional typicality logic. Lecture Notes in Computer Science, 7519:107– 119, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33353-8_9/.

Bozzelli:2012:COA

[262] Laura Bozzelli, Hans van Ditmarsch, and Sophie Pinchinat. The complexity of one-agent refinement modal logic. Lecture Notes in Computer Science, 7519:120–133, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33353-8_10/.

Caroprese:2012:VUP

[263] Luciano Caroprese, Irina Trubitsyna, Mirosław Truszczyński, and Ester Zumpano. The view-update problem for indefinite databases. Lecture Notes in Computer Science, 7519:134–146, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33353-8_11/.

Ciucci:2012:TVL

[264] Davide Ciucci and Didier Dubois. Three-valued logics for incomplete information and epistemic logic. Lecture Notes in Computer Science, 7519: 147–159, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33353-8_12/.

Eiter:2012:EUS

[265] Thomas Eiter, Michael Fink, Thomas Krennwallner, Christoph Redl, and Peter Schüller. Exploiting unfounded sets for HEX-program evaluation. Lecture Notes in Computer Science, 7519:160–175, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33353-8_13/.

Rankooh:2012:USN

[266] Masood Feyzbakhsh Rankooh, Ali Mahjoob, and Gholamreza Ghassem-Sani. Using satisfiability for non-optimal temporal planning. Lecture Notes in Computer Science, 7519:176–188, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33353-8_14/.

Finthammer:2012:HEP

[267] Marc Finthammer and Christoph Beierle. How to exploit parametric uniformity for maximum entropy reasoning in a relational probabilistic logic. Lecture Notes in Computer Science, 7519:189–201, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33353-8_15/.

Franconi:2012:EQR

[268] Enrico Franconi, Volha Kerhet, and Nhung Ngo. Exact query reformulation with first-order ontologies and databases. Lecture Notes in Computer Science, 7519:202–214, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33353-8_16/.

Gabaldon:2012:SSL

[269] Alfredo Gabaldon. A selective semantics for logic programs with preferences. Lecture Notes in Computer Science, 7519:215–227, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33353-8_17/.

Giordano:2012:MMS

[270] Laura Giordano, Valentina Gliozzi, Nicola Olivetti, and Gian Luca Pozzato. A minimal model semantics for nonmonotonic reasoning. Lecture Notes in Computer Science, 7519:228–241, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33353-8_18/.

Godo:2012:ETD

[271] Lluís Godo, Enrico Marchioni, and Pere Pardo. Extending a temporal defeasible argumentation framework with possibilistic weights. Lecture Notes in Computer Science, 7519:242–254, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33353-8_19/.

Golinska-Pilarek:2012:DLO

[272] Joanna Golińska-Pilarek. On decidability of a logic for order of magnitude qualitative reasoning with bidirectional negligibility. Lecture Notes in Computer Science, 7519:255–266, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33353-8_20/.

Anonymous:2012:FMm

[273] Anonymous. Front matter. Lecture Notes in Computer Science, 7519: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-33353-8/1.

Dezfuli:2012:XSC

[274] Mohammad G. Dezfuli and Mostafa S. Haghjoo. Xtream: a system for continuous querying over uncertain data streams. Lecture Notes in Computer Science, 7520:1–15, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33362-0_1/.

Urbanova:2012:AOR

[275] Lucie Urbanova, Vilem Vychodil, and Lena Wiese. Applications of ordinal ranks to flexible query answering. Lecture Notes in Computer Science, 7520:16–29, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33362-0_2/.

Cuzzocrea:2012:MUD

[276] Alfredo Cuzzocrea, Rubén de Juan Marín, Hendrik Decker, and Francesc D. Muñoz-Escoí. Managing uncertainty in databases and scaling it up to concurrent transactions. Lecture Notes in Computer Science, 7520:30–43, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33362-0_3/.

Nouioua:2012:GNS

[277] Farid Nouioua. Generalizing naive and stable semantics in argumentation frameworks with necessities and preferences. Lecture Notes in Computer Science, 7520:44–57, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33362-0_4/.

Amgoud:2012:SSL

[278] Leila Amgoud. Stable semantics in logic-based argumentation. Lecture Notes in Computer Science, 7520:58–71, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33362-0_5/.

Amgoud:2012:OLB

[279] Leila Amgoud. The outcomes of logic-based argumentation systems under preferred semantics. Lecture Notes in Computer Science, 7520: 72–84, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33362-0_6/.

Dvorak:2012:AAM

[280] Wolfgang Dvořák, Stefan Szeider, and Stefan Woltran. Abstract argumentation via monadic second order logic. Lecture Notes in Computer Science, 7520:85–98, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33362-0_7/.

Budan:2012:AAC

[281] Maximiliano C. D. Budán, Mauro Gómez Lucero, Carlos I. Chesñevar, and Guillermo R. Simari. An approach to argumentation considering attacks through time. Lecture Notes in Computer Science, 7520: 99–112, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33362-0_8/.

Chammas:2012:DDC

[282] Antoine Chammas, Moamar Sayed-Mouchaweh, Eric Duviella, and Stéphane Lecoeuche. Drift detection and characterization for fault diagnosis and prognosis of dynamical systems. Lecture Notes in Computer Science, 7520:113–126, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33362-0_9/.

Trawinski:2012:AEG

[283] Bogdan Trawiński, Tadeusz Lasota, Magdalena Smetek, and Grzegorz Trawiński. An attempt to employ genetic fuzzy systems to predict from a data stream of premises transactions. Lecture Notes in Computer Science, 7520:127–140, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33362-0_10/.

Lughofer:2012:NII

[284] Edwin Lughofer. Navigating interpretability issues in evolving fuzzy systems. Lecture Notes in Computer Science, 7520:141–153, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33362-0_11/.

Decan:2012:CCQ

[285] Alexandre Decan, Fabian Pijcke, and Jef Wijsen. Certain conjunctive query answering in SQL. Lecture Notes in Computer Science, 7520: 154–167, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33362-0_12/.

Caroprese:2012:RCP

[286] Luciano Caroprese and Ester Zumpano. Restoring consistency in P2P deductive databases. Lecture Notes in Computer Science, 7520:168– 179, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33362-0_13/.

Gardezi:2012:TCC

[287] Jaffer Gardezi and Leopoldo Bertossi. Tractable cases of clean query answering under entity resolution via matching dependencies. Lecture Notes in Computer Science, 7520:180–193, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33362-0_14/.

Attiaoui:2012:DBC

[288] Dorra Attiaoui, Pierre-Emmanuel Doré, Arnaud Martin, and Boutheina Ben Yaghlane. A distance between continuous belief functions. Lecture Notes in Computer Science, 7520:194–205, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33362-0_15/.

Laamari:2012:CGR

[289] Wafa Laâmari, Boutheina Ben Yaghlane, and Christophe Simon. On the complexity of the graphical representation and the belief inference in the dynamic directed evidential networks with conditional belief functions. Lecture Notes in Computer Science, 7520:206–218, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33362-0_16/.

Ma:2012:RPP

[290] Jianbing Ma, Salem Benferhat, and Weiru Liu. Revision over partial preorders: a postulational study. Lecture Notes in Computer Science, 7520: 219–232, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33362-0_17/.

Hamed:2012:RUP

[291] Mohammad Ghasemi Hamed, Mathieu Serrurier, and Nicolas Durand. Representing uncertainty by possibility distributions encoding confidence bands, tolerance and prediction intervals. Lecture Notes in Computer Science, 7520:233–246, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33362-0_18/.

Kougioumtzoglou:2012:HWB

[292] Ioannis A. Kougioumtzoglou and Pol D. Spanos. Harmonic wavelets based identification of nonlinear and time-variant systems. Lecture Notes in Computer Science, 7520:247–260, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33362-0_19/.

Patelli:2012:UAU

[293] Edoardo Patelli. An upscaling approach for uncertainty quantification of stochastic contaminant transport through fractured media. Lecture Notes in Computer Science, 7520:261–272, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33362-0_20/.

Anonymous:2012:FMn

[294] Anonymous. Front matter. Lecture Notes in Computer Science, 7520: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-33362-0/1.

Akshay:2012:SBD

[295] S. Akshay, Blaise Genest, and Loïc Hélouët. Symbolically bounding the drift in time-constrained MSC graphs. Lecture Notes in Computer Science, 7521:1–15, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32943-2_1/.

DSouza:2012:CHM

[296] Deepak D’Souza and Raj Mohan Matteplackel. A compositional hierarchical monitoring automaton construction for LTL. Lecture Notes in Computer Science, 7521:16–29, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32943-2_2/.

Sanchez:2012:HTE

[297] César Sánchez and Julian Samborski-Forlese. How to translate efficiently extensions of temporal logics into alternating automata. Lecture Notes in Computer Science, 7521:30–45, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32943-2_3/.

Foster:2012:COO

[298] Simon Foster, Ondřej Rypáček, and Georg Struth. Correctness of object oriented models by extended type inference. Lecture Notes in Computer Science, 7521:46–60, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32943-2_4/.

Dai:2012:NTS

[299] Liyun Dai and Bican Xia. Non-termination sets of simple linear loops. Lecture Notes in Computer Science, 7521:61–73, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-32943-2_5/.

Nikolic:2012:DEA

[300] – Durica Nikolić and Fausto Spoto. Definite expression aliasing analysis for Java bytecode. Lecture Notes in Computer Science, 7521:74– 89, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32943-2_6/.

Riesco:2012:USS

[301] Adrián Riesco. Using semantics specified in Maude to generate test cases. Lecture Notes in Computer Science, 7521:90–104, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32943-2_7/.

Sanchez-Gil:2012:LNR

[302] Lidia Sánchez-Gil and Mercedes Hidalgo-Herrero. A locally nameless representation for a natural semantics for lazy evaluation. Lecture Notes in Computer Science, 7521:105–119, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32943-2_8/.

Benes:2012:MPR

[303] Nikola Beneš and Jan Křetínský. Modal process rewrite systems. Lecture Notes in Computer Science, 7521:120–135, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-32943-2_9/.

Riesco:2012:NCS

[304] Adrián Riesco and Juan Rodríguez-Hortalá. S-narrowing for constructor systems. Lecture Notes in Computer Science, 7521:136–150, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32943-2_10/.

Lodha:2012:DPU

[305] Sachin Lodha, Nikhil Patwardhan, and Ashim Roy. Data privacy using MASKETEER^{T M}. Lecture Notes in Computer Science, 7521:151– 158, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32943-2_11/.

Roy:2012:CCT

[306] Suman Roy and Sidharth Bihary. A conformance checker tool CSPConCheck. Lecture Notes in Computer Science, 7521:159–163, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32943-2_12/.

Raviram:2012:STS

[307] S. Raviram, P. Peranandam, and M. Satpathy. SmartTestGen+: a test suite booster for enhanced structural coverage. Lecture Notes in Computer Science, 7521:164–167, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32943-2_13/.

Williams:2012:MCU

[308] David M. Williams and Joeri de Ruiter. Model checking under fairness in ProB and its application to fair exchange protocols. Lecture Notes in Computer Science, 7521:168–182, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32943-2_14/.

Yatake:2012:MCO

[309] Kenro Yatake and Toshiaki Aoki. Model checking of OSEK/VDX OS design model based on environment modeling. Lecture Notes in Computer Science, 7521:183–197, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32943-2_15/.

Cranen:2012:CSP

[310] Sjoerd Cranen and Jeroen J. A. Keiren. A cure for stuttering parity games. Lecture Notes in Computer Science, 7521:198–212, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32943-2_16/.

Caillaud:2012:ERD

[311] Benoît Caillaud and Jean-Baptiste Raclet. Ensuring reachability by design. Lecture Notes in Computer Science, 7521:213–227, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-32943-2_17/.

Peyronnet:2012:AVE

[312] Sylvain Peyronnet and Michel De Rougemont. Approximate verification and enumeration problems. Lecture Notes in Computer Science, 7521: 228–242, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-32943-2_18/.

Mari:2012:UQS

[313] Federico Mari, Igor Melatti, and Ivano Salvo. Undecidability of quantized state feedback control for discrete time linear hybrid systems. Lecture Notes in Computer Science, 7521:243–258, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-32943-2_19/.

Anonymous:2012:BMb

[314] Anonymous. Back matter. Lecture Notes in Computer Science, 7521: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-32943-2/1.

Anonymous:2012:FMo

[315] Anonymous. Front matter. Lecture Notes in Computer Science, 7521: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-32943-2/1.

deLima:2012:SII

[316] Edirlei Soares de Lima, Bruno Feijó, Cesar T. Pozzer, and Angelo E. M. Ciarlini. Social interaction for interactive storytelling. Lecture Notes in Computer Science, 7522:1–15, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33542-6_1/.

Muller:2012:GAD

[317] Ivana Müller, Petra Sundström, Martin Murer, and Manfred Tscheligi. Gaming after dark. Lecture Notes in Computer Science, 7522: 16–29, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33542-6_2/.

daSilva:2012:IGE

[318] Fabio A. Guilherme da Silva, Antonio L. Furtado, and Angelo E. M. Ciarlini. Information-gathering events in story plots. Lecture Notes in Computer Science, 7522:30–44, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33542-6_3/.

Apken:2012:DEP

[319] Daniel Apken, Hendrik Landwehr, Marc Herrlich, Markus Krause, and Dennis Paul. Design and evaluation of parametrizable multi-genre game mechanics. Lecture Notes in Computer Science, 7522:45–52, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33542-6_4/.

Kothgassner:2012:VTT

[320] Oswald D. Kothgassner, Anna Felnhofer, Leon Beutl, Helmut Hlavacs, and Mario Lehenbauer. A virtual training tool for giving talks. Lecture Notes in Computer Science, 7522:53–66, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33542-6_5/.

Sisler:2012:SHC

[321] Vít Šisler, Cyril Brom, Jaroslav Cuhra, Kamil Činátl, and Jakub Gemrot. Stories from the history of Czechoslovakia, a serious game for teaching history of the Czech lands in the 20th century — notes on design concepts and design process. Lecture Notes in Computer Science, 7522: 67–74, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33542-6_6/.

Blake:2012:CPI

[322] Christopher Blake, Dorothée Hefner, Christian Roth, Christoph Klimmt, and Peter Vorderer. Cognitive processes involved in video game identification. Lecture Notes in Computer Science, 7522:75–84, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33542-6_7/.

Moreno:2012:ASS

[323] Alejandro Moreno, Robby van Delden, Dennis Reidsma, Ronald Poppe, and Dirk Heylen. An annotation scheme for social interaction in digital playgrounds. Lecture Notes in Computer Science, 7522:85–99, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33542-6_8/.

Wang:2012:PME

[324] Xuan Wang, Eng Tat Khoo, Sanath Siriwardana, Horathalge Iroshan, and Ryohei Nakatsu. Philosophy meets entertainment: Designing an interactive virtual philosopher. Lecture Notes in Computer Science, 7522: 100–113, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33542-6_9/.

Conroy:2012:SDI

[325] David Conroy, Peta Wyeth, and Daniel Johnson. Spotting the difference: Identifying player opponent preferences in FPS games. Lecture Notes in Computer Science, 7522:114–121, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33542-6_10/.

Lee:2012:APF

[326] Suwon Lee, Jinki Jung, Jihye Hong, J. B. Ryu, and Hyun S. Yang. AR paint: a fusion system of a paint tool and AR. Lecture Notes in Computer Science, 7522:122–129, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33542-6_11/.

Mazalek:2012:PED

[327] Ali Mazalek, Michael Nitsche, Claudia Rébola, Paul Clifton, Andy Wu, and Nick Poirier. Pictures at an exhibition: Design of a hybrid puppetry performance piece. Lecture Notes in Computer Science, 7522: 130–143, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33542-6_12/.

Minuto:2012:FGS

[328] Andrea Minuto, Gijs Huisman, and Anton Nijholt. Follow the grass: a smart material interactive pervasive display. Lecture Notes in Computer Science, 7522:144–157, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33542-6_13/.

Demeulemeester:2012:IVM

[329] Aljosha Demeulemeester, Katriina Kilpi, Shirley A. Elprama, and Sammy Lievens. The ICOCOON virtual meeting room: a virtual environment as a support tool for multipoint teleconference systems. Lecture Notes in Computer Science, 7522:158–171, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33542-6_14/.

Reis:2012:PW

[330] Sofia Reis and Nuno Correia. Playing with the weather. Lecture Notes in Computer Science, 7522:172–184, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33542-6_15/.

Gutierrez:2012:FPL

[331] Lucio Gutierrez, Eleni Stroulia, and Ioanis Nikolaidis. fAARS: a platform for location-aware trans-reality games. Lecture Notes in Computer Science, 7522:185–192, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33542-6_16/.

Mader:2012:HAT

[332] Stéphanie Mader, Stéphane Natkin, and Guillaume Levieux. How to analyse therapeutic games: The player /game /therapy model. Lecture Notes in Computer Science, 7522:193–206, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33542-6_17/.

Matyas:2012:GBT

[333] Sebastian Matyas, Daishi Kato, Takao Shime, Kazuo Kunieda, and Keiji Yamada. Game-based trust. Lecture Notes in Computer Science, 7522: 207–220, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33542-6_18/.

vanDelden:2012:DTG

[334] Robby van Delden, Pauline Aarts, and Betsy van Dijk. Design of tangible games for children undergoing occupational and physical therapy. Lecture Notes in Computer Science, 7522:221–234, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33542-6_19/.

Gerling:2012:GDO

[335] Kathrin Maria Gerling, Frank Paul Schulte, Jan Smeddinck, and Maic Masuch. Game design for older adults: Effects of age-related changes on structural elements of digital games. Lecture Notes in Computer Science, 7522:235–242, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-33542-6_20/.

Anonymous:2012:FMp

[336] Anonymous. Front matter. Lecture Notes in Computer Science, 7522: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-33542-6/1.

Abbeel:2012:MLR

[337] Pieter Abbeel. Machine learning for robotics. Lecture Notes in Computer Science, 7523:1, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/accesspage/
chapter/10.1007/978-3-642-33460-3_1.

DeRaedt:2012:DMM

[338] Luc De Raedt. Declarative modeling for machine learning and data mining. Lecture Notes in Computer Science, 7523:2–3, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/accesspage/chapter/10.1007/978-
3-642-33460-3_2.

Eck:2012:MLM

[339] Douglas Eck. Machine learning methods for music discovery and recommendation. Lecture Notes in Computer Science, 7523:4, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/accesspage/chapter/10.1007/978-
3-642-33460-3_3.

Keim:2012:SPV

[340] Daniel Keim. Solving problems with visual analytics: Challenges and applications. Lecture Notes in Computer Science, 7523:5–6, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/accesspage/chapter/10.1007/978-
3-642-33460-3_4.

Smyth:2012:ATS

[341] Padhraic Smyth. Analyzing text and social network data with probabilistic models. Lecture Notes in Computer Science, 7523:7–8, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/accesspage/chapter/10.1007/978-
3-642-33460-3_5.

Tatti:2012:DDT

[342] Nikolaj Tatti and Jilles Vreeken. Discovering descriptive tile trees. Lecture Notes in Computer Science, 7523:9–24, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33460-3_6/.

Riondato:2012:EDA

[343] Matteo Riondato and Eli Upfal. Efficient discovery of association rules and frequent itemsets through sampling with tight performance guarantees. Lecture Notes in Computer Science, 7523:25–41, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33460-3_7/.

Siebes:2012:SCD

[344] Arno Siebes and René Kersten. Smoothing categorical data. Lecture Notes in Computer Science, 7523:42–57, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33460-3_8/.

Gasse:2012:ECH

[345] Maxime Gasse, Alex Aussem, and Haytham Elghazel. An experimental comparison of hybrid algorithms for Bayesian network structure learning. Lecture Notes in Computer Science, 7523:58–73, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.com/chapter/10.1007/978-3-642-33460-3_9/.

Tschiatschek:2012:BNC

[346] Sebastian Tschiatschek, Peter Reinprecht, Manfred Mücke, and Franz Pernkopf. Bayesian network classifiers with reduced precision parameters. Lecture Notes in Computer Science, 7523:74–89, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33460-3_10/.

Papai:2012:CSP

[347] Tivadar Pápai, Shalini Ghosh, and Henry Kautz. Combining subjective probabilities and data in training Markov logic networks. Lecture Notes in Computer Science, 7523:90–105, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33460-3_11/.

Guo:2012:SBB

[348] Shengbo Guo, Scott Sanner, Thore Graepel, and Wray Buntine. Scorebased Bayesian skill learning. Lecture Notes in Computer Science, 7523: 106–121, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33460-3_12/.

Dogan:2012:NEG

[349] Ürün Dogan, Tobias Glasmachers, and Christian Igel. A note on extending generalization bounds for binary large-margin classifiers to multiple classes. Lecture Notes in Computer Science, 7523:122–129, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33460-3_13/.

Mantrach:2012:ERC

[350] Amin Mantrach and Jean-Michel Renders. Extension of the Rocchio classification method to multi-modal categorization of documents in social media. Lecture Notes in Computer Science, 7523:130–142, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33460-3_14/.

Bootkrajang:2012:LNR

[351] Jakramate Bootkrajang and Ata Kabán. Label-noise robust logistic regression and its applications. Lecture Notes in Computer Science, 7523: 143–158, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33460-3_15/.

Bespalov:2012:SCS

[352] Dmitriy Bespalov, Yanjun Qi, Bing Bai, and Ali Shokoufandeh. Sentiment classification with supervised sequence embedding. Lecture Notes in Computer Science, 7523:159–174, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33460-3_16/.

Edelkamp:2012:BMF

[353] Stefan Edelkamp and Martin Stommel. The bitvector machine: a fast and robust machine learning algorithm for non-linear problems. Lecture Notes in Computer Science, 7523:175–190, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33460-3_17/.

Maes:2012:EMC

[354] Francis Maes, Pierre Geurts, and Louis Wehenkel. Embedding Monte Carlo search of features in tree-based ensemble methods. Lecture Notes in Computer Science, 7523:191–206, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33460-3_18/.

Zhang:2012:HSS

[355] Zhihong Zhang, Edwin R. Hancock, and Xiao Bai. Hypergraph spectra for semi-supervised feature selection. Lecture Notes in Computer Science, 7523:207–222, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-33460-3_19/.

Wang:2012:LNM

[356] Jun Wang, Adam Woznica, and Alexandros Kalousis. Learning neighborhoods for metric learning. Lecture Notes in Computer Science, 7523: 223–236, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33460-3_20/.

Anonymous:2012:FMq

[357] Anonymous. Front matter. Lecture Notes in Computer Science, 7523: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-33460-3/1.

Liu:2012:AIF

[358] Ruilin Liu, Hui (Wendy) Wang, Anna Monreale, Dino Pedreschi, Fosca Giannotti, and Wenge Guo. AUDIO: An integrity \underline{Audi}ting framework of \underline{O}utlier-mining-as-a-service systems. Lecture Notes in Computer Science, 7524:1–18, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33486-3_1/.

Vinterbo:2012:DPP

[359] Staal A. Vinterbo. Differentially private projected histograms: Construction and use for prediction. Lecture Notes in Computer Science, 7524: 19–34, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33486-3_2/.

Kamishima:2012:FAC

[360] Toshihiro Kamishima, Shotaro Akaho, Hideki Asoh, and Jun Sakuma. Fairness-aware classifier with prejudice remover regularizer. Lecture Notes in Computer Science, 7524:35–50, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33486-3_3/.

Kirshenbaum:2012:LCM

[361] Evan Kirshenbaum, George Forman, and Michael Dugan. A live comparison of methods for personalized article recommendation at forbes.com. Lecture Notes in Computer Science, 7524:51–66, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-33486-3_4/.

Hidasi:2012:FBT

[362] Balázs Hidasi and Domonkos Tikk. Fast ALS-based tensor factorization for context-aware recommendation from implicit feedback. Lecture Notes in Computer Science, 7524:67–82, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33486-3_5/.

Cheng:2012:PEM

[363] Weiwei Cheng and Eyke Hüllermeier. Probability estimation for multiclass classification based on label ranking. Lecture Notes in Computer Science, 7524:83–98, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33486-3_6/.

Ure:2012:APM

[364] N. Kemal Ure, Alborz Geramifard, Girish Chowdhary, and Jonathan P. How. Adaptive planning for Markov decision processes with uncertain transition models via incremental feature dependency discovery. Lecture Notes in Computer Science, 7524:99–115, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33486-3_7/.

Akrour:2012:AAP

[365] Riad Akrour, Marc Schoenauer, and Michèle Sebag. APRIL: Active preference learning-based reinforcement learning. Lecture Notes in Computer Science, 7524:116–131, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33486-3_8/.

Peters:2012:ADD

[366] Markus Peters, Wolfgang Ketter, Maytal Saar-Tsechansky, and John Collins. Autonomous data-driven decision-making in smart electricity markets. Lecture Notes in Computer Science, 7524:132–147, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33486-3_9/.

Michini:2012:BNI

[367] Bernard Michini and Jonathan P. How. Bayesian nonparametric inverse reinforcement learning. Lecture Notes in Computer Science, 7524: 148–163, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33486-3_10/.

Nguyen:2012:BMC

[368] Truong-Huy Dinh Nguyen, Wee-Sun Lee, and Tze-Yun Leong. Bootstrapping Monte Carlo tree search with an imperfect heuristic. Lecture Notes in Computer Science, 7524:164–179, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33486-3_11/.

Dulac-Arnold:2012:FRL

[369] Gabriel Dulac-Arnold, Ludovic Denoyer, Philippe Preux, and Patrick Gallinari. Fast reinforcement learning with large action sets using errorcorrecting output codes for MDP factorization. Lecture Notes in Computer Science, 7524:180–194, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33486-3_12/.

Ermon:2012:LPB

[370] Stefano Ermon, Yexiang Xue, Carla Gomes, and Bart Selman. Learning policies for battery usage optimization in electric vehicles. Lecture Notes in Computer Science, 7524:195–210, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33486-3_13/.

Ramavajjala:2012:PIB

[371] Vivek Ramavajjala and Charles Elkan. Policy iteration based on a learned transition model. Lecture Notes in Computer Science, 7524: 211–226, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33486-3_14/.

Boularias:2012:SAL

[372] Abdeslam Boularias, Oliver Krömer, and Jan Peters. Structured apprenticeship learning. Lecture Notes in Computer Science, 7524:227– 242, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33486-3_15/.

Gay:2012:BAC

[373] Dominique Gay and Marc Boullé. A Bayesian approach for classification rule mining in quantitative databases. Lecture Notes in Computer Science, 7524:243–259, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-33486-3_16/.

Batal:2012:BST

[374] Iyad Batal, Gregory Cooper, and Milos Hauskrecht. A Bayesian scoring technique for mining predictive and non-spurious rules. Lecture Notes in Computer Science, 7524:260–276, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33486-3_17/.

Lemmerich:2012:GPT

[375] Florian Lemmerich, Martin Becker, and Martin Atzmueller. Generic pattern trees for exhaustive exceptional model mining. Lecture Notes in Computer Science, 7524:277–292, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33486-3_18/.

Iwata:2012:BSS

[376] Tomoharu Iwata and Kevin Duh. Bidirectional semi-supervised learning with graphs. Lecture Notes in Computer Science, 7524:293–306, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33486-3_19/.

Verma:2012:CBS

[377] Saurabh Verma and Estevam R. Hruschka, Jr. Coupled Bayesian sets algorithm for semi-supervised learning and information extraction. Lecture Notes in Computer Science, 7524:307–322, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33486-3_20/.

Anonymous:2012:FMr

[378] Anonymous. Front matter. Lecture Notes in Computer Science, 7524: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-33486-3/1.

Armbrust:2012:VBN

[379] Christopher Armbrust, Lisa Kiekbusch, Thorsten Ropertz, and Karsten Berns. Verification of behaviour networks using finite-state automata. Lecture Notes in Computer Science, 7526:1–12, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-33347-7_1/.

Beckert:2012:FSM

[380] Bernhard Beckert and Daniel Bruns. Formal semantics of model fields in annotation-based specifications. Lecture Notes in Computer Science, 7526:13–24, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33347-7_2/.

Edelkamp:2012:SPB

[381] Stefan Edelkamp, Tim Federholzner, and Peter Kissmann. Searching with partial belief states in general games with incomplete information. Lecture Notes in Computer Science, 7526:25–36, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33347-7_3/.

Federmann:2012:MLF

[382] Christian Federmann. A machine-learning framework for hybrid machine translation. Lecture Notes in Computer Science, 7526:37–48, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33347-7_4/.

Finthammer:2012:UEW

[383] Marc Finthammer and Christoph Beierle. Using equivalences of worlds for aggregation semantics of relational conditionals. Lecture Notes in Computer Science, 7526:49–60, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33347-7_5/.

Hennig:2012:DMI

[384] Ben Hennig and Norbert Reithinger. Developing of a multimodal interactive training system in therapeutic calisthenics for elderly people. Lecture Notes in Computer Science, 7526:61–72, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33347-7_6/.

Jabbari:2012:PLG

[385] Shahin Jabbari, Robert C. Holte, and Sandra Zilles. PAC-learning with general class noise models. Lecture Notes in Computer Science, 7526: 73–84, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33347-7_7/.

Kessler:2012:AMP

[386] Jens Kessler, Jürgen Strobel, and Horst-Michael Gross. Avoiding moving persons by using simple trajectory prediction and spatio temporal planning. Lecture Notes in Computer Science, 7526:85–96, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33347-7_8/.

Kramer:2012:UNN

[387] Oliver Kramer. Unsupervised nearest neighbors with kernels. Lecture Notes in Computer Science, 7526:97–106, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33347-7_9/.

Holldobler:2012:CEP

[388] Steffen Hölldobler, Norbert Manthey, and Peter Steinke. A compact encoding of pseudo-Boolean constraints into SAT. Lecture Notes in Computer Science, 7526:107–118, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33347-7_10/.

Mattar:2012:STM

[389] Nikita Mattar and Ipke Wachsmuth. Small talk is more than chitchat. Lecture Notes in Computer Science, 7526:119–130, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33347-7_11/.

Peherstorfer:2012:CBD

[390] Benjamin Peherstorfer, Dirk Pflüger, and Hans-Joachim Bungartz. Clustering based on density estimation with sparse grids. Lecture Notes in Computer Science, 7526:131–142, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33347-7_12/.

Schiller:2012:CBC

[391] Marvin R. G. Schiller and Fernand R. Gobet. A comparison between cognitive and AI models of blackjack strategy learning. Lecture Notes in Computer Science, 7526:143–155, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33347-7_13/.

Schwering:2012:PRP

[392] Christoph Schwering, Daniel Beck, Stefan Schiffer, and Gerhard Lakemeyer. Plan recognition by program execution in continuous temporal domains. Lecture Notes in Computer Science, 7526:156–167, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33347-7_14/.

Vollmer:2012:MHM

[393] Christian Vollmer, Julian P. Eggert, and Horst-Michael Gross. Modeling human motion trajectories by sparse activation of motion primitives learned from unpartitioned data. Lecture Notes in Computer Science, 7526:168–179, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-33347-7_15/.

Westphal:2012:NQC

[394] Matthias Westphal and Julien Hué. Nogoods in qualitative constraintbased reasoning. Lecture Notes in Computer Science, 7526:180–192, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33347-7_16/.

Zastrau:2012:SGD

[395] David Zastrau and Stefan Edelkamp. Stochastic gradient descent with GPGPU. Lecture Notes in Computer Science, 7526:193–204, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33347-7_17/.

Zhang:2012:IQS

[396] Xiaomin Zhang, Sandra Zilles, and Robert C. Holte. Improved query suggestion by query search. Lecture Notes in Computer Science, 7526: 205–216, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33347-7_18/.

Ozcep:2012:KBR

[397] Özgür Lütfü Özçep. Knowledge-base revision using implications as hypotheses. Lecture Notes in Computer Science, 7526:217–228, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33347-7_19/.

Lee:2012:ICD

[398] Sangkyun Lee. Improving confidence of dual averaging stochastic online learning via aggregation. Lecture Notes in Computer Science, 7526: 229–232, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33347-7_20/.

Anonymous:2012:FMs

[399] Anonymous. Front matter. Lecture Notes in Computer Science, 7526: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-33347-7/1.

Lilis:2012:IRE

[400] Yannis Lilis and Anthony Savidis. Implementing reusable exception handling patterns with compile-time metaprogramming. Lecture Notes in Computer Science, 7527:1–15, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33176-3_1/.

Pereverzeva:2012:CSF

[401] Inna Pereverzeva, Elena Troubitsyna, and Linas Laibinis. A case study in formal development of a fault tolerant multi-robotic system. Lecture Notes in Computer Science, 7527:16–31, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33176-3_2/.

Fayollas:2012:FTI

[402] Camille Fayollas, Jean-Charles Fabre, David Navarre, Philippe Palanque, and Yannick Deleris. Fault-tolerant interactive cockpits for critical applications: Overall approach. Lecture Notes in Computer Science, 7527: 32–46, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33176-3_3/.

Prokhorova:2012:LME

[403] Yuliya Prokhorova and Elena Troubitsyna. Linking modelling in event-B with safety cases. Lecture Notes in Computer Science, 7527:47–62, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33176-3_4/.

Hamid:2012:SLD

[404] Brahim Hamid, Jacob Geisel, Adel Ziani, and David Gonzalez. Safety lifecycle development process modeling for embedded systems- example of railway domain. Lecture Notes in Computer Science, 7527:63– 75, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33176-3_5/.

Khan:2012:LER

[405] Yasir Imtiaz Khan and Matteo Risoldi. Language enrichment for resilient MDE. Lecture Notes in Computer Science, 7527:76–90, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33176-3_6/.

terBeek:2012:AGT

[406] Maurice H. ter Beek, Henry Muccini, and Patrizio Pelliccione. Assumeguarantee testing of evolving software product line architectures. Lecture Notes in Computer Science, 7527:91–105, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33176-3_7/.

Gulcu:2012:FIS

[407] Koray Gülcü, Hasan Sözer, and Barış Aktemur. FAS: Introducing a service for avoiding faults in composite services. Lecture Notes in Computer Science, 7527:106–120, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33176-3_8/.

Gorbenko:2012:DSO

[408] Anatoliy Gorbenko, Alexander Romanovsky, Vyacheslav Kharchenko, and Olga Tarasyuk. Dependability of service-oriented computing: Timeprobabilistic failure modelling. Lecture Notes in Computer Science, 7527: 121–133, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33176-3_9/.

BenHamida:2012:MSC

[409] Amira Ben Hamida, Antonia Bertolino, Antonello Calabrò, and Guglielmo De Angelis. Monitoring service choreographies from multiple sources. Lecture Notes in Computer Science, 7527:134–149, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-33176-3_10/.

Masci:2012:SFI

[410] Paolo Masci, Dominic Furniss, Paul Curzon, Michael D. Harrison, and Ann Blandford. Supporting field investigators with PVS: a case study in the healthcare domain. Lecture Notes in Computer Science, 7527: 150–164, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33176-3_11/.

Ferrari:2012:MBE

[411] Alessio Ferrari, Massimiliano L. Itria, Silvano Chiaradonna, and Giorgio O. Spagnolo. Model-based evaluation of the availability of a CBTC system. Lecture Notes in Computer Science, 7527:165–179, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33176-3_12/.

Anonymous:2012:BMc

[412] Anonymous. Back matter. Lecture Notes in Computer Science, 7527: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-33176-3/1.

Anonymous:2012:FMt

[413] Anonymous. Front matter. Lecture Notes in Computer Science, 7527: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-33176-3/1.

DiLoreto:2012:DPE

[414] Ines Di Loreto, Simone Mora, and Monica Divitini. Don’t panic: Enhancing soft skills for civil protection workers. Lecture Notes in Computer Science, 7528:1–12, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33687-4_1/.

Gekker:2012:HG

[415] Alex Gekker. Health games. Lecture Notes in Computer Science, 7528: 13–30, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33687-4_2/.

Menelas:2012:SGT

[416] Bob-Antoine J. Menelas and Martin J. D. Otis. A serious game for training balance control over different types of soil. Lecture Notes in Computer Science, 7528:31–42, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33687-4_3/.

Ma:2012:CLA

[417] Minhua Ma, Kim Bale, and Paul Rea. Constructionist learning in anatomy education. Lecture Notes in Computer Science, 7528:43–58, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33687-4_4/.

Romero:2012:IIA

[418] Margarida Romero, Mireia Usart, and Maria Popescu. Interdisciplinary and international adaption and personalization of the MetaVals serious games. Lecture Notes in Computer Science, 7528:59–73, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33687-4_5/.

Azadegan:2012:SGA

[419] Aida Azadegan and Johann C. K. H. Riedel. Serious games adoption in corporate training. Lecture Notes in Computer Science, 7528:

74–85, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33687-4_6/.

ElMawas:2012:TPK

[420] Nour El Mawas and Jean-Pierre Cahier. Towards participative and knowledge-intensive serious games. Lecture Notes in Computer Science, 7528:86–97, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33687-4_7/.

vanderSpek:2012:TDC

[421] Erik D. van der Spek. Towards designing for competence and engagement in serious games. Lecture Notes in Computer Science, 7528: 98–109, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33687-4_8/.

Marsh:2012:BGG

[422] Tim Marsh, Li Zhiqiang Nickole, and Eric Klopfer. Blended in-game and off-game learning: Assimilating serious games in the classroom and curriculum. Lecture Notes in Computer Science, 7528:110–122, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33687-4_9/.

Mathew:2012:CGB

[423] Tintu Mathew, Jochen Zange, and Joern Rittweger. A computer game based motivation system for human physiology studies. Lecture Notes in Computer Science, 7528:123–134, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33687-4_10/.

Oliveira:2012:LLC

[424] Manuel Fradinho Oliveira and Heiko Duin. Lessons learnt from contextualized interactive story driven development methodology. Lecture Notes in Computer Science, 7528:135–149, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33687-4_11/.

Sanchez:2012:VPS

[425] Rosa García Sánchez and Alasdair G. Thin. Value propositions for serious games in health and well-being. Lecture Notes in Computer Science, 7528:

150–157, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33687-4_12/.

Bocconi:2012:DTC

[426] Stefano Bocconi and Yulia Bachvarova. Dealing with threshold concepts in serious games for competence development. Lecture Notes in Computer Science, 7528:158–169, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33687-4_13/.

Koplin:2012:BMP

[427] Martin Koplin and Carl Skelton. Betaville — a massively participatory mirror world game. Lecture Notes in Computer Science, 7528: 170–173, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33687-4_14/.

Wahner:2012:LTP

[428] Thorsten Wahner and Moritz Kartheuser. Logical thinking by play using the example of the game “Space Goats”. Lecture Notes in Computer Science, 7528:174–182, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33687-4_15/.

Bry:2012:SSE

[429] François Bry and Christoph Wieser. Squaring and scripting the ESP game: Trimming a GWAP to deep semantics. Lecture Notes in Computer Science, 7528:183–192, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33687-4_16/.

Cerinsek:2012:ACM

[430] Gregor Cerinsek, Heiko Duin, and Fiorella Colombo. The application of the CISD^{2} methodology for the definition of a serious game competencebased learning scenario in the domain of sustainable manufacturing. Lecture Notes in Computer Science, 7528:193–207, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-33687-4_17/.

Seitlinger:2012:EVN

[431] Paul C. Seitlinger and Michael A. Bedek. Evaluating the validity of a noninvasive assessment procedure. Lecture Notes in Computer Science, 7528:

208–218, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33687-4_18/.

Petersen:2012:COE

[432] Sobah A. Petersen and Michael A. Bedek. Challenges and opportunities in evaluating learning in serious games: a look at behavioural aspects. Lecture Notes in Computer Science, 7528:219–230, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33687-4_19/.

Hyndman:2012:AEL

[433] Jennifer Hyndman, Tom Lunney, and Paul Mc Kevitt. AmbiLearn: Enhancing the learning environment for primary school education. Lecture Notes in Computer Science, 7528:231–242, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33687-4_20/.

Anonymous:2012:FMu

[434] Anonymous. Front matter. Lecture Notes in Computer Science, 7528: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-33687-4/1.

Mizutani:2012:ESP

[435] Tetsuya Mizutani, Yuki Shinagawa, Naoki Murakami, and Shigeru Igarashi. Expressive secondo performances of a realtime personcomputer ensemble system. Lecture Notes in Computer Science, 7530: 1–8, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33478-8_1/.

Zhang:2012:DLS

[436] Yunong Zhang, Huarong Wu, Zhijun Zhang, Senbo Fu, and Yonghua Yin. Different-level schemes’ equivalence for self-motion planning of robot manipulators. Lecture Notes in Computer Science, 7530:9–16, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33478-8_2/.

Kong:2012:GIO

[437] Lifang Kong, Ying Zhao, and Xinbin Liu. Grey incidence optimization model based on hybrid differential evolution algorithm. Lecture Notes in Computer Science, 7530:17–24, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33478-8_3/.

Li:2012:DMB

[438] Yongfeng Li and Liping Zhu. A data mining based approach to research the relationship between Kansei and usability: a case study of mobile phones. Lecture Notes in Computer Science, 7530:25–33, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33478-8_4/.

Qin:2012:MUB

[439] Xiongpai Qin. Making use of the big data: Next generation of algorithm trading. Lecture Notes in Computer Science, 7530:34–41, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33478-8_5/.

Cao:2012:DCA

[440] Rongrong Cao, Dongjuan Zhu, Qinqin Huang, Xunheng Wang, and Zongcai Ruan. Dynamic causality analysis on default mode network. Lecture Notes in Computer Science, 7530:42–49, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33478-8_6/.

Ma:2012:CED

[441] Xiaoju Ma, Bo Li, Ying Zhang, and Ming Yan. The canny edge detection and its improvement. Lecture Notes in Computer Science, 7530: 50–58, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33478-8_7/.

Tan:2012:MMM

[442] Taizhe Tan, Hailing Liu, Yinwei Zhan, and Yuzhen Jin. Multimodality medical image registration based on improved I-alpha information (SNI) with gradient. Lecture Notes in Computer Science, 7530: 59–66, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33478-8_8/.

Li:2012:DSH

[443] Lingling Li, Weicheng Xie, Ziyang He, Xin Xu, Changmin Chen, and Xiaorong Cui. Design of smart home control system based on ZigBee and embedded Web technology. Lecture Notes in Computer Science, 7530:

67–74, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33478-8_9/.

Zhang:2012:UGP

[444] Ye Zhang, Kang Cao, Kangrui Wu, and Tenglong Yu. Using Gaussian potential function for underdetermined blind sources separation based on DUET. Lecture Notes in Computer Science, 7530:75–81, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33478-8_10/.

Li:2012:NBT

[445] Xiaochen Li, Xizhong Lou, Ting Peng, Jia Xu, Qian Zhou, and Daorong Wu. A new balancing type of wireless sensor network routing algorithm. Lecture Notes in Computer Science, 7530:82–89, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33478-8_11/.

Xu:2012:DPG

[446] Juanjuan Xu and Jian Wu. Development of portable gait analysis system based on a DSP. Lecture Notes in Computer Science, 7530: 90–96, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33478-8_12/.

Tan:2012:INN

[447] Dongjie Tan and An Zhang. Improved neural networks based method for infrared focal plane arrays nonuniformity correction. Lecture Notes in Computer Science, 7530:97–104, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33478-8_13/.

Wang:2012:ICB

[448] Kegang Wang, Liying Qi, and Guohua Geng. Images classifications based on color-texture feature. Lecture Notes in Computer Science, 7530: 105–112, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33478-8_14/.

Bai:2012:SSR

[449] Yu Bai and Yanlong Liu. A synchronization strengthen RFID authentication protocol based on key array. Lecture Notes in Computer Science, 7530:113–119, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-33478-8_15/.

Yang:2012:SFC

[450] Wenchuan Yang, Bei Jia, and Bowei Cao. Study of a fuzzy comprehensive evaluation problem in cloud computing. Lecture Notes in Computer Science, 7530:120–125, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33478-8_16/.

Wang:2012:KGC

[451] Xiao Wang, Fusheng Yu, Huixin Zhang, and Yuming Liu. Knowledgeguided clustering of large-scale time series under wavelet transformation. Lecture Notes in Computer Science, 7530:126–133, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33478-8_17/.

Yan:2012:SDF

[452] Leiming Yan, Jinwei Wang, Jin Han, and Yuxiang Wang. A significancedriven framework for characterizing and finding evolving patterns of news networks. Lecture Notes in Computer Science, 7530:134–141, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33478-8_18/.

Wang:2012:AAR

[453] Dexing Wang, Qian Xie, Dongmei Huang, and Hongchun Yuan. Analysis of association rule mining on quantitative concept lattice. Lecture Notes in Computer Science, 7530:142–149, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33478-8_19/.

Zhang:2012:CRS

[454] Junbo Zhang, Tianrui Li, and Hongmei Chen. Composite rough sets. Lecture Notes in Computer Science, 7530:150–159, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-33478-8_20/.

Anonymous:2012:FMv

[455] Anonymous. Front matter. Lecture Notes in Computer Science, 7530: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-33478-8/1.

Livshits:2012:FMW

[456] Benjamin Livshits. Finding malware on a Web scale. Lecture Notes in Computer Science, 7531:1–2, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

accesspage/chapter/10.1007/978-3-642-33704-8_1.

Wang:2012:ESR

[457] Zhaohui Wang, Ryan Johnson, Rahul Murmuria, and Angelos Stavrou. Exposing security risks for commercial mobile devices. Lecture Notes in Computer Science, 7531:3–21, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33704-8_2/.

Martinelli:2012:QQE

[458] Fabio Martinelli, Ilaria Matteucci, and Charles Morisset. From qualitative to quantitative enforcement of security policy. Lecture Notes in Computer Science, 7531:22–35, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33704-8_3/.

Cadenhead:2012:DIC

[459] Tyrone Cadenhead, Murat Kantarcioglu, Vaibhav Khadilkar, and Bhavani Thuraisingham. Design and implementation of a cloud-based assured information sharing system. Lecture Notes in Computer Science, 7531:36–50, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33704-8_4/.

Yakovlev:2012:OKD

[460] Victor Yakovlev, Valery Korzhik, Mihail Bakaev, and Guillermo Morales-Luna. Optimization of key distribution protocols based on extractors for noisy channels within active adversaries. Lecture Notes in Computer Science, 7531:51–64, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33704-8_5/.

Tsay:2012:VUL

[461] Joe-Kai Tsay and Stig F. Mjølsnes. A vulnerability in the UMTS and LTE authentication and key agreement protocols. Lecture Notes in Computer Science, 7531:65–76, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33704-8_6/.

Moldovyan:2012:BBD

[462] Alexandr Moldovyan, Nikolay Moldovyan, and Evgenia Novikova. Blind 384-bit digital signature scheme. Lecture Notes in Computer Science, 7531:77–83, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33704-8_7/.

Jin:2012:RRC

[463] Xin Jin, Ravi Sandhu, and Ram Krishnan. RABAC: Role-centric attribute-based access control. Lecture Notes in Computer Science, 7531: 84–96, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33704-8_8/.

Oleshchuk:2012:TAR

[464] Vladimir Oleshchuk. Trust-aware RBAC. Lecture Notes in Computer Science, 7531:97–107, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33704-8_9/.

Grusho:2012:AMI

[465] Alexander Grusho, Nick Grusho, and Elena Timonina. Alternative mechanisms for information security. Lecture Notes in Computer Science, 7531: 108–113, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33704-8_10/.

Desharnais:2012:EIF

[466] Josée Desharnais, Erwanne P. Kanyabwero, and Nadia Tawbi. Enforcing information flow policies by a three-valued analysis. Lecture Notes in Computer Science, 7531:114–129, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33704-8_11/.

Avanesov:2012:TOS

[467] Tigran Avanesov, Yannick Chevalier, Michaël Rusinowitch, and Mathieu Turuani. Towards the orchestration of secured services under nondisclosure policies. Lecture Notes in Computer Science, 7531:130–145, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33704-8_12/.

Chechulin:2012:ANI

[468] Andrey Chechulin, Igor Kotenko, and Vasily Desnitsky. An approach for network information flow analysis for systems of embedded components. Lecture Notes in Computer Science, 7531:146–155, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33704-8_13/.

Granadillo:2012:ICS

[469] Gustavo Gonzalez Granadillo, Hervé Débar, Grégoire Jacob, and Chrystel Gaber. Individual countermeasure selection based on the return on response investment index. Lecture Notes in Computer Science, 7531: 156–170, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33704-8_14/.

Rieke:2012:SRR

[470] Roland Rieke, Luigi Coppolino, Andrew Hutchison, Elsa Prieto, and Chrystel Gaber. Security and reliability requirements for advanced security event management. Lecture Notes in Computer Science, 7531: 171–180, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33704-8_15/.

Schutte:2012:MBS

[471] Julian Schütte, Roland Rieke, and Timo Winkelvos. Model-based security event management. Lecture Notes in Computer Science, 7531: 181–190, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33704-8_16/.

Dolgikh:2012:UBM

[472] Andrey Dolgikh, Tomas Nykodym, Victor Skormin, and Zachary Birnbaum. Using behavioral modeling and customized normalcy profiles as protection against targeted cyber-attacks. Lecture Notes in Computer Science, 7531:191–202, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33704-8_17/.

Mustapha:2012:LHH

[473] Yosra Ben Mustapha, Hervé Débar, and Grégoire Jacob. Limitation of honeypot/honeynet databases to enhance alert correlation. Lecture Notes in Computer Science, 7531:203–217, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33704-8_18/.

Zegzhda:2012:SMI

[474] Dmitry P. Zegzhda and Tatiana V. Stepanova. Stochastic model of interaction between botnets and distributed computer defense systems. Lecture Notes in Computer Science, 7531:218–225, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33704-8_19/.

Yavvari:2012:MCU

[475] Chaitanya Yavvari, Arnur Tokhtabayev, Huzefa Rangwala, and Angelos Stavrou. Malware characterization using behavioral components. Lecture Notes in Computer Science, 7531:226–239, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33704-8_20/.

Anonymous:2012:FMw

[476] Anonymous. Front matter. Lecture Notes in Computer Science, 7531: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-33704-8/1.

Ceri:2012:MMB

[477] Stefano Ceri, Emanuele Della Valle, Dino Pedreschi, and Roberto Trasarti. Mega-modeling for big data analytics. Lecture Notes in Computer Science, 7532:1–15, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34002-4_1/.

Cao:2012:SKQ

[478] Xin Cao, Lisi Chen, Gao Cong, Christian S. Jensen, Qiang Qu, and Anders Skovsgaard. Spatial keyword querying. Lecture Notes in Computer Science, 7532:16–29, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-34002-4_2/.

Dayal:2012:CDH

[479] Umeshwar Dayal, Chetan Gupta, Malu Castellanos, Song Wang, and Manolo Garcia-Solaco. Of cubes, DAGs and hierarchical correlations: a novel conceptual model for analyzing social media data. Lecture Notes in Computer Science, 7532:30–49, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34002-4_3/.

Villegas:2012:UCE

[480] Antonio Villegas, Antoni Olivé, and Maria-Ribera Sancho. Understanding constraint expressions in large conceptual schemas by automatic filtering. Lecture Notes in Computer Science, 7532:50–63, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34002-4_4/.

Houy:2012:UUC

[481] Constantin Houy, Peter Fettke, and Peter Loos. Understanding understandability of conceptual models — what are we actually talking about? Lecture Notes in Computer Science, 7532:64–77, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34002-4_5/.

Stark:2012:RCC

[482] Jeannette Stark and Werner Esswein. Rules from cognition for conceptual modelling. Lecture Notes in Computer Science, 7532:78–87, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34002-4_6/.

Anderlik:2012:UDO

[483] Stefan Anderlik, Bernd Neumayr, and Michael Schrefl. Using domain ontologies as semantic dimensions in data warehouses. Lecture Notes in Computer Science, 7532:88–101, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34002-4_7/.

Sekhavat:2012:SCS

[484] Yoones A. Sekhavat and Jeffrey Parsons. Sliced column-store (SCS): Ontological foundations and practical implications. Lecture Notes in Computer Science, 7532:102–115, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-34002-4_8/.

Albrecht:2012:SDL

[485] Alexander Albrecht and Felix Naumann. Schema decryption for large extract-transform-load systems. Lecture Notes in Computer Science, 7532: 116–125, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34002-4_9/.

Ntoutsi:2012:FGR

[486] Eirini Ntoutsi, Kostas Stefanidis, Kjetil Nørvåg, and Hans-Peter Kriegel. Fast group recommendations by applying user clustering. Lecture Notes in Computer Science, 7532:126–140, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34002-4_10/.

Wang:2012:UTW

[487] Jingjing Wang, Haixun Wang, Zhongyuan Wang, and Kenny Q. Zhu. Understanding tables on the Web. Lecture Notes in Computer Science, 7532:141–155, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-34002-4_11/.

Yogev:2012:BGT

[488] Sivan Yogev and Haggai Roitman. Bridging the gaps towards advanced data discovery over semi-structured data. Lecture Notes in Computer Science, 7532:156–165, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-34002-4_12/.

Hernandez:2012:TDC

[489] Inma Hernández, Carlos R. Rivero, David Ruiz, and Rafael Corchuelo. Towards discovering conceptual models behind Web sites. Lecture Notes in Computer Science, 7532:166–175, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34002-4_13/.

Li:2012:DBS

[490] Sheng Li, Junhu Wang, Kewen Wang, and Jiang Li. A distance-based spelling suggestion method for XML keyword search. Lecture Notes in Computer Science, 7532:176–189, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34002-4_14/.

Embley:2012:CLH

[491] David W. Embley, Stephen W. Liddle, Deryle W. Lonsdale, Joseph S. Park, and Byung-Joo Shin. Cross-language hybrid keyword and semantic search. Lecture Notes in Computer Science, 7532:190–203, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34002-4_15/.

Britell:2012:MSW

[492] Scott Britell and Lois M. L. Delcambre. Mapping semantic widgets to Web-based, domain-specific collections. Lecture Notes in Computer Science, 7532:204–213, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-34002-4_16/.

Guentert:2012:EMS

[493] Markus Guentert, Matthias Kunze, and Mathias Weske. Evaluation measures for similarity search results in process model repositories. Lecture Notes in Computer Science, 7532:214–227, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-34002-4_17/.

Cervera:2012:MAP

[494] Mario Cervera, Manoli Albert, Victoria Torres, and Vicente Pelechano. The MOSKitt4ME approach: Providing process support in a method engineering context. Lecture Notes in Computer Science, 7532:228–241, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-34002-4_18/.

Lumineau:2012:ECD

[495] Nicolas Lumineau, Frédérique Laforest, Yann Gripay, and Jean-Marc Petit. Extending conceptual data model for dynamic environment. Lecture Notes in Computer Science, 7532:242–251, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-34002-4_19/.

Le:2012:CGR

[496] Lam-Son Lê and Aditya Ghose. Contracts + goals = roles? Lecture Notes in Computer Science, 7532:252–266, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34002-4_20/.

Anonymous:2012:FMx

[497] Anonymous. Front matter. Lecture Notes in Computer Science, 7532: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-34002-4/1.

Fouque:2012:IHB

[498] Pierre-Alain Fouque and Mehdi Tibouchi. Indifferentiable hashing to Barreto–Naehrig curves. Lecture Notes in Computer Science, 7533:1– 17, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33481-8_1/.

Mesnager:2012:SBF

[499] Sihem Mesnager. Semi-bent functions with multiple trace terms and hyperelliptic curves. Lecture Notes in Computer Science, 7533:18–36, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33481-8_2/.

Abarzua:2012:CAB

[500] Rodrigo Abarzúa and Nicolas Thériault. Complete atomic blocks for elliptic curves in Jacobian coordinates over prime fields. Lecture Notes in Computer Science, 7533:37–55, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33481-8_3/.

Phan:2012:MBT

[501] Duong Hieu Phan, David Pointcheval, and Mario Strefler. Message-based traitor tracing with optimal ciphertext rate. Lecture Notes in Computer Science, 7533:56–77, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33481-8_4/.

Abdalla:2012:LRS

[502] Michel Abdalla and Jill-Jênn Vie. Leakage-resilient spatial encryption. Lecture Notes in Computer Science, 7533:78–99, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-33481-8_5/.

Boureanu:2012:PFA

[503] Ioana Boureanu, Aikaterini Mitrokotsa, and Serge Vaudenay. On the pseudorandom function assumption in (secure) distance-bounding protocols. Lecture Notes in Computer Science, 7533:100–120, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33481-8_6/.

Abdalla:2012:LBH

[504] Michel Abdalla, Angelo De Caro, and Karina Mochetti. Lattice-based hierarchical inner product encryption. Lecture Notes in Computer Science, 7533:121–138, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-33481-8_7/.

Poppelmann:2012:TEA

[505] Thomas Pöppelmann and Tim Güneysu. Towards efficient arithmetic for lattice-based cryptography on reconfigurable hardware. Lecture Notes in Computer Science, 7533:139–158, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33481-8_8/.

Bernstein:2012:SIN

[506] Daniel J. Bernstein, Tanja Lange, and Peter Schwabe. The security impact of a new cryptographic library. Lecture Notes in Computer Science, 7533: 159–176, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33481-8_9/.

Aranha:2012:FIS

[507] Diego F. Aranha, Armando Faz-Hernández, Julio López, and Francisco Rodríguez-Henríquez. Faster implementation of scalar multiplication on Koblitz curves. Lecture Notes in Computer Science, 7533: 177–193, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33481-8_10/.

Nachef:2012:ZKM

[508] Valérie Nachef, Jacques Patarin, and Emmanuel Volte. Zero-knowledge for multivariate polynomials. Lecture Notes in Computer Science, 7533: 194–213, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33481-8_11/.

Dixon:2012:IEK

[509] Vanessa Dixon, Michael J. Jacobson Jr., and Renate Scheidler. Improved exponentiation and key agreement in the infrastructure of a real quadratic field. Lecture Notes in Computer Science, 7533:214–233, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33481-8_12/.

Barhum:2012:UOT

[510] Kfir Barhum and Ueli Maurer. UOWHFs from OWFs: Trading regularity for efficiency. Lecture Notes in Computer Science, 7533:234– 253, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33481-8_13/.

MacFie:2012:RMR

[511] Andrew MacFie and Daniel Panario. Random mappings with restricted preimages. Lecture Notes in Computer Science, 7533:254–270, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33481-8_14/.

Kircanski:2012:SRK

[512] Aleksandar Kircanski and Amr M. Youssef. On the Sosemanuk related key-IV sets. Lecture Notes in Computer Science, 7533:271–287, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33481-8_15/.

Gouvea:2012:HSI

[513] Conrado P. L. Gouvêa and Julio López. High speed implementation of authenticated encryption for the MSP430X microcontroller. Lecture Notes in Computer Science, 7533:288–304, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33481-8_16/.

Gierlichs:2012:ICD

[514] Benedikt Gierlichs, Jörn-Marc Schmidt, and Michael Tunstall. Infective computation and dummy rounds: Fault protection for block ciphers without check-before-output. Lecture Notes in Computer Science, 7533: 305–321, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33481-8_17/.

Anonymous:2012:BMd

[515] Anonymous. Back matter. Lecture Notes in Computer Science, 7533: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-33481-8/1.

Anonymous:2012:FMy

[516] Anonymous. Front matter. Lecture Notes in Computer Science, 7533: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-33481-8/1.

Bernt:2012:PIP

[517] Matthias Bernt, Kun-Mao Chao, and Jyun-Wei Kao. Preserving inversion phylogeny reconstruction. Lecture Notes in Computer Science, 7534:1–13, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33122-0_1/.

Brown:2012:FPT

[518] Daniel G. Brown and Jakub Truszkowski. Fast phylogenetic tree reconstruction using locality-sensitive hashing. Lecture Notes in Computer Science, 7534:14–29, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-33122-0_2/.

Tsirogiannis:2012:ECP

[519] Constantinos Tsirogiannis, Brody Sandel, and Dimitris Cheliotis. Efficient computation of popular phylogenetic tree measures. Lecture Notes in Computer Science, 7534:30–43, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33122-0_3/.

Brown:2012:SFH

[520] Daniel G. Brown and Daniel Dexter. SibJoin: a fast heuristic for half-sibling reconstruction. Lecture Notes in Computer Science, 7534: 44–56, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33122-0_4/.

Zhu:2012:REM

[521] Yun Zhu and Luay Nakhleh. Reconstructing the evolution of molecular interaction networks under the DMC and link dynamics models. Lecture Notes in Computer Science, 7534:57–68, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33122-0_5/.

Halldorsson:2012:EPS

[522] Bjarni V. Halldórsson, Dima Blokh, and Roded Sharan. Estimating population size via line graph reconstruction. Lecture Notes in Computer Science, 7534:69–80, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33122-0_6/.

Deepak:2012:ECF

[523] Akshay Deepak, David Fernández-Baca, and Michelle M. McMahon. Extracting conflict-free information from multi-labeled trees. Lecture Notes in Computer Science, 7534:81–92, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33122-0_7/.

Gysel:2012:RPU

[524] Rob Gysel, Kristian Stevens, and Dan Gusfield. Reducing problems in unrooted tree compatibility to restricted triangulations of intersection graphs. Lecture Notes in Computer Science, 7534:93–105, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33122-0_8/.

Lafond:2012:ORA

[525] Manuel Lafond, Krister M. Swenson, and Nadia El-Mabrouk. An optimal reconciliation algorithm for gene trees with polytomies. Lecture Notes in Computer Science, 7534:106–122, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33122-0_9/.

Nguyen:2012:AGT

[526] Thi Hau Nguyen, Jean-Philippe Doyon, and Stéphanie Pointet. Accounting for gene tree uncertainties improves gene trees and reconciliation inference. Lecture Notes in Computer Science, 7534:123–134, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33122-0_10/.

Milo:2012:RTC

[527] Nimrod Milo, Shay Zakov, Erez Katzenelson, and Eitan Bachmat. RNA tree comparisons via unrooted unordered alignments. Lecture Notes in Computer Science, 7534:135–148, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33122-0_11/.

Rinaudo:2012:TDP

[528] Philippe Rinaudo, Yann Ponty, Dominique Barth, and Alain Denise. Tree decomposition and parameterized algorithms for RNA structure-sequence alignment including tertiary interactions and pseudoknots. Lecture Notes in Computer Science, 7534:149–164, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33122-0_12/.

Bhar:2012:TET

[529] Anirban Bhar, Martin Haubrock, and Anirban Mukhopadhyay. δ-TRIMAX: Extracting triclusters and analysing coregulation in time series gene expression data. Lecture Notes in Computer Science, 7534: 165–177, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33122-0_13/.

Lin:2012:CAC

[530] Yen-Yi Lin, Phuong Dao, Faraz Hach, Marzieh Bakhshi, and Fan Mo. CLIIQ: Accurate comparative detection and quantification of expressed isoforms in a population. Lecture Notes in Computer Science, 7534: 178–189, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33122-0_14/.

Shutters:2012:ILB

[531] Brad Shutters, Sudheer Vakati, and David Fernández-Baca. Improved lower bounds on the compatibility of quartets, triplets, and multistate characters. Lecture Notes in Computer Science, 7534:190–200, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33122-0_15/.

Tabei:2012:SMT

[532] Yasuo Tabei. Succinct multibit tree: Compact representation of multibit trees by using succinct data structures in chemical fingerprint searches. Lecture Notes in Computer Science, 7534:201–213, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.com/chapter/10.1007/978-3-642-33122-0_16/.

Cox:2012:CDS

[533] Anthony J. Cox, Tobias Jakobi, and Giovanna Rosone. Comparing DNA sequence collections by direct comparison of compressed text indexes. Lecture Notes in Computer Science, 7534:214–224, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33122-0_17/.

Bowe:2012:SBG

[534] Alexander Bowe, Taku Onodera, and Kunihiko Sadakane. Succinct de Bruijn graphs. Lecture Notes in Computer Science, 7534:225–235, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33122-0_18/.

Chikhi:2012:SEE

[535] Rayan Chikhi and Guillaume Rizk. Space-efficient and exact de Bruijn graph representation based on a Bloom filter. Lecture Notes in Computer Science, 7534:236–248, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33122-0_19/.

Vyahhi:2012:BGR

[536] Nikolay Vyahhi, Alex Pyshkin, Son Pham, and Pavel A. Pevzner. From de Bruijn graphs to rectangle graphs for genome assembly. Lecture Notes in Computer Science, 7534:249–261, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33122-0_20/.

Anonymous:2012:FMz

[537] Anonymous. Front matter. Lecture Notes in Computer Science, 7534: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-33122-0/1.

Bjorklund:2012:PTP

[538] Andreas Björklund. The path taken for k-path. Lecture Notes in Computer Science, 7535:1, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/accesspage/
chapter/10.1007/978-3-642-33293-7_1.

Marx:2012:RTP

[539] Dániel Marx. Randomized techniques for parameterized algorithms. Lecture Notes in Computer Science, 7535:2, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/accesspage/chapter/10.1007/978-3-642-33293-7_2.

Pilipczuk:2012:FMI

[540] Marcin Pilipczuk and Michał Pilipczuk. Finding a maximum induced degenerate subgraph faster than 2^{n}. Lecture Notes in Computer Science, 7535:3–12, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33293-7_3/.

Chen:2012:ETH

[541] Yijia Chen, Kord Eickmeyer, and Jörg Flum. The exponential time hypothesis and the parameterized clique problem. Lecture Notes in Computer Science, 7535:13–24, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33293-7_4/.

Escoffier:2012:NRP

[542] Bruno Escoffier, Jérôme Monnot, and Vangelis Th. Paschos. New results on polynomial inapproximability and fixed parameter approximability of edge dominating set. Lecture Notes in Computer Science, 7535: 25–36, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33293-7_5/.

Bliznets:2012:NAP

[543] Ivan Bliznets and Alexander Golovnev. A new algorithm for parameterized MAX-SAT. Lecture Notes in Computer Science, 7535:37–48, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33293-7_6/.

Bonizzoni:2012:RSC

[544] Paola Bonizzoni, Riccardo Dondi, and Giancarlo Mauri. Restricted and swap common superstring: a parameterized view. Lecture Notes in Computer Science, 7535:49–60, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33293-7_7/.

Kowalik:2012:NHM

[545] Łukasz Kowalik. Nonblocker in H-minor free graphs: Kernelization meets discharging. Lecture Notes in Computer Science, 7535:61–72, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33293-7_8/.

Flum:2012:SDS

[546] Jörg Flum and Moritz Müller. Some definitorial suggestions for parameterized proof complexity. Lecture Notes in Computer Science, 7535: 73–84, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33293-7_9/.

Golovach:2012:EAS

[547] Petr A. Golovach, Pinar Heggernes, and Dieter Kratsch. An exact algorithm for subset feedback vertex set on chordal graphs. Lecture Notes in Computer Science, 7535:85–96, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33293-7_10/.

Fomin:2012:PSM

[548] Fedor V. Fomin, Bart M. P. Jansen, and Michał Pilipczuk. Preprocessing subgraph and minor problems: When does a small vertex cover help? Lecture Notes in Computer Science, 7535:97–108, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33293-7_11/.

Bentz:2012:PTA

[549] Cédric Bentz. A polynomial-time algorithm for planar multicuts with few source-sink pairs. Lecture Notes in Computer Science, 7535:109– 119, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33293-7_12/.

Chakraborty:2012:ICP

[550] Chiranjit Chakraborty and Rahul Santhanam. Instance compression for the polynomial hierarchy and beyond. Lecture Notes in Computer Science, 7535:120–134, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-33293-7_13/.

Adiga:2012:PTP

[551] Abhijin Adiga, Jasine Babu, and L. Sunil Chandran. Polynomial time and parameterized approximation algorithms for boxicity. Lecture Notes in Computer Science, 7535:135–146, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33293-7_14/.

Kaski:2012:HHS

[552] Petteri Kaski, Mikko Koivisto, and Jesper Nederlof. Homomorphic hashing for sparse coefficient extraction. Lecture Notes in Computer Science, 7535:147–158, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-33293-7_15/.

Kaski:2012:FMS

[553] Petteri Kaski, Mikko Koivisto, and Janne H. Korhonen. Fast monotone summation over disjoint sets. Lecture Notes in Computer Science, 7535: 159–170, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33293-7_16/.

Blaser:2012:WCM

[554] Markus Bläser and Radu Curticapean. Weighted counting of k-matchings is #W [1]-hard. Lecture Notes in Computer Science, 7535:171–181, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33293-7_17/.

Kitsunai:2012:CDP

[555] Kenta Kitsunai, Yasuaki Kobayashi, and Keita Komuro. Computing directed pathwidth in O(1.89^{n}) time. Lecture Notes in Computer Science, 7535:182–193, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-33293-7_18/.

Abello:2012:MRC

[556] James Abello, Pavel Klavík, and Jan Kratochvíl. MSOL restricted contractibility to planar graphs. Lecture Notes in Computer Science, 7535: 194–205, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33293-7_19/.

Elberfeld:2012:SCP

[557] Michael Elberfeld, Christoph Stockhusen, and Till Tantau. On the space complexity of parameterized problems. Lecture Notes in Computer Science, 7535:206–217, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33293-7_20/.

Anonymous:2012:FMba

[558] Anonymous. Front matter. Lecture Notes in Computer Science, 7535: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-33293-7/1.

Blanco:2012:WDR

[559] Lorenzo Blanco, Valter Crescenzi, Paolo Merialdo, and Paolo Papotti. Web data reconciliation: Models and experiences. Lecture Notes in Computer Science, 7538:1–15, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34213-4_1/.

Mulwad:2012:DIF

[560] Varish Mulwad, Tim Finin, and Anupam Joshi. A domain independent framework for extracting linked semantic data from tables. Lecture Notes in Computer Science, 7538:16–33, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34213-4_2/.

Unbehauen:2012:KES

[561] Jörg Unbehauen, Sebastian Hellmann, Sören Auer, and Claus Stadler. Knowledge extraction from structured sources. Lecture Notes in Computer Science, 7538:34–52, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-34213-4_3/.

Brambilla:2012:EIG

[562] Marco Brambilla, Stefano Ceri, Nicola Cinefra, Anish Das Sarma, and Fabio Forghieri. Extracting information from Google Fusion Tables. Lecture Notes in Computer Science, 7538:53–67, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-34213-4_4/.

Bozzon:2012:MWD

[563] Alessandro Bozzon, Stefano Ceri, and Sr– dan Zagorac. Materialization of Web data sources. Lecture Notes in Computer Science, 7538: 68–81, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34213-4_5/.

Guerrisi:2012:NLI

[564] Vincenzo Guerrisi, Pietro La Torre, and Silvia Quarteroni. Natural language interfaces to data services. Lecture Notes in Computer Science, 7538:82–97, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34213-4_6/.

Aral:2012:MMD

[565] Atakan Aral, Ilker Zafer Akin, and Marco Brambilla. Mobile multi-domain search over structured Web data. Lecture Notes in Computer Science, 7538:98–110, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34213-4_7/.

Brambilla:2012:CLM

[566] Marco Brambilla and Massimiliano Zanoni. Clustering and labeling of multi-dimensional mixed structured data. Lecture Notes in Computer Science, 7538:111–126, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-34213-4_8/.

Morales-Chaparro:2012:VSR

[567] Rober Morales-Chaparro, Juan Carlos Preciado, and Fernando Sánchez-Figueroa. Visualizing search results: Engineering visual patterns development for the Web. Lecture Notes in Computer Science, 7538: 127–142, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34213-4_9/.

Bozzon:2012:ESA

[568] Alessandro Bozzon, Emanuele Della Valle, and Sara Magliacane. Extending SPARQL algebra to support efficient evaluation of top-K SPARQL queries. Lecture Notes in Computer Science, 7538:143–156, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34213-4_10/.

Castano:2012:TCE

[569] Silvana Castano, Alfio Ferrara, and Stefano Montanelli. Thematic clustering and exploration of linked data. Lecture Notes in Computer Science, 7538:157–175, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-34213-4_11/.

Cohen:2012:SRE

[570] Marcelo Cohen and Daniel Schwabe. Support for reusable explorations of linked data in the semantic Web. Lecture Notes in Computer Science, 7538:176–190, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-34213-4_12/.

Cohen:2012:SPM

[571] Sara Cohen, Benny Kimelfeld, and Georgia Koutrika. A survey on proximity measures for social networks. Lecture Notes in Computer Science, 7538:191–206, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-34213-4_13/.

Bozzon:2012:ESC

[572] Alessandro Bozzon, Marco Brambilla, Stefano Ceri, and Andrea Mauri. Extending search to crowds: a model-driven approach. Lecture Notes in Computer Science, 7538:207–222, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34213-4_14/.

Hees:2012:BCA

[573] Jörn Hees, Thomas Roth-Berghofer, Ralf Biedert, Benjamin Adrian, and Andreas Dengel. BetterRelations: Collecting association strengths for linked data triples with a game. Lecture Notes in Computer Science, 7538:223–239, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-34213-4_15/.

Brambilla:2012:ICR

[574] Marco Brambilla, Sofia Ceppi, Nicola Gatti, and Enrico H. Gerding. An incentive–compatible revenue–sharing mechanism for the economic sustainability of multi–domain search based on advertising. Lecture Notes in Computer Science, 7538:240–254, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34213-4_16/.

Anonymous:2012:BMe

[575] Anonymous. Back matter. Lecture Notes in Computer Science, 7538: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-34213-4/1.

Anonymous:2012:FMbb

[576] Anonymous. Front matter. Lecture Notes in Computer Science, 7538: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-34213-4/1.

Broy:2012:CPS

[577] Manfred Broy, María Victoria Cengarle, and Eva Geisberger. Cyberphysical systems: Imminent challenges. Lecture Notes in Computer Science, 7539:1–28, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-34059-8_1/.

Cliff:2012:GFM

[578] Dave Cliff and Linda Northrop. The global financial markets: An ultralarge-scale systems perspective. Lecture Notes in Computer Science, 7539: 29–70, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34059-8_2/.

Keen:2012:WCP

[579] Justin Keen. What is a care pathway? Lecture Notes in Computer Science, 7539:71–80, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34059-8_3/.

Lange:2012:CCT

[580] Douglas S. Lange, Phillip Verbancsics, Robert S. Gutzwiller, John Reeder, and Cullen Sarles. Command and control of teams of autonomous systems. Lecture Notes in Computer Science, 7539:81–93, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34059-8_4/.

McDermid:2012:RLO

[581] John A. McDermid. The risks of LSCITS: The odds are stacked against us. Lecture Notes in Computer Science, 7539:94–117, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34059-8_5/.

Autili:2012:IAS

[582] Marco Autili, Vittorio Cortellessa, Davide Di Ruscio, and Paola Inverardi. Integration architecture synthesis for taming uncertainty in the digital space. Lecture Notes in Computer Science, 7539:118–131, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34059-8_6/.

Chen:2012:SNI

[583] Bangdao Chen and A. W. Roscoe. Social networks for importing and exporting security. Lecture Notes in Computer Science, 7539:132–147, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-34059-8_7/.

Faleiro:2012:CPM

[584] Jose Faleiro, Sriram Rajamani, Kaushik Rajan, G. Ramalingam, and Kapil Vaswani. CScale — a programming model for scalable and reliable distributed applications. Lecture Notes in Computer Science, 7539: 148–156, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34059-8_8/.

Garlan:2012:FTE

[585] David Garlan, Vishal Dwivedi, Ivan Ruchkin, and Bradley Schmerl. Foundations and tools for end-user architecting. Lecture Notes in Computer Science, 7539:157–182, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-34059-8_9/.

Haber:2012:EDO

[586] Arne Haber, Holger Rendel, Bernhard Rumpe, and Ina Schaefer. Evolving delta-oriented software product line architectures. Lecture Notes in Computer Science, 7539:183–208, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34059-8_10/.

vonHanxleden:2012:MVM

[587] Reinhard von Hanxleden, Edward A. Lee, Christian Motika, and Hauke Fuhrmann. Multi-view modeling and pragmatics in 2020. Lecture Notes in Computer Science, 7539:209–223, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34059-8_11/.

Hennicker:2012:VBD

[588] Rolf Hennicker and Matthias Ludwig. View-based development of a simulation framework for multi-disciplinary environmental modelling. Lecture Notes in Computer Science, 7539:224–250, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-34059-8_12/.

Paige:2012:RCT

[589] Richard F. Paige, Phillip J. Brooke, Xiaocheng Ge, and Christopher D. S. Power. Revealing complexity through domain-specific modelling and analysis. Lecture Notes in Computer Science, 7539:251–265, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34059-8_13/.

Sommerville:2012:IRE

[590] Ian Sommerville, Russell Lock, and Tim Storer. Information requirements for enterprise systems. Lecture Notes in Computer Science, 7539: 266–282, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34059-8_14/.

Andre:2012:CBI

[591] Étienne André, Kais Klai, Hanen Ochi, and Laure Petrucci. A counterexample-based incremental and modular verification approach. Lecture Notes in Computer Science, 7539:283–302, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.com/chapter/10.1007/978-3-642-34059-8_15/.

Calinescu:2012:CRP

[592] Radu Calinescu, Shinji Kikuchi, and Kenneth Johnson. Compositional reverification of probabilistic safety properties for large-scale complex IT systems. Lecture Notes in Computer Science, 7539:303–329, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-34059-8_16/.

Colange:2012:ESC

[593] Maximilien Colange, Lom-Messan Hillah, Fabrice Kordon, and Pierre Parutto. Extreme symmetries in complex distributed systems: The bag-oriented approach. Lecture Notes in Computer Science, 7539:330– 352, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34059-8_17/.

Drager:2012:TCB

[594] Klaus Dräger and Marta Kwiatkowska. Towards communication-based steering of complex distributed systems. Lecture Notes in Computer Science, 7539:353–368, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-34059-8_18/.

Ghezzi:2012:EAQ

[595] Carlo Ghezzi. Evolution, adaptation, and the quest for incrementality. Lecture Notes in Computer Science, 7539:369–379, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34059-8_19/.

Henzinger:2012:IIV

[596] Thomas A. Henzinger and Dejan Ničković. Independent implementability of viewpoints. Lecture Notes in Computer Science, 7539:380–395, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-34059-8_20/.

Anonymous:2012:BMf

[597] Anonymous. Back matter. Lecture Notes in Computer Science, 7539: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-34059-8/1.

Anonymous:2012:FMbc

[598] Anonymous. Front matter. Lecture Notes in Computer Science, 7539: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-34059-8/1.

Ammar:2012:RLT

[599] Haitham Bou Ammar, Matthew E. Taylor, Karl Tuyls, and Gerhard Weiss. Reinforcement learning transfer using a sparse coded inter-task mapping. Lecture Notes in Computer Science, 7541:1–16, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34799-3_1/.

An:2012:GTS

[600] Bo An and Milind Tambe. Game theory for security: An important challenge for multiagent systems. Lecture Notes in Computer Science, 7541:17–30, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34799-3_2/.

Brys:2012:LCO

[601] Tim Brys, Yann-Michaël De Hauwere, Ann Nowé, and Peter Vrancx. Local coordination in online distributed constraint optimization problems. Lecture Notes in Computer Science, 7541:31–47, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34799-3_3/.

Carrera:2012:IDA

[602] Álvaro Carrera and Carlos A. Iglesias. Improving diagnosis agents with hybrid hypotheses confirmation reasoning techniques. Lecture Notes in Computer Science, 7541:48–62, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34799-3_4/.

Cossentino:2012:TDP

[603] Massimo Cossentino, Carmelo Lodato, Salvatore Lopes, Patrizia Ribino, and Valeria Seidita. Towards a design process for modeling MAS organizations. Lecture Notes in Computer Science, 7541:63–79, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34799-3_5/.

Emele:2012:AST

[604] Chukwuemeka David Emele, Timothy J. Norman, and Simon Parsons. Argumentation strategies for task delegation. Lecture Notes in Computer Science, 7541:80–96, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-34799-3_6/.

Fornara:2012:UOD

[605] Nicoletta Fornara, Daniel Okouya, and Marco Colombetti. Using OWL 2 DL for expressing ACL content and semantics. Lecture Notes in Computer Science, 7541:97–113, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-34799-3_7/.

Gleizes:2012:SAC

[606] Marie-Pierre Gleizes. Self-adaptive complex systems. Lecture Notes in Computer Science, 7541:114–128, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34799-3_8/.

Gratie:2012:ASA

[607] Cristian Gratie and Adina Magda Florea. Argumentation semantics for agents. Lecture Notes in Computer Science, 7541:129–144, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34799-3_9/.

Kaisers:2012:MAL

[608] Michael Kaisers, Daan Bloembergen, and Karl Tuyls. Multi-agent learning and the reinforcement gradient. Lecture Notes in Computer Science, 7541: 145–159, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34799-3_10/.

Kok:2012:MGM

[609] Eric M. Kok, John-Jules Ch. Meyer, Herre van Oostendorp, and Henry Prakken. A methodology for the generation of multi-agent argumentation dialogue scenarios. Lecture Notes in Computer Science, 7541: 160–174, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34799-3_11/.

McBurney:2012:WM

[610] Peter McBurney. What are models for? Lecture Notes in Computer Science, 7541:175–188, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-34799-3_12/.

Melo:2012:QPB

[611] Francisco S. Melo, Matthijs T. J. Spaan, and Stefan J. Witwicki. Query-POMDP: POMDP-based communication in multiagent systems. Lecture Notes in Computer Science, 7541:189–204, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-34799-3_13/.

Persson:2012:MAB

[612] Camille Persson, Gauthier Picard, Fano Ramparany, and Olivier Boissier. A multi-agent based governance of machine-to-machine systems. Lecture Notes in Computer Science, 7541:205–220, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-34799-3_14/.

Seidita:2012:PPF

[613] Valeria Seidita, Massimo Cossentino, and Antonio Chella. A proposal of process fragment definition and documentation. Lecture Notes in Computer Science, 7541:221–237, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34799-3_15/.

Perotto:2012:RIS

[614] Filipo Studzinski Perotto. Recognizing internal states of other agents to anticipate and coordinate interactions. Lecture Notes in Computer Science, 7541:238–258, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-34799-3_16/.

Anonymous:2012:BMg

[615] Anonymous. Back matter. Lecture Notes in Computer Science, 7541: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-34799-3/1.

Anonymous:2012:FMbd

[616] Anonymous. Front matter. Lecture Notes in Computer Science, 7541: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-34799-3/1.

Gratie:2012:FLA

[617] Cristian Gratie and Adina Magda Florea. Fuzzy labeling for argumentation frameworks. Lecture Notes in Computer Science, 7543:1–8, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33152-7_1/.

Kakas:2012:AAB

[618] A. Kakas, L. Amgoud, G. Kern-Isberner, N. Maudet, and P. Moraitis. ABA: Argumentation based agents. Lecture Notes in Computer Science, 7543:9–27, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33152-7_2/.

Kido:2012:PAS

[619] Hiroyuki Kido. Practical argumentation semantics for Pareto optimality and its relationships with values. Lecture Notes in Computer Science, 7543:28–45, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33152-7_3/.

Maruyama:2012:SAM

[620] Yoshifumi Maruyama, Taichi Hasegawa, Takeshi Hagiwara, and Hajime Sawamura. Syncretic argumentation for multi-agents by lattice homomorphism, fusion and sum. Lecture Notes in Computer Science, 7543: 46–65, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33152-7_4/.

Takahashi:2012:SDA

[621] Kazuko Takahashi and Yu Nambu. A semantics for dynamic argumentation frameworks. Lecture Notes in Computer Science, 7543:66– 85, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33152-7_5/.

Bonzon:2012:OMP

[622] Elise Bonzon and Nicolas Maudet. On the outcomes of multiparty persuasion. Lecture Notes in Computer Science, 7543:86–101, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33152-7_6/.

Letia:2012:AJB

[623] Ioan Alfred Letia and Adrian Groza. Arguing with justifications between collaborating agents. Lecture Notes in Computer Science, 7543: 102–116, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33152-7_7/.

vanderWeide:2012:RAD

[624] T. L. van der Weide and F. Dignum. Reasoning about and discussing preferences between arguments. Lecture Notes in Computer Science, 7543: 117–135, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33152-7_8/.

vanderWeide:2012:MCA

[625] T. L. van der Weide, F. Dignum, J.-J. Ch. Meyer, H. Prakken, and G. A. W. Vreeswijk. Multi-criteria argument selection in persuasion dialogues. Lecture Notes in Computer Science, 7543:136–153, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33152-7_9/.

Emele:2012:ASC

[626] Chukwuemeka David Emele, Timothy J. Norman, and Simon Parsons. Argumentation strategies for collaborative plan resourcing. Lecture Notes in Computer Science, 7543:154–173, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33152-7_10/.

Pardo:2012:CDD

[627] Pere Pardo, Sergio Pajares Ferrando, Eva Onaindia, Lluís Godo, and Pilar Dellunde. Cooperative dialogues for defeasible argumentation-based planning. Lecture Notes in Computer Science, 7543:174–193, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33152-7_11/.

Parsons:2012:UAR

[628] Simon Parsons, Elizabeth Sklar, and Peter McBurney. Using argumentation to reason with and about trust. Lecture Notes in Computer Science, 7543:194–212, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33152-7_12/.

Sklar:2012:TAA

[629] Elizabeth Sklar and M. Q. Azhar. Toward the application of argumentation to interactive learning systems. Lecture Notes in Computer Science, 7543:213–230, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33152-7_13/.

Anonymous:2012:BMh

[630] Anonymous. Back matter. Lecture Notes in Computer Science, 7543: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-33152-7/1.

Anonymous:2012:FMbe

[631] Anonymous. Front matter. Lecture Notes in Computer Science, 7543: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-33152-7/1.

Ficarra:2012:LTC

[632] Francisco V. Cipolla Ficarra. Local tourism and cultural heritage internalization: Myths and realities from software. Lecture Notes in Computer Science, 7546:1–14, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33944-8_1/.

Herrera:2012:ECC

[633] Susana I. Herrera and Silvia del V. Zuaín. Emotion and communicability in e-culture applications. Lecture Notes in Computer Science, 7546: 15–24, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33944-8_2/.

Ficarra:2012:EEC

[634] Francisco V. Cipolla Ficarra. The expansion era of the communicability: First nations for the local and global promotion of cultural and natural heritage. Lecture Notes in Computer Science, 7546:25–37, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33944-8_3/.

Digion:2012:CIU

[635] Leda B. Digión and Mabel Sosa. Conceptual integration of usability and communicability for the interface maintenance of E-learning type collaborative systems. Lecture Notes in Computer Science, 7546: 38–48, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33944-8_4/.

Ficarra:2012:TAC

[636] Francisco V. Cipolla Ficarra. Trichotomic analysis: Communicability, distance learning and hypermedia systems off-line. Lecture Notes in Computer Science, 7546:49–63, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33944-8_5/.

Trejo:2012:FOW

[637] Natalia Trejo, Sandra Casas, and Karim Hallar. A feature-oriented WSDL extension for describing grid services. Lecture Notes in Computer Science, 7546:64–72, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33944-8_6/.

Ficarra:2012:DPG

[638] Francisco V. Cipolla Ficarra. Digital photography and geographical information in the Web 2.0: a quality evaluation of the contents. Lecture Notes in Computer Science, 7546:73–88, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33944-8_7/.

Kratky:2012:PNS

[639] Andreas Kratky. Playing nature — a short history of our mediated relationship to nature. Lecture Notes in Computer Science, 7546:89– 98, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33944-8_8/.

Sanchez:2012:VRW

[640] Gonzalo Martín Sánchez. Virtual reconstruction of the wall and alcazar of Molina de Segura. Lecture Notes in Computer Science, 7546: 99–112, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33944-8_9/.

deAbajo:2012:ERD

[641] Beatriz Sainz de Abajo and Lucas D. P. Mendes. eGovernment: Real democracy in digital society. Lecture Notes in Computer Science, 7546: 113–123, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33944-8_10/.

Jimenez:2012:RNS

[642] Mauricio Pérez Jiménez. Reviews on the narrative status of video games. Lecture Notes in Computer Science, 7546:124–132, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33944-8_11/.

Ridwan:2012:SOB

[643] Sohaila Binte Ridwan and Hasan Shahid Ferdous. The state of Open-StreetMap in Bangladesh. Lecture Notes in Computer Science, 7546: 133–143, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33944-8_12/.

Benotti:2012:BVG

[644] Luciana Benotti and Alexandre Denis. Building virtual guides for virtual worlds. Lecture Notes in Computer Science, 7546:144–153, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33944-8_13/.

Cerratto-Pargman:2012:UCD

[645] T. Cerratto-Pargman and O. Knutsson. User centered development of automatic E-mail answering for the public sector. Lecture Notes in Computer Science, 7546:154–156, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33944-8_14/.

Anonymous:2012:BMi

[646] Anonymous. Back matter. Lecture Notes in Computer Science, 7546: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-33944-8/1.

Anonymous:2012:FMbf

[647] Anonymous. Front matter. Lecture Notes in Computer Science, 7546: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-33944-8/1.

Ficarra:2012:AUC

[648] Francisco V. Cipolla Ficarra. The argentinization of the user centered design. Lecture Notes in Computer Science, 7547:1–14, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34010-9_1/.

Clusella:2012:ISC

[649] María M. Clusella and María G. Mitre. Integrated synergy for cultural contents recursivity in e-culture system. Lecture Notes in Computer Science, 7547:15–25, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34010-9_2/.

Buzzi:2012:FSA

[650] M. Claudia Buzzi and Francesco Donini. Federation and security aspects for the management of the EHR in Italy. Lecture Notes in Computer Science, 7547:26–37, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-34010-9_3/.

Rodriguez:2012:SRU

[651] Juan Manuel Rodríguez and Cristian Mateos. Are Smartphones really useful for scientific computing? Lecture Notes in Computer Science, 7547:38–47, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34010-9_4/.

Podgorelec:2012:ORB

[652] Vili Podgorelec. Ontology and rule based inferring on project teams. Lecture Notes in Computer Science, 7547:48–57, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-34010-9_5/.

Pacini:2012:SCC

[653] Elina Pacini, Melisa Ribero, and Cristian Mateos. Simulation on cloud computing infrastructures of parametric studies of nonlinear solids problems. Lecture Notes in Computer Science, 7547:58–70, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34010-9_6/.

Ficarra:2012:NTI

[654] Francisco V. Cipolla Ficarra. New technologies of the information and communication: Analysis of the constructors and destructors of the European educational system. Lecture Notes in Computer Science, 7547: 71–84, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34010-9_7/.

Fares:2012:MAM

[655] Rubén Fares and Rosanna Costaguta. A multi-agent model that promotes team-role balance in computer supported collaborative learning. Lecture Notes in Computer Science, 7547:85–91, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-34010-9_8/.

Yannibelli:2012:MAC

[656] Virginia Yannibelli and Analía Amandi. A memetic algorithm for collaborative learning team formation in the context of software engineering courses. Lecture Notes in Computer Science, 7547:92–103, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34010-9_9/.

Alharbi:2012:EDG

[657] Saad Alharbi. Empirically derived guidelines for audio-visual Email browsing. Lecture Notes in Computer Science, 7547:104–113, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-34010-9_10/.

Corbellini:2012:IAU

[658] Alejandro Corbellini and Silvia Schiaffino. Intelligent analysis of user interactions in a collaborative software engineering context. Lecture Notes in Computer Science, 7547:114–123, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34010-9_11/.

Ficarra:2012:MNG

[659] Francisco V. Cipolla Ficarra. Motivation for next generation of users versus parochialism in software engineering. Lecture Notes in Computer Science, 7547:124–133, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-34010-9_12/.

Costaguta:2012:GSP

[660] Rosanna Costaguta and Elena Durán. Group and students profiles to support collaborative learning in a multiagent model. Lecture Notes in Computer Science, 7547:134–139, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34010-9_13/.

Rodriguez:2012:TSS

[661] Guillermo Rodríguez and Alvaro Soria. Teaching Scrum to software engineering students with virtual reality support. Lecture Notes in Computer Science, 7547:140–150, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-34010-9_14/.

Ficarra:2012:SAI

[662] Francisco V. Cipolla Ficarra and Andreas Kratky. Security of the automatic information on-line: a study of the controls forbid. Lecture Notes in Computer Science, 7547:151–164, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34010-9_15/.

Digion:2012:CUI

[663] Leda B. Digión and Mabel Sosa. Communicability and usability for the interface in e-learning. Lecture Notes in Computer Science, 7547: 165–175, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34010-9_16/.

Nicoletti:2012:TSA

[664] Matías Nicoletti and J. Andrés Diaz-Pace. Towards software architecture documents matching stakeholders’ interests. Lecture Notes in Computer Science, 7547:176–185, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-34010-9_17/.

Salinas:2012:ARB

[665] Sergio Ariel Salinas and Carlos García Garino. An architecture for resource behavior prediction to improve scheduling systems performance on enterprise desktop grids. Lecture Notes in Computer Science, 7547: 186–196, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34010-9_18/.

Logrono:2012:DSA

[666] Juan Francisco Silva Logroño and Luis Berdún. Discrete sequences analysis for detecting software design patterns. Lecture Notes in Computer Science, 7547:197–207, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-34010-9_19/.

Crasso:2012:PMS

[667] Marco Crasso and Cristian Mateos. A programming model for the semantic Web. Lecture Notes in Computer Science, 7547:208–218, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34010-9_20/.

Anonymous:2012:FMbg

[668] Anonymous. Front matter. Lecture Notes in Computer Science, 7547: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-34010-9/1.

Kasabov:2012:MEG

[669] Nikola Kasabov, Stefan Schliebs, and Ammar Mohemmed. Modelling the effect of genes on the dynamics of probabilistic spiking neural networks for computational neurogenetic modelling. Lecture Notes in Computer Science, 7548:1–9, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-35686-5_1/.

DiSerio:2012:BMB

[670] Clelia Di Serio, Danilo Pellin, Alessandro Ambrosi, Ingrid Glad, and Arnoldo Frigessi. Biostatistics meets bioinformatics in integrating information from highdimensional heterogeneous genomic data: Two examples from rare genetic diseases and infectious diseases. Lecture Notes in Computer Science, 7548:10–20, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-35686-5_2/.

Angelini:2012:BMM

[671] Claudia Angelini, Daniela De Canditiis, Marianna Pensky, and Naomi Brownstein. Bayesian models for the multi-sample time-course microarray experiments. Lecture Notes in Computer Science, 7548:21–35, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-35686-5_3/.

Barla:2012:MLP

[672] Annalisa Barla, Giuseppe Jurman, Roberto Visintainer, and Margherita Squillario. A machine learning pipeline for discriminant pathways identification. Lecture Notes in Computer Science, 7548:36–48, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-35686-5_4/.

Lisboa:2012:DHP

[673] Paulo J. G. Lisboa, Ian H. Jarman, Terence A. Etchells, and Simon J. Chambers. Discovering hidden pathways in bioinformatics. Lecture Notes in Computer Science, 7548:49–60, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-35686-5_5/.

Bassani:2012:RMM

[674] Niccolò Bassani, Federico Ambrogi, Cristina Battaglia, and Elia Biganzoli. Reliability of miRNA microarray platforms: An approach based on random effects linear models. Lecture Notes in Computer Science, 7548:61–72, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-35686-5_6/.

Spinelli:2012:BPI

[675] Roberta Spinelli, Rocco Piazza, Alessandra Pirola, Simona Valletta, and Roberta Rostagno. A bioinformatics procedure to identify and annotate somatic mutations in whole-exome sequencing data. Lecture Notes in Computer Science, 7548:73–82, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-35686-5_7/.

Gonzalez-Navarro:2012:FSP

[676] Félix Fernando González-Navarro and Lluís A. Belanche-Muñoz. Feature selection for the prediction and visualization of brain tumor types using proton magnetic resonance spectroscopy data. Lecture Notes in Computer Science, 7548:83–97, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-35686-5_8/.

Ribas:2012:UGM

[677] Vicent J. Ribas, Jesús Caballero López, Anna Sáez de Tejada, and Juan Carlos Ruiz-Rodríguez. On the use of graphical models to study ICU outcome prediction in septic patients treated with statins. Lecture Notes in Computer Science, 7548:98–111, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-35686-5_9/.

Canakoglu:2012:IBI

[678] Arif Canakoglu, Giorgio Ghisalberti, and Marco Masseroli. Integration of biomolecular interaction data in a genomic and proteomic data warehouse to support biomedical knowledge discovery. Lecture Notes in Computer Science, 7548:112–126, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-35686-5_10/.

Savojardo:2012:MLM

[679] Castrense Savojardo, Piero Fariselli, Damiano Piovesan, and Pier Luigi Martelli. Machine-learning methods to predict protein interaction sites in folded proteins. Lecture Notes in Computer Science, 7548:127–135, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-35686-5_11/.

Cardenas:2012:CKB

[680] Martha Ivón Cárdenas, Alfredo Vellido, Iván Olier, Xavier Rovira, and Jesús Giraldo. Complementing kernel-based visualization of protein sequences with their phylogenetic tree. Lecture Notes in Computer Science, 7548:136–149, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-35686-5_12/.

Jancura:2012:DSF

[681] Pavol Jancura, Dimitrios Mavroeidis, and Elena Marchiori. DEEN: a simple and fast algorithm for network community detection. Lecture Notes in Computer Science, 7548:150–163, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-35686-5_13/.

Castiglioni:2012:SSP

[682] Paolo Castiglioni. Self-similarity in physiological time series: New perspectives from the temporal spectrum of scale exponents. Lecture Notes in Computer Science, 7548:164–175, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-35686-5_14/.

Eleuteri:2012:SVM

[683] Antonio Eleuteri and Azzam F. G. Taktak. Support vector machines for survival regression. Lecture Notes in Computer Science, 7548:176– 189, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-35686-5_15/.

Floares:2012:BCT

[684] Alexandru George Floares, Irina Luludachi, Colin Dinney, and Liana Adam. Boosted C5 trees i-biomarkers panel for invasive bladder cancer progression prediction. Lecture Notes in Computer Science, 7548: 190–200, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-35686-5_16/.

Zambelli:2012:FAM

[685] Federico Zambelli and Giulio Pavesi. A faster algorithm for motif finding in sequences from ChIP-Seq data. Lecture Notes in Computer Science, 7548: 201–212, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-35686-5_17/.

Colas:2012:CCP

[686] Fabrice Colas and Jeanine J. Houwing-Duistermaat. Case/control prediction from Illumina Methylation Microarray’s β and two-color channels in the presence of batch effects. Lecture Notes in Computer Science, 7548: 213–225, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-35686-5_18/.

Napolitano:2012:SDC

[687] Francesco Napolitano and Roberto Tagliaferri. Supporting the design, communication and management of bioinformatic protocols through the leaf tool. Lecture Notes in Computer Science, 7548:226–237, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-35686-5_19/.

Chicco:2012:GAP

[688] Davide Chicco, Marco Tagliasacchi, and Marco Masseroli. Genomic annotation prediction based on integrated information. Lecture Notes in Computer Science, 7548:238–252, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-35686-5_20/.

Anonymous:2012:BMj

[689] Anonymous. Back matter. Lecture Notes in Computer Science, 7548: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-35686-5/1.

Anonymous:2012:FMbh

[690] Anonymous. Front matter. Lecture Notes in Computer Science, 7548: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-35686-5/1.

Nyberg:2012:PSA

[691] Kaisa Nyberg. “provable” security against differential and linear cryptanalysis. Lecture Notes in Computer Science, 7549:1–8, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34047-5_1/.

Dinur:2012:IAF

[692] Itai Dinur, Orr Dunkelman, and Adi Shamir. Improved attacks on full GOST. Lecture Notes in Computer Science, 7549:9–28, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34047-5_2/.

Bogdanov:2012:ZCL

[693] Andrey Bogdanov and Meiqin Wang. Zero correlation linear cryptanalysis with reduced data complexity. Lecture Notes in Computer Science, 7549:29–48, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34047-5_3/.

Wang:2012:MSA

[694] Meiqin Wang, Yue Sun, Elmar Tischhauser, and Bart Preneel. A model for structure attacks, with applications to PRESENT and Serpent. Lecture Notes in Computer Science, 7549:49–68, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-34047-5_4/.

Lu:2012:MDL

[695] Jiqiang Lu. A methodology for differential-linear cryptanalysis and its applications. Lecture Notes in Computer Science, 7549:69–89, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34047-5_5/.

Liu:2012:NOI

[696] Ya Liu, Leibo Li, Dawu Gu, Xiaoyun Wang, Zhiqiang Liu, Jiazhe Chen, and Wei Li. New observations on impossible differential cryptanalysis of reduced-round Camellia. Lecture Notes in Computer Science, 7549: 90–109, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34047-5_6/.

Jean:2012:IRA

[697] Jérémy Jean, María Naya-Plasencia, and Thomas Peyrin. Improved rebound attack on the finalist Grøstl. Lecture Notes in Computer Science, 7549:110–126, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-34047-5_7/.

Wu:2012:PPA

[698] Shuang Wu, Dengguo Feng, Wenling Wu, Jian Guo, Le Dong, and Jian Zou. (pseudo) preimage attack on round-reduced Grøstl hash function and others. Lecture Notes in Computer Science, 7549:127–145, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34047-5_8/.

Naya-Plasencia:2012:PCA

[699] María Naya-Plasencia and Thomas Peyrin. Practical cryptanalysis of ARMADILLO2. Lecture Notes in Computer Science, 7549:146–162, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-34047-5_9/.

Wei:2012:SIV

[700] Lei Wei, Thomas Peyrin, Przemysław Sokołowski, San Ling, Josef Pieprzyk, and Huaxiong Wang. On the (in)security of IDEA in various hashing modes. Lecture Notes in Computer Science, 7549:163–179, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-34047-5_10/.

Rogaway:2012:SCS

[701] Phillip Rogaway, Mark Wooding, and Haibin Zhang. The security of ciphertext stealing. Lecture Notes in Computer Science, 7549:180–195, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-34047-5_11/.

Fleischmann:2012:MFA

[702] Ewan Fleischmann, Christian Forler, and Stefan Lucks. McOE: a family of almost foolproof on-line authenticated encryption schemes. Lecture Notes in Computer Science, 7549:196–215, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34047-5_12/.

Saarinen:2012:CAG

[703] Markku-Juhani Olavi Saarinen. Cycling attacks on GCM, GHASH and other polynomial MACs and hashes. Lecture Notes in Computer Science, 7549:216–225, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-34047-5_13/.

Mendel:2012:CAR

[704] Florian Mendel, Tomislav Nad, and Martin Schläffer. Collision attacks on the reduced dual-stream hash function RIPEMD-128. Lecture Notes in Computer Science, 7549:226–243, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34047-5_14/.

Khovratovich:2012:BPA

[705] Dmitry Khovratovich, Christian Rechberger, and Alexandra Savelieva. Bicliques for preimages: Attacks on Skein-512 and the SHA-2 family. Lecture Notes in Computer Science, 7549:244–263, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-34047-5_15/.

Li:2012:CMM

[706] Ji Li, Takanori Isobe, and Kyoji Shibutani. Converting meet-in-the-middle preimage attack into pseudo collision attack: Application to SHA-2. Lecture Notes in Computer Science, 7549:264–286, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34047-5_16/.

Velichkov:2012:USS

[707] Vesselin Velichkov, Nicky Mouha, Christophe De Cannière, and Bart Preneel. UNAF: a special set of additive differences with application to the differential analysis of ARX. Lecture Notes in Computer Science, 7549: 287–305, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34047-5_17/.

Courtois:2012:EAR

[708] Nicolas T. Courtois, Pouyan Sepehrdad, Petr Sušil, and Serge Vaudenay. ElimLin algorithm revisited. Lecture Notes in Computer Science, 7549: 306–325, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34047-5_18/.

Nguyen:2012:SOU

[709] Long Hoang Nguyen and A. W. Roscoe. Short-output universal hash functions and their use in fast and secure data authentication. Lecture Notes in Computer Science, 7549:326–345, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-34047-5_19/.

Heyse:2012:LEA

[710] Stefan Heyse, Eike Kiltz, Vadim Lyubashevsky, Christof Paar, and Krzysztof Pietrzak. Lapin: An efficient authentication protocol based on ring-LPN. Lecture Notes in Computer Science, 7549:346–365, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-34047-5_20/.

Anonymous:2012:FMbi

[711] Anonymous. Front matter. Lecture Notes in Computer Science, 7549: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-34047-5/1.

Nakano:2012:IVA

[712] Yukiko Nakano, Michael Neff, Ana Paiva, and Marilyn Walker, editors. Intelligent Virtual Agents: 12th International Conference, IVA 2012, Santa Cruz, CA, USA, September, 12–14, 2012. Proceedings, volume 7502 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33196-3 (print), 3-642-33197-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ????

URL http://www.springerlink.com/content/978-3-642-33197-8.

Morzy:2012:ADI

[713] Tadeusz Morzy, Theo Härder, and Robert Wrembel, editors. Advances in Databases and Information Systems: 16th East European Conference, ADBIS 2012, Poznań, Poland, September 18–21, 2012. Proceedings, volume 7503 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33073-8 (print), 3-642-33074-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN ???? URL http://www.springerlink.com/content/978-3-642-
33074-2.

Su:2012:IRA

[714] Chun-Yi Su, Subhash Rakheja, and Honghai Liu, editors. Intelligent Robotics and Applications: 5th International Conference, ICIRA 2012, Montreal, Canada, October 3–5, 2012, Proceedings, Part II, volume 7507 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33514-4 (print), 3-642-33515-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN ???? URL http://www.springerlink.com/content/978-3-642-
33515-0.

Yap:2012:MBI

[715] Pew-Thian Yap, Tianming Liu, Dinggang Shen, Carl-Fredrik Westin, and Li Shen, editors. Multimodal Brain Image Analysis: Second International Workshop, MBIA 2012, Held in Conjunction with MICCAI 2012, Nice, France, October 1–5, 2012. Proceedings, volume 7509 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33529-2 (print), 3-642-33530-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.com/content/978-3-642-33530-3.

Ayache:2012:MICa

[716] Nicholas Ayache, Hervé Delingette, Polina Golland, and Kensaku Mori, editors. Medical Image Computing and Computer-Assisted Intervention — MICCAI 2012: 15th International Conference, Nice, France, October 1–5, 2012, Proceedings, Part I, volume 7510 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33414-8 (print), 3-642-33415-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://

www.springerlink.com/content/978-3-642-33415-3.

Ayache:2012:MICb

[717] Nicholas Ayache, Hervé Delingette, Polina Golland, and Kensaku Mori, editors. Medical Image Computing and Computer-Assisted Intervention — MICCAI 2012: 15th International Conference, Nice, France, October 1–5, 2012, Proceedings, Part II, volume 7511 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33417-2 (print), 3-642-33418-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://

www.springerlink.com/content/978-3-642-33418-4.

Ayache:2012:MICc

[718] Nicholas Ayache, Hervé Delingette, Polina Golland, and Kensaku Mori, editors. Medical Image Computing and Computer-Assisted Intervention — MICCAI 2012: 15th International Conference, Nice, France, October 1–5, 2012, Proceedings, Part III, volume 7512 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-33453-9 (print), 3-642-33454-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://www.

springerlink.com/content/978-3-642-33454-2.

Park:2012:NPC

[719] James J. Park, Albert Zomaya, Sang-Soo Yeo, and Sartaj Sahni, editors. Network and Parallel Computing: 9th IFIP International Conference, NPC 2012, Gwangju, Korea, September 6–8, 2012. Proceedings, volume 7513 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-35605-2 (print), 3-642-35606-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN ???? URL http://www.springerlink.com/content/978-3-642-
35606-3.

Fraser:2012:SBS

[720] Gordon Fraser and Jerffeson Teixeira de Souza, editors. Search Based Software Engineering: 4th International Symposium, SSBSE 2012, Riva del Garda, Italy, September 28–30, 2012. Proceedings, volume 7515 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33118-1 (print), 3-642-33119-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ????

URL http://www.springerlink.com/content/978-3-642-33119-0.

Gobel:2012:LGT

[721] Stefan Göbel, Wolfgang Müller, Bodo Urban, and Josef Wiemeyer, editors. E-Learning and Games for Training, Education, Health and Sports: 7th International Conference, Edutainment 2012 and 3rd International Conference, GameDays 2012, Darmstadt, Germany, September 18–20, 2012. Proceedings, volume 7516 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33465-2 (print), 3-642-33466-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://www.springerlink.

com/content/978-3-642-33466-5.

Blanc-Talon:2012:ACI

[722] Jacques Blanc-Talon, Wilfried Philips, Dan Popescu, Paul Scheunders, and Pavel Zemčík, editors. Advanced Concepts for Intelligent Vision Systems: 14th International Conference, ACIVS 2012, Brno, Czech Republic, September 4–7, 2012. Proceedings, volume 7517 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-33139-4 (print), 3-642-33140-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://www.

springerlink.com/content/978-3-642-33140-4.

Castano:2012:ACM

[723] Silvana Castano, Panos Vassiliadis, Laks V. Lakshmanan, and Mong Li Lee, editors. Advances in Conceptual Modeling: ER 2012 Workshops CMS, ECDM-NoCoDA, MoDIC, MORE-BI, RIGiM, SeCoGIS, WISM, Florence, Italy, October 15–18, 2012. Proceedings, volume 7518 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33998-0 (print), 3-642-33999-9 (ebook). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL

http://www.springerlink.com/content/978-3-642-33999-8.

delCerro:2012:LAI

[724] Luis Fariñas del Cerro, Andreas Herzig, and Jérôme Mengin, editors. Logics in Artificial Intelligence: 13th European Conference, JELIA 2012, Toulouse, France, September 26–28, 2012. Proceedings, volume 7519 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33352-4 (print), 3-642-33353-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ????

URL http://www.springerlink.com/content/978-3-642-33353-8.

Hullermeier:2012:SUM

[725] Eyke Hüllermeier, Sebastian Link, Thomas Fober, and Bernhard Seeger, editors. Scalable Uncertainty Management: 6th International Conference, SUM 2012, Marburg, Germany, September 17–19, 2012. Proceedings, volume 7520 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33361-3 (print), 3-642-33362-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN ???? URL http://www.springerlink.com/content/978-3-642-
33362-0.

Roychoudhury:2012:TAC

[726] Abhik Roychoudhury and Meenakshi D’Souza, editors. Theoretical Aspects of Computing — ICTAC 2012: 9th International Colloquium, Bangalore, India, September 24–27, 2012. Proceedings, volume 7521 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32942-X (print), 3-642-32943-8 (ebook). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL

http://www.springerlink.com/content/978-3-642-32943-2.

Herrlich:2012:ECI

[727] Marc Herrlich, Rainer Malaka, and Maic Masuch, editors. Entertainment Computing — ICEC 2012: 11th International Conference, ICEC 2012, Bremen, Germany, September 26–29, 2012. Proceedings, volume 7522 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33541-1 (print), 3-642-33542-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ????

URL http://www.springerlink.com/content/978-3-642-33542-6.

Flach:2012:MLKa

[728] Peter A. Flach, Tijl De Bie, and Nello Cristianini, editors. Machine Learning and Knowledge Discovery in Databases: European Conference, ECML PKDD 2012, Bristol, UK, September 24–28, 2012. Proceedings, Part I, volume 7523 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33459-8 (print), 3-642-33460-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). LCCN ???? URL http://www.springerlink.com/content/
978-3-642-33460-3.

Flach:2012:MLKb

[729] Peter A. Flach, Tijl De Bie, and Nello Cristianini, editors. Machine Learning and Knowledge Discovery in Databases: European Conference, ECML PKDD 2012, Bristol, UK, September 24–28, 2012. Proceedings, Part II, volume 7524 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33485-7 (print), 3-642-33486-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). LCCN ???? URL http://www.springerlink.com/content/
978-3-642-33486-3.

Glimm:2012:KAA

[730] Birte Glimm and Antonio Krüger, editors. KI 2012: Advances in Artificial Intelligence: 35th Annual German Conference on AI, Saarbrücken, Germany, September 24–27, 2012. Proceedings, volume 7526 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33346-X (print), 3-642-33347-8 (ebook). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL

http://www.springerlink.com/content/978-3-642-33347-7.

Avgeriou:2012:SER

[731] Paris Avgeriou, editor. Software Engineering for Resilient Systems: 4th International Workshop, SERENE 2012, Pisa, Italy, September 27–28, 2012. Proceedings, volume 7527 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33175-0 (print), 3-642-33176-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://www.springerlink.

com/content/978-3-642-33176-3.

Ma:2012:SGD

[732] Minhua Ma, Manuel Fradinho Oliveira, Jannicke Baalsrud Hauge, Heiko Duin, and Klaus-Dieter Thoben, editors. Serious Games Development and Applications: Third International Conference, SGDA 2012, Bremen, Germany, September 26–29, 2012. Proceedings, volume 7528 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33686-8 (print), 3-642-33687-6 (ebook). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL

http://www.springerlink.com/content/978-3-642-33687-4.

Lei:2012:AIC

[733] Jingsheng Lei, Fu Lee Wang, Hepu Deng, and Duoqian Miao, editors. Artificial Intelligence and Computational Intelligence: 4th International Conference, AICI 2012, Chengdu, China, October 26–28, 2012. Proceedings, volume 7530 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33477-6 (print), 3-642-33478-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). LCCN ???? URL http://www.springerlink.com/content/
978-3-642-33478-8.

Kotenko:2012:CNS

[734] Igor Kotenko and Victor Skormin, editors. Computer Network Security: 6th International Conference on Mathematical Methods, Models and Architectures for Computer Network Security, MMM-ACNS 2012, St. Petersburg, Russia, October 17–19, 2012. Proceedings, volume 7531 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33703-1 (print), 3-642-33704-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ????

URL http://www.springerlink.com/content/978-3-642-33704-8.

Atzeni:2012:CMI

[735] Paolo Atzeni, David Cheung, and Sudha Ram, editors. Conceptual Modeling: 31st International Conference ER 2012, Florence, Italy, October 15– 18, 2012. Proceedings, volume 7532 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-34001-6 (print), 3-642-34002-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://www.springerlink.

com/content/978-3-642-34002-4.

Hevia:2012:PCL

[736] Alejandro Hevia and Gregory Neven, editors. Progress in Cryptology — LATINCRYPT 2012: 2nd International Conference on Cryptology and Information Security in Latin America, Santiago, Chile, October 7–10, 2012. Proceedings, volume 7533 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33480-6 (print), 3-642-33481-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://www.springerlink.

com/content/978-3-642-33481-8.

Raphael:2012:ABI

[737] Ben Raphael and Jijun Tang, editors. Algorithms in Bioinformatics: 12th International Workshop, WABI 2012, Ljubljana, Slovenia, September 10– 12, 2012. Proceedings, volume 7534 of Lecture Notes in Computer Science.

Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33121-1 (print), 3-642-33122-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://www.springerlink.

com/content/978-3-642-33122-0.

Thilikos:2012:PEC

[738] Dimitrios M. Thilikos and Gerhard J. Woeginger, editors. Parameterized and Exact Computation: 7th International Symposium, IPEC 2012, Ljubljana, Slovenia, September 12–14, 2012. Proceedings, volume 7535 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33292-7 (print), 3-642-33293-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ????

URL http://www.springerlink.com/content/978-3-642-33293-7.

Ceri:2012:SCB

[739] Stefano Ceri and Marco Brambilla, editors. Search Computing: Broadening Web Search, volume 7538 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-34212-4 (print), 3-642-34213-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://www.springerlink.

com/content/978-3-642-34213-4.

Calinescu:2012:LSC

[740] Radu Calinescu and David Garlan, editors. Large-Scale Complex IT Systems. Development, Operation and Management: 17th Monterey Workshop 2012, Oxford, UK, March 19–21, 2012, Revised Selected Papers, volume 7539 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-34058-X (print), 3-642-34059-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN ???? URL http://www.springerlink.com/content/978-3-642-
34059-8.

Cossentino:2012:MAS

[741] Massimo Cossentino, Michael Kaisers, Karl Tuyls, and Gerhard Weiss, editors. Multi-Agent Systems: 9th European Workshop, EUMAS 2011, Maastricht, The Netherlands, November 14–15, 2011. Revised Selected Papers, volume 7541 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-34798-3 (print), 3-642-34799-1 (e-book). ISSN 0302-9743 (print), 1611-

3349 (electronic). LCCN ???? URL http://www.springerlink.com/
content/978-3-642-34799-3.

McBurney:2012:AMA

[742] Peter McBurney, Simon Parsons, and Iyad Rahwan, editors. Argumentation in Multi-Agent Systems: 8th International Workshop, ArgMAS 2011, Taipei, Taiwan, May 3, 2011, Revised Selected Papers, volume 7543 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33151-3 (print), 3-642-33152-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ????

URL http://www.springerlink.com/content/978-3-642-33152-7.

Cipolla-Ficarra:2012:HCI

[743] Francisco Cipolla-Ficarra, Kim Veltman, Huang Chih-Fang, Miguel Cipolla-Ficarra, and Andreas Kratky, editors. Human-Computer Interaction, Tourism and Cultural Heritage: Second International Workshop, HCITOCH 2011, Córdoba, Argentina, September 14–15, 2011, Revised Selected Papers, volume 7546 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33943-3 (print), 3-642-33944-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://www.springerlink.

com/content/978-3-642-33944-8.

Cipolla-Ficarra:2012:ANT

[744] Francisco Cipolla-Ficarra, Kim Veltman, Domen Verber, Miguel Cipolla-Ficarra, and Florian Kammüller, editors. Advances in New Technologies, Interactive Interfaces and Communicability: Second International Conference, ADNTIIC 2011, Huerta Grande, Argentina, December 5–7, 2011, Revised Selected Papers, volume 7547 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-34009-1 (print), 3-642-34010-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://

www.springerlink.com/content/978-3-642-34010-9.

Biganzoli:2012:CIM

[745] Elia Biganzoli, Alfredo Vellido, Federico Ambrogi, and Roberto Tagliaferri, editors. Computational Intelligence Methods for Bioinformatics and Biostatistics: 8th International Meeting, CIBB 2011, Gargnano del Garda, Italy, June 30–July 2, 2011, Revised Selected Papers, volume 7548 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-35685-0 (print), 3-642-35686-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN ???? URL http://www.springerlink.com/content/978-3-642-
35686-5.

Canteaut:2012:FSE

[746] Anne Canteaut, editor. Fast Software Encryption: 19th International Workshop, FSE 2012, Washington, DC, USA, March 19–21, 2012. Revised Selected Papers, volume 7549 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-34046-6 (print), 3-642-34047-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://www.springerlink.

com/content/978-3-642-34047-5.